### v1

# Galled Perfect Transfer Networks

Abstract
Predicting horizontal gene transfers often requires comparative sequence data, but recent work has shown that character-based approaches could also be useful for this task. Notably, perfect transfer networks (PTN) explain the character diversity of a set of taxa for traits that are gained once, rarely lost, but that can be transferred laterally. Characterizing the structure of such characters is an important step towards understanding more complex characters. Although efficient algorithms can infer such networks from character data, they can sometimes predict overly complicated transfer histories.
With the goal of recovering the simplest possible scenarios in this model, we introduce galled perfect transfer networks, which are PTNs that are galled trees. Such networks are useful for characters that are incompatible in terms of tree-like evolution, but that do fit in an almost-tree scenario. We provide polynomial-time algorithms for two problems: deciding whether one can add transfer edges to a tree to transform it into a galled PTN, and deciding whether a set of characters are galled-compatible, that is, they can be explained by some galled PTN. We also analyze a real dataset comprising of a bacterial species trees and KEGG functions as characters, and derive several conclusions on the difficulty of explaining characters in a galled tree, which provide several directions for future research.
Keywords:
Galled trees Horizontal Gene Transfer Algorithms
1 Introduction
Trees have served as a conventional representation of evolution for centuries in biology. However, contemporary evidence has found frequent exchanges of genetic material between co-existing species, indicating that evolution should rather be expressed as a “web of life”. Horizontal Gene Transfer (HGT) is an important force of innovation between and within in all the domains of life [Soucy2015]. They are known to occur routinely between procaryotes [koonin2001horizontal, thomas2005mechanisms] but also happen between different domains. For example, the thermotogale bacteria, which thrive in extreme environments, are believed to have acquired several genes from archea [Gogarten2005, nesbo2001phylogenetic]. HGTs also affect eukaryotes [Keeling2008], with examples including the acquisition of fructophily from bacteria by yeasts [Gonalves2018] and transfers from parasitic plants to their hosts [Wickell2019].
Owing to their central role in evolution, several algorithmic approaches have been developed to identify HGTs [Ravenhall2015]. Parametric methods seek DNA regions that exhibit a signature that differs from the rest of the genome [LAWRENCE20021], whereas phylogenetic methods rely on the comparison of reconstructed species and gene trees, often using reconciliation [Menet2022-ak, Bansal2012]. Some approaches also use sequence divergence patterns to infer timing discrepancies that correspond to transfers [Schaller2021, Gei2018, Lafond2020, Jones2017].
The vast majority of these methods rely on sequence comparisons. However, sequence-based methods are known to struggle when highly divergent sequences are involved, especially in the presence of ancient transfer events [boto2010horizontal]. An alternative is to predict HGTs with characters, which are morphological or molecular traits that a taxa may possess or not. Character-based methods have been successfully applied to recover recombination or hybridation events [gusfield2014recombinatorics]. A fundamental example of character-based data is gene expression, where the trait is whether or not a gene is expressed in a condition of interest [de1995phenotypic, pontes2013configurable, rawat2008novel], which can sometimes exhibit better phylogenetic signals than similarity measures [Alexander2007].
These approaches aim to explain the diversity of a set of taxa $\mathcal{S}$ that each possess a subset of characters from a set $\mathcal{C}$. Ideally, there should be a phylogeny in which, for each character $C\in\mathcal{C}$, the taxa that possess $C$ form a clade. If such a tree exists, it is called a perfect phylogeny [bodlaender1992two, fernandez2001perfect, bafna2003haplotyping, iersel2019third]. In this setting, perfect phylogenies assume that each character has a unique origin (no-homoplasy), and is always inherited vertically once acquired (no-losses). Of course, these are strong assumptions that rarely apply to real biological datasets. However, understanding this theoretical model has led to multiple extensions with practical applications. Examples include the reconstruction of evolution from Short Interspersed Nuclear Elements (SINE) using partial characters [pe2004incomplete]; haplotyping [bafna2003haplotyping]; or the inference of cancer phylogenies, which were modeled as an extension of perfect phylogenies in [el2016inferring], and broadened to Dollo parsimonies in [el2018sphyr]. Therefore, gaining a deeper understanding of such restricted models can often serve as a stepping stone to reconstruct more complex evolutionary scenarios.
When a perfect phylogeny does not exist for a set of characters, one may instead consider network-like structures to explain this diversity. To this end, Nakhleh et al. introduced Perfect Phylogenetic Networks (PPNs) in [nakhleh, nakhlehtesis]. These networks allow multi-state characters and require that, for each character $C$, the network contains a tree in which nodes in the same state are connected. This is a powerful model that is, unfortunately, difficult to work with, since even deciding whether a known network explains a set of characters is NP-hard. We note that in parallel to our work, this question was studied in the context of binary characters on galled trees [warnow2024statistically], which we discuss in more detail below.
Importantly, PPNs were introduced as trees with additional transfer edges. This implies knowledge of the vertical versus horizontal transmissions, and that these networks belong to the class of tree-based networks [Pons2018, Fischer2020] (in fact, they are LGT networks, see next section). In [ptn], we introduced perfect transfer networks (PTNs), a specialization of PPNs where characters satisfy the same assumptions as perfect phylogenies: they are binary, have a unique origin, and are assumed to always be inherited vertically once acquired. An important difference of PTNs is that one cannot choose which subtree of the network can be used to explain a character, as the tree of vertical inheritance is fixed. Because of this, as we showed in [ptn], even PPNs on binary characters are different from PTNs. In the latter, a character may or may not be transferred horizontally to other species. This can be useful for transferrable characters that are difficult to revert, such as material acquired horizontally from mitochondria or chloroplasts resulting from endosymbiotic events [Zachar2020, anselmetti2021gene], or in the case of metabolites when HGT plays a role in the generation of new metabolic pathways in bacteria [Goyal2022]. From an algorithmic standpoint, an advantage of the more restricted PTN model is that the problems of deciding whether a network explains a set of characters, or whether a tree can be augmented with transfers to do so, become polynomial-time solvable.
Thus, PTNs are a promising model for HGT inference from characters. However, PTNs have demonstrated a tendency to introduce an excessive number of transfer events. In [ptn], we provide an algorithm that shows that any tree can be made to explain any set of characters by adding transfers, although the resulting networks may be overly complicated and bloated with HGT edges. This raises the question of explaining a set of characters with the simplest possible network. One way would be to build a network or augment a tree with a minimum number of transfers, which is an unsolved complexity problem. Another direction is to impose structural conditions on the desired network.
Our contribution. In this work, we focus on the second direction. We explore the evolutionary structure that many consider as the simplest beyond trees, namely galled trees. These consist of networks in which all underlying cycles are independent, and were first used in the context of hybridization and recombination [gusfield2004optimal]. When they fit the data, galled trees are desirable because of their parsimonious nature and ease of interpretation. They are also a popular graph-theoretical structure that serves as a first-step towards the development of more structurally complex networks [Huson2009, Cardona2020]. In our case, characters that can be explained by a galled tree can be thought of “not quite tree-like, but almost”.
We present galled perfect transfer networks (galled PTNs) which are galled tree-based networks with unidirectional edges that explain a set of characters. We then provide polynomial-time algorithms for two problems. In the galled-completion problem, we ask whether it is possible to complete a given tree with transfers to obtain a galled PTN. We also study the galled-compatibility problem in which, given a set of characters, we must decide whether a galled PTN can explain them. During the process, we provide several structural characterizations of characters that can be displayed on a given tree or a network. To improve the reading experience, we provide a sketch for most proofs, and refer to the appendix for detailed proofs.
We also propose a case study of perfect transfer networks on a real dataset consisting of bacterial species and functional characters obtained the Kyoto Encyclopedia of Genes and Genomes (KEGG) database [KEGG]. We show that the conditions required to explain a set of characters with a galled tree are difficult to achieve, as most characters prevent such an explanation, even when taken individually. On the other hand, the case study lets us see that character losses and unresolved species trees can be responsible for this phenomenon. This leads to future questions such as how to model losses and transfers together, and how character-based transfer prediction can be used to resolve trees.
Related work. Aside from PPNs, other models have been proposed to explain characters via networks. In recombination networks [gusfield2014recombinatorics], characters are explained by an ancestral recombination graph (ARG) in which hybridation nodes represent recombination events through crossovers. A fundamental difference is that such hybrids do not consider a donor/recipient relationship whereas in HGTs, it can be important to distinguish between the parental and lateral acquisition. As we showed in [ptn, Figure 3], PTNs and recombination graphs explain different sets of characters, even on galled trees with a single transfer/hybridation event, the main reason being that crossover events are different from transfer events. Nonetheless, it is worth mentioning that in [gusfield2004optimal], the author shows how to reconstruct a galled ancestral recombination graph from a set of $m$ characters and $n$ taxa, if possible, in time $O(nm+n^{3})$. Using this approach, the aforementioned work of Warnow et al. [warnow2024statistically] shows how a galled tree can be reconstructed (or not) from a set of characters in the PPN model, where characters are interpreted as bipartitions of the trees contained in the network, achieving the same time complexity. In Section 2, we show how this formulation differs from ours. In a similar vein, in [vanIersel2010, kelk2012] the authors study the question of reconstructing a network that displays a set of characters in the softwired sense, meaning that for each character, some tree contained in the network contains it as a clade (characters are called clusters therein). In particular, one can reconstruct in polynomial time a level-1 network that explains a set of characters, if one exists. As level-1 networks are closely related to galled trees, it is likely that the latter can be used in our setting. But, as also argued in [ptn, Figure 2], softwired characters can explain different sets of characters than PTNs. The main reason is that distinguishing between horizontal and vertical edges does not allow choosing which tree of the network should be used to explain a character.
2 Preliminaries
A phylogenetic network, or simply a network for short, is a directed acyclic graph $N=(V,E)$ with one node $\rho(N)$ of in-degree zero, called the root. A node of in-degree one and outdegree zero is a leaf, and $L(N)$ denotes the set of leaves of $N$. A node of in-degree and out-degree $1$ is a subdivision node, which we allow. For $W\subseteq V$, $N-W$ is the directed graph obtained after removing $W$ and incident edges. An underlying cycle of $N$ is a set of nodes and edges that form a cycle when ignoring the edge directions. A network $N$ is a galled tree if no two distinct underlying cycles of $N$ contain a common node (see Figure 1.(c)).
A tree $T$ is a network with no underlying cycle. We write $u\preceq_{T}v$ if $v$ is on the path from $\rho(T)$ to $u$, in which case $v$ is an ancestor of $u$ and $u$ a descendant of $v$. Note that $v$ is an ancestor and descendant of itself. Two nodes $u$,$v$ are incomparable in $T$ if none descends from the other, i.e. if neither $u\preceq v$ nor $v\preceq u$. The parent of a node $v$ in $T$ is $p_{T}(v)$. For $v\in V(T)$, we use $T(v)$ for the subtree of $T$ rooted at $v$, that is, $T(v)$ contains $v$ and all of its descendants. We may write $L_{T}(v)$ as a shorthand for $L(T(v))$, or just $L(v)$ if $T$ is understood. The set $L(v)$ is called a clade of $T$.
An LGT network (where LGT comes from Lateral, or Horizontal, Gene Transfers[Cardona_2015] is a network $N=(V,E_{S}\cup E_{T})$, where $\{E_{S},E_{T}\}$ is a specified partition of the edge-set of $N$, such that the subgraph $\mathcal{T}_{N}:=(V,E_{S})$ is a tree with the same set of nodes as $N$. The tree $\mathcal{T}_{N}$ is called the support tree of $N$. The edges in $E_{S}$ are called support edges and the edges in $E_{T}$ are called transfer edges. A node that is the endpoint of a transfer edge is called a transfer node. We assume that for each transfer edge $(u,v)\in E_{T}$, the nodes $u$ and $v$ are incomparable in $\mathcal{T}_{N}$. We also assume that transfer nodes have exactly one out-neighbor in $\mathcal{T}_{N}$. The tree obtained from $\mathcal{T}_{N}$ by suppressing its subdivision nodes is called the base tree of $G$ [^1] Suppressing a subdivision node $u$ with parent $p$ and child $v$ consists of removing $u$ and adding an edge from $p$ to $v$. For $v\in V(N)\setminus\{\rho(N)\}$, we use $p_{N}(v)$ to denote the unique in-neighbor of $v$ in the support tree of $N$. For simplicity, an LGT network that is also a galled tree will be called a galled LGT network.
2.1 Perfect transfer networks
Let $\mathcal{S}$ be a set of taxa. A character $C$ is a subset of $\mathcal{S}$, which represents the set of taxa that possess the common trait. We usually denote a set of characters by $\mathcal{C}$. To formalize PTNs, given an LGT network $N$, a $\mathcal{C}$-$labeling$ of $N$ is a function $l:V(N)\to 2^{\mathcal{C}}$ that maps each node to the subset of characters it possesses.
Definition 1.
Let $\mathcal{S}$ be a set of taxa, let $\mathcal{C}\subseteq 2^{\mathcal{S}}$ be a set of characters, and let $N=(V,E_{S}\cup E_{T})$ be an LGT network with leafset $\mathcal{S}$. We say that a $\mathcal{C}$-labeling $l$ of $N$ explains $\mathcal{C}$ if the following conditions hold:
  1. 1.
    for each leaf $x$, $l(x)=\{C\in\mathcal{C}:x\in C\}$ (leaves are labeled by their characters);
  2. 2.
    for each support edge $(u,v)\in E_{S}$, $C\in l(u)$ implies that $C\in l(v)$ (never lost once acquired);
  3. 3.
    for each $C\in\mathcal{C}$, there exists a unique node $v\in V$ that reaches every node $w$ satisfying $C\in l(w)$ (single origin).
Furthermore, we call $N$ a perfect transfer network (PTN) for $\mathcal{C}$ if there exists a $\mathcal{C}$-labeling of $N$ that explains $\mathcal{C}$.
Figure 1: (a) A tree and characters $C_{1},C_{2},C_{3}$. (b) A PTN with $T$ as base tree that is not a galled tree (two underlying cycles contain the left child of the root). (c) A galled-completion of $T$. Note that in (b) and (c) for every character $C$, there exists a node (in gray) that transmits to all its vertical descendants, and reaches every leaf in $C$.
Figure 1 shows two PTNs for the same set of characters. It does not exhibit the full $\mathcal{C}$-labeling, but a possible origin of each character is annotated on the internal nodes. In [ptn], it is shown that any set of characters can be explained by some PTN. In fact, any tree can become a PTN by adding enough transfer edges. Our goal is to constrain the PTNs to avoid overcomplicated solutions.
Let $T$ be a tree on leafset $\mathcal{S}$, and let $\mathcal{C}$ be a set of characters. We say that $T$ is galled-completable for $\mathcal{C}$ if there exists a galled LGT network $N$ that explains $\mathcal{S}$ and whose base tree is $T$. We call such a galled PTN a galled-completion of $T$. An example of this problem is shown on Figure 1. For a set of characters $\mathcal{C}$ on taxa $\mathcal{S}$, we say that $\mathcal{C}$ is galled-compatible if there exists a galled tree on leafset $\mathcal{S}$ that explains $\mathcal{C}$. Our problems of interest are the following:
  • The Galled Completion problem: given a tree $T$ on leafset $\mathcal{S}$ and a character set $\mathcal{C}$, is $T$ galled-completable for $\mathcal{C}$?
  • The Galled Compatibility problem. Given a set of taxa $\mathcal{S}$ and a character set $\mathcal{C}$, is $\mathcal{C}$ galled-compatible?
2.2 Properties of galled PTNs
We begin by stating properties of galled PTNs that will be useful throughout. Let $N$ be a network that explains a set of characters $\mathcal{C}$. Observe that if a leaf $l$ does not possess a character $C\in\mathcal{C}$, then by the “never lost once acquired” condition, the parent of $l$ in $\mathcal{T}_{N}$ cannot possess $C$ either (otherwise, the parent would be required to transfer $C$ to $v$). In fact, no ancestor of $l$ in $\mathcal{T}_{N}$ can possess $C$. This lets us deduce a subset $F_{C}$ of nodes that forbid $C$, as follows:
$F_{C}(N)=\{v\in V(N):\exists l\in L_{\mathcal{T}_{N}}(v)\mbox{ such that $l\notin C$}\}.$
It follows that the origin of $C$ must be a node of $N-F_{C}$. By the “single origin” condition, it is also necessary that this origin reaches every leaf in $C$. As shown in [ptn], the existence of such a node is also sufficient to explain $C$. Since it is easier to deal with, we will heavily use the following characterization [^2] Note that we adapted this characterization, since in the original definition of PTNs, the taxa were treated as sets of characters instead of the other way around..
Lemma 1 ([ptn]).
Let $N$ be an LGT network and let $\mathcal{C}$ be a set of characters. Then $N$ is a perfect transfer network for $\mathcal{C}$ if and only if for every character $C\in\mathcal{C}$, the network $N-F_{C}(N)$ contains a node $v$ that reaches every leaf in $C$.
{proofsketch}
If $N$ is a PTN for $\mathcal{C}$, then for a character $C\in\mathcal{C}$, as mentioned the nodes in $F_{C}(N)$ cannot possess $C$. This means that the origin of $C$, that is, a node that satisfies the third condition of PTNs, must be in $N-F_{C}(N)$. Conversely, if some node of $N-F_{C}(N)$ reaches every leaf, then it can serve as the origin of the character. ∎
We also describe two useful generic properties of galled LGT networks. The first one says that a node $v$ cannot have two descending transfer nodes that each go outside of the $v$ subtree (otherwise, $v$ would be part of two distinct cycles).
{lemmarep}
Let $N=(V,E_{S}\cup E_{T})$ be a galled LGT network and let $v\in V(N)$ be a node with two distinct descendants $x$ and $y$ that are transfer nodes (with $v\in\{x,y\}$ being possible). Then one of the transfer edges incident to $x$ or $y$ has both endpoints in the subtree $\mathcal{T}_{N}(v)$.
{proofsketch}
If $v$ has two descendants $x,y$ with transfer neighbors that are external to the $v$ subtree of $\mathcal{T}_{N}$, then the two outgoing transfers inevitably create two distinct cycles that both contain $v$, contradicting the galled property. ∎
Proof.
We prove this by contraposition. Let $x^{\prime}$ and $y^{\prime}$ be nodes that share a transfer edge with $x$ and $y$, respectively. Suppose that neither $x^{\prime}$ nor $y^{\prime}$ are contained in $\mathcal{T}_{N}(v)$. Write $lca(a,b)$ for the lowest common ancestor of $a,b$ in $\mathcal{T}_{N}$. By our supposition, in $\mathcal{T}_{N}$ we have that $v$ is on the path from $lca(x,x^{\prime})$ to $x$, and from $lca(y,y^{\prime})$ to $y$. This implies the existence of two underlying cycles in $N$: there is a cycle formed by $lca(x,x^{\prime})-x^{\prime}-x-lca(x,x^{\prime})$ and the cycle formed by $lca(y,y^{\prime})-y^{\prime}-y-lca(y,y^{\prime})$ (where the edges between and $x^{\prime},x$, and between $y^{\prime},y$ are used, and the rest use edges of $\mathcal{T}_{N}$). These cycles are distinct since they use a different (unique) transfer edge. However, they intersect at $v$, and thus $N$ is not galled. ∎∎
The next property states that if some $x$ is able to reach a node $y$ that does not descend from it, then $x$ must of course achieve this through a transfer edge.
{lemmarep}
Let $N=(V,E_{S}\cup E_{T})$ be a galled LGT network. Let $x,y\in V$ be two nodes that are not comparable in $\mathcal{T}_{N}$ and such that $x$ reaches $y$ in $N$. Then there exists an edge $(x^{\prime},y^{\prime})\in E_{T}$ such that $x^{\prime}\preceq_{\mathcal{T}_{N}}x$ and $y\preceq_{\mathcal{T}_{N}}y^{\prime}$.
{proofsketch}
Consider the first transfer edge $(x^{\prime},y^{\prime})$ on a path from $x$ to $y$ in $N$. Note that $x^{\prime}$ must descend from $x$ in $\mathcal{T}_{N}$. If $y$ is not a descendant of $y^{\prime}$, then from $y^{\prime}$ the path needs to borrow another transfer edge that goes out of the $y^{\prime}$ subtree of $\mathcal{T}_{N}$. Thus, the $y^{\prime}$ subtree has two transfers with external endpoints, contradicting Lemma 2.2. ∎
Proof.
Suppose for a contradiction that such $(x^{\prime},y^{\prime})$ does not exist in $N$. Since $x$ reaches $y$ in $N$, we know that there must exist at least another node $w\in V(N)$ incomparable to both $x$ and $y$ that serves as an intermediary node from $x$ to $y$. More specifically, there exists a transfer edge with an endpoint in $\mathcal{T}_{N}(x)$ and whose receiving end is $w$. Nevertheless, $y$ is not a descendant of $w$ (otherwise, $w=y^{\prime}$ would end the proof), which implies the existence of a transfer edge from some descendant of $w$ to a node incomparable with $w$. However this creates a situation where $\mathcal{T}_{N}(w)$ is a subtree with two transfer edges that have their endpoints outside of the subtree, which is forbidden by Lemma 2.2 ∎∎
2.3 Differences with reconstructions from bipartitions
As we pointed out in the introduction, Warnow et al. [warnow2024statistically] also independently developed galled tree reconstructions algorithms from character data (which are assumed to be SNPs in the paper). Therein, for each character, the species can be in one of two states, and the states are assumed to correspond to a bipartition of a galled tree $N$, where a bipartition is a partition of the leaves that can be obtained by taking a tree displayed by $N$, removing an edge, and taking the two sets of leaves of the resulting connected components (here, “displayed” means a tree that can be obtained by removing one parent edge from each node of in-degree two). They discuss algorithms to reconstruct a galled tree that contains a set of given characters, that is, that contains all the corresponding bipartitions. More specifically, they show that if all the bipartitions of a galled tree $N$ are known (but $N$ is unknown), then $N$ can be reconstructed in polynomial time. They also show that this approach is statistically consistent if characters evolve on a galled tree and may or may not be transmitted on edges labeled as transfer.
We refer to this as the Perfect Phylogenetic Network (PPN) model, as the notion of explaining a character is the same as in [nakhlehtesis]. If we interpret one state as “presence” and the other as “absence”, this is very similar to our formulation, with two notable exceptions. First, we assume an ordering in states from “absence” to “presence”, whereas no such ordering is imposed in the PPN model. More importantly, the notion of displaying a tree does not distinguish between transfer and vertical descent edges. Figure 2 illustrates this difference. Suppose that a given set of characters (i.e., bipartitions) results in the network shown on the left, and consider the character $C$ denoted by the colored square. In the model of [warnow2024statistically], this character is considered as explained, since the displayed tree on the right shows that the character indeed corresponds to a bipartition. On the other hand, this network does not explain the character under the PTN model (see caption). Roughly speaking, displaying a tree allows removing a vertical edge, whereas we would require transmitting characters along that removed vertical edge. See the figure caption for details. We do note that [warnow2024statistically] consider an evolutionary model in which characters evolve down a tree in which transfer edges are explicitly labeled. However, this is only to describe the statistically consistent model, as unlike us the algorithms do not consider this labeling. Let us also mention that [warnow2024statistically] do not consider the tree completion problem.
Figure 2: (a) An example that shows that galled PPNs are not necessarily galled PTNs. The colored square represents a character present at the two rightmost leaves (or in the PPN formulation, there are two states: with or without a square). (b) A tree displayed by the network, obtained by deleting one edge (we left the resulting nodes with a single child, although they could be suppressed). The illustrated character labeling shows that this network explains the character under the PPN model. As for PTNs, one can see from the definition of $F_{C}(N)$ that under the PTN formulation, no internal node of the initial network can contain the character, and thus that the network cannot explain it.
3 The Galled Completion Problem
In this section, let $T$ be a tree on taxa $\mathcal{S}$ and let $\mathcal{C}$ be a set of characters. We assume that $T$ has no subdivision node. We will first describe the necessary conditions for $T$ to be galled-completable. The key factor lies in the ancestor relationship that exists between any set of first-appearance (FA) nodes for distinct characters.
Definition 2.
Let $T$ be a tree on leafset $\mathcal{S}$ and let $C\subseteq\mathcal{S}$ be a character. A node $v\in V(T)$ is a first-appearance (FA) node for $C$ if $L_{T}(v)\subseteq C$ and $v$ is either the root, or its parent $u$ satisfies $L_{T}(u)\not\subseteq C$. The set of FA nodes for $C$ in $T$ is denoted as $\alpha_{T}(C)$.
In other words, $v$ is a FA for $C$ if it roots a maximal subtree of taxa that contain $C$. An example of this definition is shown in Figure 3.
Figure 3: A tree $T$ on species $\mathcal{S}=\{s_{1},s_{2},s_{3},s_{4},s_{5},s_{6},s_{7}\}$ with character set $\mathcal{C}=\{C_{1},C_{2},C_{3},C_{4}\}$. Every colored shape indicates to which character a specific taxa belongs to. The corresponding sets of FAs for every character are as follows: $\alpha_{T}(C_{1})=\{x,y\},\alpha_{T}(C_{2})=\{y,s_{2}\},\alpha_{T}(C_{3})=\{v\},\alpha_{T}(C_{4})=\{u\}$.
For a character $C\in\mathcal{C}$, we will say that a node $v$ is an origin for $C$ if $v$ reaches every leaf in $C$ in $N-F_{C}(N)$. By Lemma 1, our goal is to ensure that every character has an origin. We first show that in any galled-completion of a tree $T$, an origin of a character $C$ must descend from a FA node (or it could be a transfer node added just above it), and it must “give” its character to all other FAs by transferring just above them.
{lemmarep}
Let $N$ be a galled-completion of a tree $T$ that explains $\mathcal{C}$. Let $C\in\mathcal{C}$ be a character and let $w$ be an origin for $C$ in $N$. Then there is $\alpha_{i}\in\alpha_{T}(C)$ such that both of the following hold:
  • either $w\preceq\alpha_{i}$ or $w$ is a transfer node whose child in $\mathcal{T}_{N}$ is $\alpha_{i}$.
  • for every $\alpha_{j}\in\alpha_{T}(C)\setminus\{\alpha_{i}\}$ there is a transfer edge $(u,v)$ such that $u\preceq w$ and $v=p_{N}(\alpha_{j})$.
{proofsketch}
By definition, in $T$ the ancestors of FAs are in $F_{C}(N)$ since they have descendants not in $C$. Thus, an origin $w$ in $N$ must be in a FA subtree, or it could be a transfer node added above when adding transfer edges from $T$ to $N$. If there are multiple FAs, $w$ must be able to reach the other ones, and Lemma 2.2 lets us establish that transfer edges are needed to achieve this. ∎
Proof.
We focus on the first condition. First observe that if $v\in V(T)$, then $v\in V(N)$ as well. Moreover, the parent $p_{v}$ of $v$ in $T$ is either also the parent of $v$ in $\mathcal{T}_{N}$, or the latter is a transfer node created to complete $T$ into $N$. In the latter case, note that at most one transfer node could could be present on the path between $p_{v}$ and $v$ in $\mathcal{T}_{N}$, as otherwise, there would be two consecutive transfer nodes which, as one can check, would violate Lemma 2.2. In other words, some edges of $T$ are subdivided to obtain $\mathcal{T}_{N}$, but each edge can be subdivided at most once.
Let us next observe that if $w$ is an origin for $C$, by definition it must be a node in $N-F_{C}(N)$, which is a graph that only contains FA nodes of $T$, their support tree descendants in $N$, and possibly transfer nodes that are the parents of FA nodes. Thus either $w$ descends from some $\alpha_{i}\in\alpha_{T}(C)$, or results from a subdivision of $(p_{T}(\alpha_{i}),\alpha_{i})$. In the latter case, $w$ is the parent of $\alpha_{i}$ since an edge can only be subdivided once, as explained above. Therefore, the first condition of our statement holds.
For the second condition, as we know the origin $w$ descends from some $\alpha_{i}$ or is its parent. Let $\alpha_{j}$ be a FA other than $\alpha_{i}$. Let $l$ be a leaf descending from $\alpha_{j}$ in $\mathcal{T}_{N}$. Since $w$ is an origin, it reaches $l$ in $N-F_{C}(N)$ and thus also in $N$. As $w$ is not an ancestor of $l$, by Lemma 2.2 there is a transfer edge $(u,v)$ in $N$ such that $u\preceq w$ and $l\preceq v$ in $\mathcal{T}_{N}$. The endpoints of this transfer edge must also be in $N-F_{C}(N)$ for $w$ to be able to use it. Next, note that $p_{T}(\alpha_{j})$ has descendants not in $C$. Thus in $N$, $p_{T}(\alpha_{j})$ and its support tree ancestors are in $F_{C}(N)$, and so it follows that $v$ descends from $\alpha_{j}$ or is a transfer node on its parent edge, in which case $v=p_{N}(\alpha_{j})$ (again since edges can only be subdivided once). If $\alpha_{j}$ is a leaf, the latter must hold and we are done. So assume that $\alpha_{j}$ is not a leaf and let $l_{1},l_{2}$ be two leaves descending from $\alpha_{j}$ in $\mathcal{T}_{N}$. Then there are two transfer edges $(u_{1},v_{1})$ and $(u_{2},v_{2})$ that exist in $N-F_{C}(N)$, where $u_{1},u_{2}$ descend from $w$ and $v_{1},v_{2}$ are ancestors of $l_{1}$ and $l_{2}$, respectively. Since these two transfers go out of the $\mathcal{T}_{N}(w)$ subtree, they must in fact be equal by Lemma 2.2. In other words, $w$ reaches every leaf below $\alpha_{j}$ through a single transfer $(u,v)$ such that $v$ is an ancestor of all those leaves. Since $v$ is in $N-F_{C}(N)$, it follows that $v$ must be the parent of $\alpha_{j}$ (and not $\alpha_{j}$ itself because it cannot be a transfer node as its has at least two children, by assumption on $T$). ∎∎
We next argue that the galled requirement places an important limitation on FAs, as there can be at most two per character. This means that a character $C$ must either be a clade of $T$, or it could be split in two clades.
{lemmarep}
Let $T$ be a galled-completable tree for a character set $\mathcal{C}$. Then $\alpha_{T}(C)\leq 2$ for any character $C\in\mathcal{C}$.
{proofsketch}
By Lemma 3, an origin for $C$ must be in one of the FA subtrees (or just above). If there are three FAs, that origin must also reach the other two FAs, but this requires two descending transfers that contradict Lemma 2.2. ∎
Proof.
Let $N$ be a galled-completion of $T$ that explains $\mathcal{C}$. Suppose for contradiction that $\alpha_{T}(C)\geq 3$ for some $C\in\mathcal{C}$. Let $w$ be an origin for $C$ in $N-F_{C}(N)$. By Lemma 3, $w$ descends from or is the parent of one of the FAs, say $x\in\alpha_{T}(C)$. Let $y,z$ be two other FAs of $C$. Note that, by definition of FAs, $x,y,z$ are pairwise incomparable in $T$. By the second statement of Lemma 3, there are transfer edges $(u_{1},p_{N}(y))$ and $(u_{2},p_{N}(z))$ where $u_{1}$ and $u_{2}$ descend from $w$ in $\mathcal{T}_{N}$. This implies that the subtree $\mathcal{T}_{N}(w)$ contains two outgoing transfers, which contradicts Lemma 2.2. Thus $|\alpha_{T}(C)|\leq 2$ for all $C\in\mathcal{C}$. ∎∎
We next show that FAs of distinct characters have limited ancestry relationships. The proof relies on a case analysis and the previous properties.
{lemmarep}
Let $T$ be a galled-completable tree for $\mathcal{C}$. Suppose that there exists two distinct characters $A$ and $B$ with $\alpha_{T}(A)=\{a_{1},a_{2}\}$ and $\alpha_{T}(B)=\{b_{1},b_{2}\}$. Then the two following statements hold:
  1. 1.
    If $b_{1}\prec a_{1}$ and $b_{2}$ is not comparable to $a_{1}$, then $a_{2}=b_{2}$.
  2. 2.
    If $a_{1}=b_{1}$, then either $b_{2}\prec a_{2}$ or $a_{2}\prec b_{2}$.
{proofsketch}
For the first statement, if $b_{1}\prec a_{1}$ and $b_{2}$ is outside of the $a_{1}$ subtree, then $b_{1}$ needs to use the same transfer edge as $a_{1}$ to exchange material (because of the galled property). This is only possible if they send the character to the same clade, that is, if $a_{2}=b_{2}$.
For the second statement, if $a_{1}=b_{1}$, then again the two characters must use the same transfer edge to exchange material. In fact, one can argue that $a_{1}=b_{1}$ must be the receiving end of the transfer, and that $a_{2},b_{2}$ must be comparable to be able to use the same node to send the character. ∎
Proof.
We begin with the first statement. Suppose for a contradiction that the conditions of the lemma hold, but that $a_{2}\neq b_{2}$. Let $N$ be a galled-completion of $T$ that explains $\mathcal{C}$. This leads to the following two cases.
  • Case 1: $a_{2}$ and $b_{2}$ are comparable. Suppose $b_{2}\prec a_{2}$. Let $w$ be an origin for $A$ in $N-F_{A}(N)$. By Lemma 3, $w$ descends from $a_{1}$ or $a_{2}$, or is a transfer node with one of those as a child. Suppose first that $w\preceq a_{1}$ or $w=p_{N}(a_{1})$. We know by Lemma 3 that there exists a transfer edge from some descendant of $w$ that has $p_{N}(a_{2})$ as endpoint. Also by Lemma 3, $p_{N}(b_{2})$ or a descendant of $b_{2}$ is a transfer node whose other endpoint is $p_{N}(b_{1})$ or a descendant of $b_{1}$. This implies that there are two transfer nodes in the subtree $\mathcal{T}_{N}(p_{N}(a_{2}))$ whose other endpoint is outside, which is forbidden by Lemma 2.2. The case where $w\preceq a_{2}$ or is the parent of $a_{2}$ is symmetric.
    Next suppose that $a_{2}\prec b_{2}$. By Lemma 3, one of $p_{N}(a_{1})$ or $p_{N}(a_{2})$ is the receiving end of a transfer, whose sending end is on the other side. If $p_{N}(a_{1})$ is a transfer node, we also know that $p_{N}(b_{1})$ or one of its descendants is a transfer node, with the other endpoint on the $b_{2}$ side. Either way, $\mathcal{T}_{N}(p_{N}(a_{1}))$ has two descending transfer nodes with external endpoints, contradicting Lemma 3. Thus $p_{N}(a_{2})$ is the receiving end of a transfer. By a symmetric argument, $p_{N}(b_{1})$ is the receiving end of a transfer. This implies that there are transfers between $\mathcal{T}_{N}(a_{1})$ and $\mathcal{T}_{N}(b_{2})$ (or their parents) going in opposite directions. Because transfer edges are unidirectional, these transfers are distinct, from which it can easily be seen that there are two underlying cycles intersecting. Therefore, this case is not possible.
  • Case 2: $a_{2}$ and $b_{2}$ are incomparable. By Lemma 3, there is one transfer with one endpoint being $p_{N}(a_{1})$ or a descendant of $a_{1}$, and the other being $p_{N}(a_{2})$ or a descendant of $a_{2}$. Likewise, there is a similar transfer connecting $b_{1}$ and $b_{2}$. Since $a_{2}$ and $b_{2}$ are incomparable, these two transfers must be distinct, they are both in $\mathcal{T}_{N}(a_{1})$ or $\mathcal{T}_{N}(p_{N}(a_{1}))$, and they both have external endpoints, contradicting Lemma 2.2.
Since neither case is possible, we must have $a_{2}=b_{2}$.
Next, consider the second statement. Let $N$ be a galled-completion of $T$ that explains $\mathcal{C}$. Suppose for a contradiction that $a_{1}=b_{1}$, but neither $a_{2}\prec b_{2}$ nor $b_{2}\prec a_{2}$ holds i.e. $a_{2}$ and $b_{2}$ are incomparable in $\mathcal{T}_{N}$. By Lemma 3, there is a transfer edge between descendants of $a_{1}$ and $a_{2}$ (or their parents), and a transfer edge between descendants of $b_{1}$ and $b_{2}$ (or their parents). Since $a_{2}$ and $b_{2}$ are incomparable, these transfers must be distinct. However, they imply the existence of two transfer nodes descending from $a_{1}=b_{1}$ (or its parent) that have an external endpoint, contradicting Lemma 2.2. ∎∎
Note that the first statement of the proof of Lemma 3 requires transfer edges to be unidirectional. If we allowed bidirectional transfer edges, then we can devise examples in which the lemma does not hold (not shown due to space constraints). Allowing such edges already leads to more complex structures.
We can begin describing our algorithmic strategy. The first step is to locate and count the FAs for each character to verify the condition established by Lemma 3. A subtree which is rooted at a FA for a specific character $C$ is in fact a maximal clade for $C$. An intuitive way of joining these clades is to add transfer edges between the different subtrees to fulfill the connectivity requirement in Lemma 1. However, this may add superfluous transfer edges, as only the minimal ones are required.
To make this precise, let $v\in V(T)$. We say that another node $w\in V(T)$ is an FA neighbor of $v$ if there exists a character $C\in\mathcal{C}$ such that $\alpha_{T}(C)=\{v,w\}$. We also say that $w$ is a minimal FA neighbor of $v$ if $w\preceq_{T}w^{\prime}$ for every FA neighbor of $v$. A pair of FA nodes $\{v,w\}$ is called simple if $w$ is the unique FA neighbor of $v$ and $v$ is the unique FA neighbor of $w$.
Definition 3.
Let $T$ be a tree on leafset $\mathcal{S}$ with character set $\mathcal{C}$. Let $T^{\prime}$ be the tree obtained by subdividing every edge of $T$ once. Then the redundancy-free network for $T$ is the LGT-network $N=(V,E_{S}\cup E_{T})$ with $T^{\prime}$ as support tree obtained as follows:
  • for each $v\in V(T)$ with at least two FA neighbors, and for every minimal FA neighbor $w$ of $v$, add the transfer edge $(p_{T^{\prime}}(w),p_{T^{\prime}}(v))$;
  • for each pair $\{v,w\}$ of simple FA nodes, add one of the transfer edges $(p_{T^{\prime}}(v),p_{T^{\prime}}(w))$ or $(p_{T^{\prime}}(w),p_{T^{\prime}}(v))$ arbitrarily (but not both).
Figure 4: (a) A given tree $T$ on four characters. (b) The FA neighbors between every FAs for each character are indicated by the dashed lines. (c) The redundancy free network of $T$. Note that $C_{3}$ has simple FAs, so the direction for the transfer edge is arbitrary. On the other hand, for $C_{1}$ and $C_{2}$ note that $z$ is the minimal FA neighbor of $y$ so the direction of the transfer edge is not arbitrary.
The idea of this network is that if a node $y$ has multiple FA neighbors, as in Figure 4, then there are many characters with $y$ as a FA. By adding a transfer edge whose sender is the parent of the minimal FA neighbor $z$ of $y$, again as in the figure, all those characters can have their origin on the $z$ side and transfer to $y$. In the figure, we add a transfer from the parent of $z$ to that of $y$ for character $C_{2}$. At the same time, $C_{1}$ can use that same transfer edge because it also wants $y$ and its descendants to inherit the character. For characters whose FAs correspond to a simple pair, no such transfer re-use is needed and the direction is arbitrary.
Before proving our main characterization in terms of redundancy-free networks, we need an intermediary result.
{lemmarep}
Let $T$ be a tree on leafset $\mathcal{S}$ and character set $\mathcal{C}$. If $T$ is galled-completable, then in the redundancy-free network $N=(V_{T},E_{T})$ of $T$, every node is incident to at most one transfer edge.
{proofsketch}
Suppose that some node $u^{\prime}$ of $N$ is incident to two distinct transfer edges, with $v^{\prime}$ and $w^{\prime}$ as the other endpoints (where $u^{\prime},v^{\prime},w^{\prime}$ are subdivision nodes added when transforming $T$ to $N$). Since $N$ only adds transfers between minimal FA neighbors, this means that $v^{\prime}$ and $w^{\prime}$ are incomparable nodes of $\mathcal{T}_{N}$ (a special case arises when $v^{\prime}=w^{\prime}$, see the full proof for details). Roughly speaking, this means that two characters need to exchange material between the clades of $u^{\prime}$ and $v^{\prime}$, and also between the clades of $u^{\prime}$ and $w^{\prime}$. This is only possible in a galled-completion if the $u^{\prime}$ subtree contains at least two transfers, one exchanging with $v^{\prime}$ and the other with $w^{\prime}$. This creates two transfers below $u^{\prime}$ with external endpoints, contradicting the galled property by Lemma 2.2. ∎
Proof.
Suppose for contradiction that some node $u^{\prime}$ of $N$ is incident to two distinct transfer edges, with $v^{\prime},w^{\prime}$ the other endpoints of those edges (note that $v^{\prime}=w^{\prime}$ is possible). By construction, $u^{\prime},v^{\prime},w^{\prime}$ are subdivision vertices of $\mathcal{T}_{N}$, and are respective parents of FA nodes $u,v,w$ in $T$. Moreover, $v$ and $w$ must be FA neighbors of $u$, and so there are characters $C_{uv}$ (resp. $C_{uw}$) with FAs $\{u,v\}$ (resp. $\{u,w\}$) in $T$.
Suppose first that $v\neq w$. In that case, $C_{uv}$ and $C_{vw}$ are distinct as their FAs differ. By the second part of Lemma 3, the two characters $C_{uv}$ and $C_{uw}$ have a common FA $u$, and thus $v$ and $w$ are comparable. Suppose without loss of generality that $w\prec_{T}v$. We can further assume that $w$ is a minimal FA neighbor of $u$ (otherwise, we choose $w$ to be such a neighbor, which does not affect the presence of the edge between $u^{\prime}$ and $v^{\prime}$), which guarantees that $(w^{\prime},u^{\prime})\in E_{T}$. By the construction of $N$, the edge $(v^{\prime},u^{\prime})$ cannot be in $E_{T}$, as $u^{\prime}$ only receives transfers from its minimal neighbors. Therefore, the transfer between $u^{\prime}$ and $v^{\prime}$ is in the $(u^{\prime},v^{\prime})$ direction. As $v$ is not in a simple pair, this edge is present because $v$ has multiple FA neighbors and $u$ is minimal. This means that there is another character $C_{vz}$ with FAs $\{v,z\}$, where $z\neq u$. Using Lemma 3 in the same way as before, we get $u\prec_{T}z$. We now have two pairs of FAs $\{z,v\}$ and $\{u,w\}$, where $u\prec_{T}z$, $w$ is not comparable to $z$ (because $v$ is not), and $w\neq v$. By putting $\{z,v\}=\{a_{1},a_{2}\}$ and $\{u,w\}=\{b_{1},b_{2}\}$, we obtain a contradiction of the first part of Lemma 3.
So suppose that $v=w$, and hence $C_{uv}=C_{uw}$. This means that both edges $(u^{\prime},v^{\prime})$ and $(v^{\prime},u^{\prime})$ are present. This is not possible if $\{u,v\}$ is simple, so $u,v$ both have at least two FA neighbors, $v$ must be a minimal FA neighbor of $u$ and $u$ a minimal FA neighbor of $v$. Let $x\neq v$ be another FA neighbor of $u$. Then some character $C_{ux}$ has $\{u,x\}$ as FAs. Since $C_{uv}$ has FAs $\{u,v\}$, by the second part of Lemma 3, $x$ and $v$ must be comparable. Since $v$ is minimal, $v\prec_{T}x$. As $v$ also has other FA neighbors, by a symmetric argument, some character $C_{vy}$ has FAs $\{v,y\}$ with $u\prec_{T}y$. But the pair of FAs $\{y,v\},\{u,x\}$ has $u\prec_{T}y$, $x$ incomparable to $y$ ($x\prec_{T}y$ is not possible since $v$ would descend from $y$, and $y\prec_{T}x$ is not possible as $u$ would descend from $x$), and $v\neq x$, a contradiction of the first part of Lemma 3. ∎∎
We finally arrive to our characterization of galled-completable trees.
{lemmarep}
A tree $T$ on leafset $\mathcal{S}$ with character set $\mathcal{C}$ is galled-completable if and only if its redundancy-free network $N=(V,E_{S}\cup E_{T})$ is a galled tree.
{proofsketch}
In the forward direction, if $T$ can be completed into some LGT network $N^{\prime}$, we can argue that for every transfer edge $(u,v)$ added to $N$, there is a corresponding transfer edge in $N^{\prime}$ whose endpoints are either $u$ and $v$, or their support tree descendants. Roughly speaking, this means that the cycles of $N$ have corresponding cycles in $N^{\prime}$, and since $N^{\prime}$ is a galled tree, so is $N$. Conversely, if $N$ is galled, it explains $\mathcal{C}$ by construction. That is, for every character split into two clades in $T$, we either added an edge above the subtrees to connect them, or there is an edge that was added in the descendants due to some minimal FA neighbor, which allow meeting the connectivity requirements of Lemma 1. ∎
Proof.
($\Rightarrow$) Suppose that $T$ is galled-completable and let $N^{\prime}=(V^{\prime},E_{S}^{\prime}\cup E_{T}^{\prime})$ be a galled-completion for $T$ that explains $\mathcal{C}$. We need to show that $N$ is a galled tree. We assume that $N^{\prime}$ is obtained from $T$ by subdividing every edge once, then adding a set of transfer edges between subdivision nodes (note that if $N^{\prime}$ exists, such a galled-completion also exists). Since $N$ also subdivides every edge once, this allows us to assume that $N^{\prime}$ and $N$ have the same set of nodes, putting the nodes in correspondence without ambiguity.
We first map each transfer edge $(u^{\prime},v^{\prime})$ of $N$ to a distinct transfer edge $(x,y)$ of $N^{\prime}$, such that $y\in\{u^{\prime},v^{\prime}\}$. Let $(u^{\prime},v^{\prime})\in E_{T}$. Note that by the construction of $N$, $u^{\prime},v^{\prime}$ are subdivision nodes of $\mathcal{T}_{N}$ and are parents of FAs $u,v$. Moreover, $u,v$ are FA neighbors, and thus there is some character $C$ with $\alpha_{T}(C)=\{u,v\}$. By Lemma 3, in $N^{\prime}$ there is a transfer edge $(x,y)$, where either $x\preceq_{\mathcal{T}_{N}}u^{\prime}$ and $y=v^{\prime}$, or $x\preceq_{\mathcal{T}_{N}}v^{\prime}$ and $y=u^{\prime}$. Either way, we map $(u^{\prime},v^{\prime})$ to $(x,y)$. Note that by Lemma 3, $u^{\prime}$ and $v^{\prime}$ are only incident to one transfer edge $(u^{\prime},v^{\prime})$. Therefore, only $(u^{\prime},v^{\prime})$ can be mapped to $(x,y)$, as $y\in\{u^{\prime},v^{\prime}\}$.
Next, consider $(u^{\prime},v^{\prime})\in E_{T}$ and its corresponding edge $(x,y)$ in $N^{\prime}$. Suppose that $(x,y)=(\tilde{u},v^{\prime})$ for some descendant $\tilde{u}$ of $u^{\prime}$ in $\mathcal{T}_{N^{\prime}}$. Consider the network $\tilde{N}$ obtained from $N^{\prime}$ by removing $(\tilde{u},v^{\prime})$ and inserting $(u^{\prime},v^{\prime})$ (which does nothing if $\tilde{u}=u^{\prime}$). Note that in $N^{\prime}$, the transfer edge $(\tilde{u},v^{\prime})$ belongs to a unique underlying cycle $H$ formed by the transfer edge, plus the paths from $\tilde{u}$ and from $v^{\prime}$ to the lowest common ancestor of $\tilde{u}$ and $v^{\prime}$ in $\mathcal{T}_{N^{\prime}}$. Since $u^{\prime}$ is on the path between that ancestor and $\tilde{u}$, in $\tilde{N}$ the incorporation of $(u^{\prime},v^{\prime})$ only creates an underlying cycle whose set of vertices is a subset of $H$. Since $H$ did not intersect with other underlying cycles in $N^{\prime}$, this new underlying cycle does not either, and it follows that $\tilde{N}$ is a galled tree as well. The same idea applies if $y=u^{\prime}$ instead.
Since all underlying cycles of $N^{\prime}$ are independent, we can apply the transformation from the previous paragraph to every mapped transfer edge of $N^{\prime}$. Notice that because each $(u^{\prime},v^{\prime})\in E_{T}$ maps to a distinct transfer edge of $N^{\prime}$, no edge of $N^{\prime}$ is “moved” twice by this process. Therefore, we obtain a galled tree whose set of transfer edges contains $E_{T}$. It then follows that $N$ is a galled tree as well.
()($\Leftarrow$)( ⇐ ) Suppose that the redundancy-free network $N$ is a galled tree. We show that for every character in $C\in\mathcal{C}$ there exists a node in $N-F_{C}(N)$ that reaches every leaf in $C$ as required by Lemma 1. Note that if $|\alpha_{N}(C)|=1$, then $C$ forms a clade in $T$ and thus in $\mathcal{T}_{N}$. If $u$ is the FA node of $C$ in $T$, then $u$ is able to reach every element of $C$ in $N-F_{C}(N)$ since every support tree descendant of $u$ is also in $N-F_{C}(N)$, which is sufficient to explain $C$.
Suppose that $|\alpha_{T}(C)|=2$. Let $u$ and $v$ be the FAs of $C$ in $T$, which are the roots of clades $C_{1}$ and $C_{2}$ such that $C=C_{1}\cup C_{2}$. Denote $p_{u}=p_{\mathcal{T}_{N}}(u)$ and $p_{v}=p_{\mathcal{T}_{N}}(v)$, where $p_{u}$ and $p_{v}$ are transfer nodes since every edge of $T$ was subdivided to produce $N$. Note that by definition, in $\mathcal{T}_{N}$ all the descendants of $p_{u}$ and of $p_{v}$ are in $C$, and thus both of them appear in $N-F_{C}(N)$. Suppose that one of the edges $(p_{u},p_{v})$ or $(p_{v},p_{u})$ is present. In the first case, $p_{u}$ is an origin for $C$ and in the second case, $p_{v}$ is an origin.
So suppose that neither transfer edge is present. Then by construction, either $u$ or $v$ has at least two FA neighbors (otherwise, we would have added one of the two edges between their parent in an arbitrary direction and be in the previous case). Suppose without loss of generality that $u$ has at least two FA neighbors. Then $v$ is not one of its minimal FA neighbors, which means that there is some minimal FA neighbor $w$ of $u$ such that $w\prec_{T}v$. Let $p_{w}=p_{\mathcal{T}_{N}}(w)$ and note that $N$ contains the transfer edge $(p_{w},p_{u})$. Then, $v$ is an origin for $C$, since it is in $N-F_{C}(N)$, it can reach all the elements of $C$ that descend from itself in $\mathcal{T}_{N}$, and all those that descend from $u$ through the transfer edge $(p_{w},p_{u})$.
We have therefore shown that all characters have an origin, which by Lemma 1 implies that $N$ is a PTN for $\mathcal{C}$, as desired. ∎∎
1 function FindGalledCompletion($T$,$\mathcal{S}$,$\mathcal{C}$)
2       //$T$ is a tree on taxa set $\mathcal{S}$, $\mathcal{C}$ is the character set.
3       Let $T^{\prime}$ be the tree that results from subdividing every edge in $T$, let $N=T^{\prime}$
4       // $FA(v)$ is a set that will keep track of FA neighbors present in $v$. for  $C\in\mathcal{C}$ do
5             Compute $\alpha_{T}(C)$, the set of FAs of $C$ in $T$
6             if $|\alpha_{T}(C)|>2$ then return “not galled-completable”
7             if $\alpha_{T}(C)=\{u,v\}$ then add $u$ to $FA(v)$ and add $v$ to $FA(u)$
8
9      for $v$ in postorder($T$) do
10             Let $c_{min}$ be an arbitrary element of $FA(v)$
11             for $u$ in $FA(v)$ do
12                   if $u$ is not comparable to $c_{min}$ then
13                         return “not galled-completable”
14
15                   if $u\prec c_{min}$ then set $c_{min}=u$
16
17            if $|FA(v)|\geq 2$ then
18                   Add $(p_{N}(c_{min}),p_{N}(v))$ to $N$
19
20            else
21                   if  $c_{min}$ is marked as “visited” then
22                         Add $(p_{N}(v),p_{N}(c_{min}))$ to $N$.
23
24                  else
25                         Mark $v$ as “visited”
26
27
28      if  $N$ is a galled tree then return $N$
29       else return “not galled-completable”
30
Algorithm 1 Check if a given tree $T$ is galled-completable.
The previous lemma implies a polynomial time verification algorithm, detailed in Algorithm 1. We build the redundancy-free network and verify that it is a galled tree. We can calculate the set of FAs for each character and use this information to assign the FA neighbors to each node in $T$. Then, in a postorder traversal pick every node $v$ that has FA neighbors. Note that if those neighbors are not all comparable, then there exists no completion, since two outgoing transfers are necessary to explain them. When $v$ has multiple FA neighbors, we find the minimal one ($c_{min}$ in the algorithm) and add the corresponding transfer edge. If $c_{min}$ is the only FA neighbor of $v$, we “mark” $v$. If $c_{min}$ is also marked, then $\{v,c_{min}\}$ is a simple pair, and if not, then either $c_{min}$ will be marked later, or it has multiple FA neighbors and will create its own transfer.
{theoremrep}
Algorithm 1 correctly solves the Galled Tree Completion problem in time $O(|V(T)||\mathcal{C}|)$.
Let us remark that the complexity of the algorithm is dominated by the computation of the set of FA nodes. Assuming a traversal of $T$ for each $C\in\mathcal{C}$, this takes time $O(|V(T)||\mathcal{C}|)$. The rest of the algorithm only adds a time of $O(|V(T)|+|\mathcal{C}|)$. We leave the problem of computing FAs in linear time, if at all possible, for a future discussion.
Proof.
We begin by arguing that the algorithm is correct. The first for loop in Line 1 ensures that each character has at most two FAs as stated in Lemma 3. We claim that the second for loop will find all the required minimal transfers. For a fixed node $v$, we traverse its set of FA neighbors, noting that those are not necessarily comparable. However, if $v$ has two incomparable FA neighbors, then $v$ must also have at least two minimal incomparable FA neighbors. This then implies that the redundancy-free network $N$ will have two transfer edges incident to $p_{N}(v)$, a contradiction of Lemma 3. Thus, Line 1 allows us to rule out these cases. After we ensure that all the FA neighbors of $v$ are comparable we will find the minimum among them, which will allow us to add the desired transfer edge. For simple characters on the other hand, the verification in Line 1 allows us to add an arbitrary direction only when its unique neighbor is a marked node. The reason for this is that whenever the node is marked we know that there do not exist other FA neighbors to compare $c_{min}$ to.
Let us now argue the complexity. Let $T=(V,E)$ be the given tree. The complexity will be dominated by the first for loop. For a given character $C\in\mathcal{C}$, computing the FAs can be done in time $O(|V|)$. Thus the first for can be done in time $O(|\mathcal{C}||V|)$. On the other hand, the postorder traversal of $T$ that adds the transfer edges can be done on time $O(|V|+|\mathcal{C}|)$ since we know that every node can have at most $|\mathcal{C}|-1$ FA neighbors, as characters determine FA neighbors. Finally, the last verification on Line 1 can be done in time $O(|V|)$ (see e.g. [gusfield2014recombinatorics]). ∎
4 The Galled Compatibility Problem
Let us recall the galled compatibility problem: we are given a set of characters $\mathcal{C}$ on a set of taxa $\mathcal{S}$ and must decide whether a galled tree $N$ explains $\mathcal{C}$. Note that if such an $N$ exists, then its base tree is galled-completable. Therefore, $\mathcal{C}$ is galled-compatible if and only if there exists a tree $T$ that is galled-completable for $\mathcal{C}$. Instead of aiming to construct $N$ directly, our strategy is to build such a $T$ using the characterizations from the previous section.
Let $T$ be a tree and $v\in V(T)$. Let $C\in\mathcal{C}$ be a character and suppose that $C$ has two FAs $x_{1},x_{2}$ in $T$. In this case, we say that $C$ is split into $L_{T}(x_{1})$ and $L_{T}(x_{2})$ (noting that the union of these two leafsets must be $C$ since there are only two FAs). A character $C\in\mathcal{C}$ is maximal if there is no $C^{\prime}\in\mathcal{C}$ such that $C\subset C^{\prime}$. Two characters $A,B\in\mathcal{C}$ are compatible if there exists a tree $T$ in which $A$ and $B$ are clades, and incompatible otherwise. It is well-known that $A,B$ are compatible if and only if one of $A\cap B=\emptyset$ or one of $A$ or $B$ is a subset of the other. This means that if $A,B$ are incompatible, then the sets $A\cap B$, $A\setminus B,B\setminus A$ are all non-empty. Recall that for any set $\mathcal{C}$ of pairwise-compatible characters, there is a tree $T$ whose set of clades is exactly $\mathcal{C}$, plus the clade of the root and the leaves (see e.g. [gusfield2014recombinatorics]).
Figure 5: An example instance with characters $\mathcal{C}=\{A,B,C,X,Y,P,Q,R\}$ on taxa $\mathcal{S}=\{a,b,c,d,e,f,g,h,i,j,k,l,m\}$ used to illustrate the main steps 1-4 of the algorithm. Step (1) splits $A$ into $A\setminus B,A\cap B$. Step (2) integrates the clades forced by $X,Y$, which intersect the two clades of $A$. Step (3) integrates $B$ and $C$ as clades. Step (4) solves for $P$ recursively, and then for $Q,R$ recursively, then the transfer required by $A,X$, and $Y$ is added.
The detailed algorithm is somewhat involved, but the main idea can be described in a few steps that are illustrated in Figure 5:
(0) if there is a maximal character $C$ that is compatible with all the others, we may assume that $T$ splits the root with child clades $C$ and $\mathcal{S}\setminus C$ (not shown in figure, see Lemma 4), and that each remaining character is contained in one of these clades. We can solve each subset of characters recursively;
(1) otherwise, every maximal character $A$ has some incompatibility with some $B$, as in Figure 5. We can show that $T$ must either split $A$ into clades $\{A\setminus B,A\cap B\}$, or split $B$ into clades $\{B\setminus A,B\cap A\}$. We try the first option, and if it leads to a dead-end we try the other option. That is, for each of these two possibilities, we initiate a tree with only the two clades and proceed to the next steps. For the rest of the description, we assume that we start a tree with $A\setminus B,A\cap B$ as in Figure 5.1. This means that in any completion, there will be a transfer edge between the two clades to provide an origin for $A$;
(2) each character that intersects both clades must use that same transfer edge in a completion. Such characters enforce new clades that refine the previous ones. For example in Figure 5.2, $X$ enforces the clade $\{d,e\}$, which is the intersection $X\cap\{a,b,c,d,e\}$, and similarly $Y$ enforces $\{c,d,e\}$. These enforced new clades must be ordered by inclusion;
(3) other characters can enforce further clades, namely those that contain the minimal clades enforced so far. For example, the character $C$ from Figure 5 contains $\{d,e\}$, and so the clade $C$ is enforced. Likewise, $B$ is forced since it contains $\{f,g,h\}$. These are added in Figure 5.3;
(4) it turns out that if $\mathcal{C}$ is galled-compatible, then any unforced character so far represents a set of leaves that have the same parent $v$ in $T$. See $P,Q,R$ in Figure 5.3. We can recurse into the leaf children of each $v$ and replace the leaves by the resulting network. In Figure 5.4, the leaf set $\{i,j,k\}$ is replaced by a network that explains $Q,R$, and $\{f,g,h\}$ by a network that explains $P$.
An important subtlety arises in Step (1). If we need to recurse on both possible splits $\{A\setminus B,A\cap B\}$ and $\{B\setminus A,B\cap A\}$, the complexity could become exponential. Our algorithm is designed so that we never have to recurse on both. That is, we actually make a series of checks before trying $\{A\setminus B,A\cap B\}$, and we only recurse after all the checks pass. These checks are designed so that if the recursion fails to find a solution, then $\{B\setminus A,B\cap A\}$ would fail too anyways. This will become apparent in the details below, to which we now proceed.
As explained in step 0 above, we can first show that maximal compatible characters are easy to deal with, see Figure 6.
{lemmarep}
Suppose that $\mathcal{C}$ contains a maximal character $C$ that is compatible with every other character. Let $\mathcal{C}_{1}=\{A\in\mathcal{C}:A\subset C\}$ and $\mathcal{C}_{2}=\{A\in\mathcal{C}:A\cap C\neq\emptyset\}$. Then $\mathcal{C}$ is galled-compatible if and only if $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are both galled-compatible.
Moreover, given galled PTNs $N_{1},N_{2}$ that explain $\mathcal{C}_{1},\mathcal{C}_{2}$, respectively, one can obtain in time $O(|C|)$ a galled PTN $N$ that explains $\mathcal{C}$.
Figure 6: Illustration of Lemma 4. If $C$ is maximal and compatible, we split the problem into two subproblems on $\mathcal{C}_{1}$, which contains only subsets of $C$, and $\mathcal{C}_{2}$, which do not intersect with $C$, and put the resulting networks under a common root.
{proofsketch}
If $C$ as stated exists, we can start constructing a tree whose root has two children, one with $C$ as a clade, and the other with everything not in $C$. This automatically explains $C$. We can then recursively solve for $\mathcal{C}_{1}$ on the $C$ side, and for $\mathcal{C}_{2}$ on the non-$C$ side, and replace the child clades with the corresponding networks. Because $C$ is compatible, every character will be on one side, this solves for every character. ∎
Proof.
The first direction of the if and only if statement is trivial: if $\mathcal{C}$ is galled-compatible, then any of its subset is galled-compatible, including $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, since a network that explains $\mathcal{C}$ also explains $\mathcal{C}_{1},\mathcal{C}_{2}$. In the other direction, suppose that $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are galled-compatible. Since $C$ is compatible with every character, for any $A\in\mathcal{C}\setminus\{C\}$, one of $A\subseteq C,C\subseteq A$, or $A\cap C\neq\emptyset$ holds. Also, because $C$ is maximal, $C\subseteq A$ does not hold. It follows that $A$ is in one of $\mathcal{C}_{1}$ or $\mathcal{C}_{2}$ and thus $\{\mathcal{C}_{1},\mathcal{C}_{2}\}$ is a partition of $\mathcal{C}\setminus\{C\}$. Moreover, by defining $S_{1}:=\bigcup_{A\in\mathcal{C}_{1}}A$ and $S_{2}:=\bigcup_{B\in\mathcal{C}_{2}}B$, we get that $S_{1}\cap C=\emptyset$ and $S_{2}\subseteq C$. In particular, $S_{1}$ and $S_{2}$ are disjoint. Let $N_{1}$ and $N_{2}$ be galled trees that explain $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, respectively. Then $L(N_{1})$ and $L(N_{2})$ are also disjoint. Consider the network $N$ obtained by (1) creating a root node $r$; (2) adding $r(N_{1})$ and $r(N_{2})$ as children of $r$; (3) for each $s\in C\setminus S_{2}$, adding $s$ as a leaf child of $r(N_{2})$. Observe that, since $N_{1}$ and $N_{2}$ are galled trees, by adding $r$ and its two child edges cannot create intersecting cycles, nor can adding leaves under $r(N_{2})$. Moreover, it is not hard to see that every character of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ is still explained by $N$, and $C$ is also explained since it is a clade of $N$. Thus $\mathcal{C}$ is galled-compatible. This construction also shows how to obtain $N$ from $N_{1}$ and $N_{2}$ in time $O(|C|)$, since it only requires adding a new root and adding up to $|C|$ leaves under $r(N_{2})$. ∎∎
The next step is to handle maximal incompatible characters. We first show a fundamental property on pairs of incompatible characters: one of them must be a clade and the other must be split.
{lemmarep}
Let $T$ be a tree that is galled-completable for $\mathcal{C}$, and let $A,B\in\mathcal{C}$ be a pair of incompatible characters. Then one of the following holds:
  • $A$ is split into $A\setminus B,A\cap B$ in $T$, and $B$ is a clade of $T$;
  • $B$ is split into $B\setminus A,A\cap B$ in $T$, and $A$ is a clade of $T$.
{proofsketch}
By incompatibility, $A$ and $B$ cannot both be clades of $T$. If only one of them is, say $A$ is a clade, then we are done since the only way to have two FAs for $B$ is to put $B\cap A$ inside of $A$, and $B\setminus A$ elsewhere. So assume that both $A$ and $B$ are split, having two FAs $a_{1},a_{2}$ and $b_{1},b_{2}$ each. Because $A$ and $B$ intersect, with some effort it can be shown that these FAs must be related by ancestry (i.e., $a_{1}\prec b_{1}$ or $b_{1}\prec a_{1}$, and also $a_{2}\prec b_{2}$ or $b_{2}\prec a_{2}$). The proof shows that each possible case leads to requiring two distinct transfers to explain $A$ and $B$, which create intersecting cycles. ∎
Proof.
Since $A$ and $B$ are incompatible, by definition $T$ cannot contain both as a clade and so at most one can be. So assume for now that neither $A$ nor $B$ is a clade of $T$. Then both characters have two FAs in $T$ (and not more, by Lemma 3). Let $a_{1},a_{2}$ and $b_{1},b_{2}$ be the FA nodes of $A$ and $B$, respectively, and let $A_{1},A_{2},B_{1},B_{2}$ be the respective clades of $a_{1},a_{2},b_{1},b_{2}$.
We claim that some $a_{i}$ is a strict ancestor of some $b_{j}$ or vice-versa. Note that $A\cap B\neq\emptyset$, so we may assume without loss of generality that $A_{1}\cap B_{1}\neq\emptyset$. Thus, $a_{1}$ has a descending leaf in common with $b_{1}$, which implies that $a_{1}$ is an ancestor of $b_{1}$ or vice-versa. If $a_{1}\neq b_{1}$, our claim holds. If $a_{1}=b_{1}$, then $a_{2}\prec b_{2}$ or $b_{2}\prec a_{2}$ by Lemma 3 (part 2), and again our claim holds.
We may therefore assume, without loss of generality, that $a_{1}$ is a strict ancestor of $b_{1}$, which implies $B_{1}\subseteq A_{1}$. Notice that $b_{2}$ cannot be a descendant of $a_{1}$, since otherwise we would have $B_{2}\subseteq A_{1}$ and $B_{1}\cup B_{2}\subseteq A_{1}$, contradicting the incompatibility of $A$ and $B$. Since $b_{2}$ is incomparable with $b_{1}$, it is also incomparable with $a_{1}$. Then by Lemma 3 (part 1), $a_{2}=b_{2}$. Those imply $B_{1}\subseteq A_{1}$ and $A_{2}=B_{2}$, in turn implying $B\subseteq A$ and contradicting that $A$ and $B$ are incompatible.
This establishes that exactly one of $A$ or $B$ is a clade of $T$. Suppose that $B$ is a clade and that $A$ is split into $A_{1}$ and $A_{2}$ in $T$, with respective FAs $a_{1},a_{2}$. Let $b$ be the (unique) FA of $B$. If $b\preceq a_{1}$ or $b\preceq a_{2}$, then $B\subseteq A$ and $A,B$ would be compatible. Moreover, $B$ intersects with at least one of $A_{1}$ and $A_{2}$ since it intersects with $A$, and thus $b$ is an ancestor of $a_{1}$ or $a_{2}$. If $b$ is an ancestor of both, then $A\subseteq B$ and again $A,B$ would be compatible. Hence $b$ is an ancestor of exactly one of $a_{1}$ or $a_{2}$. If $b$ is an ancestor of $a_{1}$, then $A_{1}=B\cap A$ and $A_{2}=A\setminus B$. If $b$ is an ancestor of $a_{2}$, then $A_{2}=B\cap A$ and $A_{1}=A\setminus B$. This shows that the first case of the statement holds.
If $A$ is a clade of $T$ instead, then using the same arguments, we get the second case of the statement. ∎∎
Note that if we allow bidirectional transfers, there are examples in which the above lemma is not always true.
Our algorithm will find maximal incompatible $A$ and $B$ and try splitting $A$, or $B$. When trying the split $A_{1}=A\cap B,A_{2}=B\setminus A$, the characters that intersect with both $A_{1}$ and $A_{2}$ must also be split. Furthermore, one of the FAs of those must be all equal to either that of $A_{1}$ or $A_{2}$, and the other FAs must form a chain under that of $A_{1}$ or $A_{2}$. This can be formalized as follows.
Let $A$ be a character and let $\{A_{1},A_{2}\}$ be a partition of $A$. Let $\mathcal{X}\subseteq\mathcal{C}$ be the set of characters that intersect both $A_{1}$ and $A_{2}$ (note that $A\in\mathcal{X}$). We say that $\mathcal{X}$ forms an $(A_{1},A_{2})$-chain if the elements of $\mathcal{X}$ can be ordered as $\mathcal{X}=\{X_{1},\ldots,X_{l}\}$ such that $X_{l}=A$, and both of the following holds:
  • $(X_{1}\cap A_{1})\subset(X_{2}\cap A_{1})\subset\ldots\subset(X_{l}\cap A_{1})=A_{1}$; and
  • for every $X_{i}\in\mathcal{X}$, $X_{i}\setminus A_{1}=A_{2}$.
We call $X_{1}\cap A_{1}$ the bottom of the chain, and we call $A_{2}$ the stable side of the chain. In Figure 5.3, $A_{1}=\{a,b,c,d,e\},A_{2}=\{f,g,h\}$ and $\mathcal{X}=\{X,Y,A\}$. One can see that $\mathcal{X}$ forms an $(A_{1},A_{2})$-chain with bottom $\{d,e\}$ and stable side $\{f,g,h\}$.
{lemmarep}
Let $A$ be a maximal character of $\mathcal{C}$. Suppose that $T$ is a galled-completable tree for $\mathcal{C}$ in which $A$ is split into the clades $A_{1}$ and $A_{2}$. Let $\mathcal{X}\subseteq\mathcal{C}$ be the subset of characters that intersect with both $A_{1}$ and $A_{2}$.
Then, after possibly exchanging the subscripts of $A_{1}$ and $A_{2}$, $\mathcal{X}$ is an $(A_{1},A_{2})$-chain. Moreover, for every $X\in\mathcal{X}$, the clades $X\cap A_{1}$ and $X\cap A_{2}$ are in $T$.
{proofsketch}
Assuming that $A$ is split into $A_{1},A_{2}$, the maximality of $A$ implies that any $X\in\mathcal{X}$ that intersects with both must also be split. In fact, to avoid creating intersecting cycles, every such $X$ must be explained using the same transfer edge that explains $A_{1}$ and $A_{2}$. This can only be achieved if, in $T$, the FAs of the $X$’s are ordered by ancestry on one side, and all lead to the same clade on the other side. Moreover, because $A$ is maximal, every such $X$ must be a subset of $A$. This results in an $(A_{1},A_{2})$-chain. ∎
Proof.
Let $a_{1},a_{2}$ be the FAs of $A$, which respectively correspond to the clades $A_{1}$ and $A_{2}$. Let $X\in\mathcal{X}\setminus\{A\}$ and note that although $X$ intersects with both $A_{1}$ and $A_{2}$, it cannot contain both, as otherwise $A$ would not be maximal. Thus one of $X\cap A_{1}\subset A_{1}$ or $X\cap A_{2}\subset A_{2}$ holds.
If $X\cap A_{1}\subset A_{1}$, then $a_{1}$ has descending leaves not in $X$, implying that $X$ has a FA $x_{1}$ that strictly descends from $a_{1}$. Since $X$ also intersects with $A_{2}$, $X$ has another FA $x_{2}$ that must be incomparable with $a_{1}$ and, by Lemma 3 (part 1), the other FA $x_{2}$ of $X$ is equal to $a_{2}$. This implies $X\setminus A_{1}=A_{2}$. If $X\cap A_{2}\subset A_{2}$, then instead this argument yields $x_{1}=a_{1},x_{2}\prec a_{2}$ and $X\cap A_{2}\subset A_{2},X\setminus A_{2}=A_{1}$. Suppose for the remainder that $x_{1}\prec a_{1}$ and $x_{2}=a_{2}$. This is without loss of generality, as otherwise we can swap the subscripts of $A_{1}$ and $A_{2}$.
Next, consider another character $X^{\prime}\in\mathcal{X}\setminus\{X,A\}$. As above, $X^{\prime}$ must have two FAs $x^{\prime}_{1},x^{\prime}_{2}$ with either $x^{\prime}_{1}\prec a_{1}$ and $x^{\prime}_{2}=a_{2}$, or vice-versa. By the latter case, we mean $x^{\prime}_{2}\prec a_{2}$ and $x^{\prime}_{1}=a^{\prime}_{1}$, which we claim cannot occur. If this were true, we get $x_{1}\prec a_{1}=x^{\prime}_{1}$ while $x_{2}=a_{2}\neq x^{\prime}_{2}$, contradicting Lemma 3 (part 1). It must then be that $x^{\prime}_{1}\prec a_{1}$ and $x^{\prime}_{2}=a_{2}$.
It follows that every $X^{\prime}\in\mathcal{X}$ has a FA equal to $a_{2}$. By Lemma 3 (part 2), the FAs other than $a_{2}$ of two characters of $\mathcal{X}$ are comparable and, by the above, descend from $a_{1}$. Therefore, the FAs on the $a_{1}$ side are pairwise-comparable in terms of strict ancestry. This implies that in $T$, there is a path from $a_{1}$ to one of its descendants that contains all the FAs of the elements of $\mathcal{X}$ on the $a_{1}$ side. By listing these FAs from the deepest up until $a_{1}$, we get an ordering $X_{1},\ldots,X_{l}=A$ of $\mathcal{X}$ such that $X_{1}\cap A_{1}\subset X_{2}\cap A_{1}\subset\ldots\subset X_{l}\cap A_{1}=A_{1}$. Also, the FAs on the $a_{2}$ side are all equal to $a_{2}$, which means that $X\setminus A_{1}=A_{2}$ for each $X$ of $\mathcal{X}$.
It follows that $\mathcal{X}$ is an $(A_{1},A_{2})$-chain, and that each $X\cap A_{1}$ and each $X\cap A_{2}=A_{2}$ forms a clade in $T$. ∎∎
Lemma 4 shows that if we choose to split $A$ into $A_{1}$ and $A_{2}$, then we know how to split the characters that intersect with both $A_{1}$ and $A_{2}$. We can make a similar deduction for the other characters that contain the bottom or stable side of the $(A_{1},A_{2})$-chain.
{lemmarep}
Let $A$ be a maximal character of $\mathcal{C}$. Suppose that $T$ is a galled-completable tree for $\mathcal{C}$ in which $A$ is split into the clades $A_{1}$ and $A_{2}$. Let $\mathcal{X}\subseteq\mathcal{C}$ be the subset of characters that intersect with both $A_{1}$ and $A_{2}$ and suppose that $\mathcal{X}$ is an $(A_{1},A_{2})$-chain. Let $X_{1}\cap A_{1}$ be the bottom of the chain and let $A_{2}$ be the stable side of the chain.
If $C\in\mathcal{C}\setminus\mathcal{X}$ contains $X_{1}\cap A_{1}$ or contains $A_{2}$, then $C$ is a clade of $T$.
{proofsketch}
Any $C$ as described intersects with exactly one of $X_{1}\cap A_{1}$ or $A_{2}$, but not both (otherwise, it would be in $\mathcal{X}$). In this case, one can show that $C$ and $X_{1}$ must be incompatible. Since $X_{1}$ is assumed to be split, we know by Lemma 4 that $C$ cannot also be split, and thus it must be a clade. ∎
Proof.
Suppose that $C$ contains $X_{1}\cap A_{1}$. If $C=X_{1}\cap A_{1}$, then $T$ contains $C$ by Lemma 4. Otherwise, $C$ is a strict superset of $X_{1}\cap A_{1}$. Note that since $C\notin\mathcal{X}$ and already intersects with $A_{1}$, $C$ does not intersect with $A_{2}$. However, $X_{1}\setminus A_{1}=A_{2}$ by the definition of an $(A_{1},A_{2})$-chain. Therefore, $C\setminus X_{1}$ and $X_{1}\setminus C$ are non-empty, which implies that $C$ and $X_{1}$ are incompatible. By Lemma 4, one of $C$ or $X_{1}$ must be a clade of $T$. We know that $X_{1}$ is already split into $X_{1}\cap A_{1}$ and $X_{1}\cap A_{2}$ in $T$, and thus $C$ must be a clade of $T$.
Suppose that $C$ contains $A_{2}$. If $C=A_{2}$, then $T$ contains $C$ as a clade by assumption. Otherwise, $C$ is a strict superset of $A_{2}=X_{1}\cap A_{2}$. As before, $C$ does not intersect with $A_{1}$, and thus does not intersect with $X_{1}\cap A_{1}$ and is therefore incompatible with $X_{1}$. Again, $C$ is a clade of $T$ by Lemma 4. ∎∎
For an example, see the characters $B$ and $C$ in Figure 5. So far, we have handled characters “forced” by a split of $A$ into $A_{1}$ and $A_{2}$. As it turns out, the other characters can be handled in a recursive manner. To put this precisely, let $A$ be a maximal character of $\mathcal{C}$ and let $\{A_{1},A_{2}\}$ be a partition of $A$ into two non-empty sets. Let $\mathcal{X}$ be the characters that intersect with $A_{1}$ and $A_{2}$ and suppose that $\mathcal{X}$ forms and $(A_{1},A_{2})$-chain. We say that a character $C\in\mathcal{C}$ is forced by $\{A_{1},A_{2}\}$ if either: $C\in\mathcal{X}$; $C$ contains the bottom of the $\mathcal{X}$ chain; or $C$ contains the stable side of the $\mathcal{X}$ chain. Furthermore, a clade $Y\subseteq\mathcal{S}$ is forced by $\{A_{1},A_{2}\}$ if either: $Y=X\cap A_{1}$ or $Y=X\cap A_{2}$ for some $X\in\mathcal{X}$; or $Y=C$ for some character $C\in\mathcal{C}$ that contains the bottom or the stable side of the $\mathcal{X}$ chain.
The next lemma is crucial: it shows that non-forced characters can be grouped and dealt with recursively according to the tree that contains the forced clades.
{lemmarep}
Let $A$ be a maximal character of $\mathcal{C}$ and suppose that there is a galled-completable tree $T^{*}$ for $\mathcal{C}$ in which $A$ is split into $A_{1}$ and $A_{2}$. Let $T$ be the tree whose set of clades is precisely the clades forced by $\{A_{1},A_{2}\}$ (plus the root clade and the leaves).
If $C\in\mathcal{C}$ is a character not forced by $\{A_{1},A_{2}\}$, then all the taxa in $C$ have the same parent in $T$.
Figure 7: An illustration of Lemma 4. Left: the tree $T$ that contains the forced clades, with $X_{1}\cap A_{1}$ the bottom of the chain and $A_{2}$ the stable side. We assume that some characters $B,D,E$ enforced other clades. If we assume that some character $C$ has elements with distinct parents, the situation on the right is unavoidable. A transfer will be needed to link the bottom and stable side, and another transfer to explain $C$. These two transfers create intersecting underlying cycles.
{proofsketch}
Suppose that $T^{*}$ is galled-completable for $\mathcal{C}$. By the previous lemmata, all forced clades must be in $T^{*}$, and thus $T^{*}$ is a “refinement” of $T$ (that is, $T^{*}$ contains the same clades, plus possibly more). We know that a transfer edge is required between the bottom of the chain and the stable side. If there is a non-forced character $C$ with taxa having distinct parents, then no matter how $T$ is refined into $T^{*}$, we will need an additional transfer edge to explain $C$, which will create a cycle that intersects with the one we created near the bottom of the chain. See Figure 7 and the caption for an illustration. ∎
Proof.
Let $T$ and $T^{*}$ be as defined in the lemma statement. Also, let $\mathcal{X}$ be the set of characters that intersect both $A_{1}$ and $A_{2}$. By Lemma 4, we may assume that $\mathcal{X}$ forms an $(A_{1},A_{2})$-chain. Let $X_{1}\in\mathcal{X}$ be such that $X_{b}:=X_{1}\cap A_{1}$ is the bottom of the chain, and $X_{s}:=X_{1}\setminus A_{1}=A_{2}$ is the stable side (using the subscripts $b$ for bottom and $s$ for stable). Let $x_{b}$ be root node of the $X_{b}$ clade in $T$ and let $x_{s}$ be the root node of the $X_{s}$ clade.
By Lemma 4, $T^{*}$ contains all the clades forced by $\{A_{1},A_{2}\}$. Thus $T^{*}$ contains all the clades of $T$, plus possibly others. We assume that a clade present in both $T$ and $T^{*}$ is rooted at the same node in both trees — that is, if $y$ is the root of some clade $Y$ of $T$, then $y$ is also present in $T^{*}$ and is the root of clade $Y$ in $T^{*}$ as well. We will therefore assume that $V(T)\subseteq V(T^{*})$. Notably, $T^{*}$ also contains $X_{b}$ and $X_{s}$, so we assume that the same nodes $x_{b}$ and $x_{s}$ are the roots of these clades in $T^{*}$ as well.
Before proceeding, we establish a fact on the structure of $T$ (as can be seen in Figure 8 below).
Fact 1. Every internal node of $T$ is an ancestor of $x_{b}$ or $x_{s}$, and the root of $T$ is the only node that is an ancestor of both.
Proof We claim that every non-trivial clade of $T$ is either a (not necessarily strict) superset of $X_{b}$ or of $X_{s}$ (a trivial clade is either a single element or all of $L(T)$). Recall that we can distinguish three types of forced clades in $T$: (1) forced clades of the form $X\cap A_{1}$ for $X\in\mathcal{X}$, which by the chain properties must be supersets of $X_{1}\cap A_{1}=X_{b}$; (2) forced clades of the form $X\cap A_{2}$, which are equal to $A_{2}=X_{s}$; (3) forced clades from those those $C^{\prime}$ that contain $X_{b}$ or $X_{s}$. In all cases, our claim holds.
Given this claim, notice that any non-root internal node $v$ of $T$ either roots the clade $A_{1}$, $A_{2}$, or it roots a forced clade. Since any such clade contains $X_{b}$ or $X_{s}$, $v$ must be an ancestor of $x_{b}$ or $x_{s}$.
Now assume that a non-root node $v$ has both $x_{b}$ and $x_{s}$ as descendants. Let $a_{1}$ be the root of the $A_{1}$ clade. Note that $a_{1}$ is an ancestor of $x_{b}$ but not $x_{s}$, so $a_{1}$ is on the path from $x_{b}$ to $x_{s}$. This implies that $v$ is an ancestor of $a_{1}$ and thus the clade corresponding to $v$ contains $A_{1}$. As $v$ is also an ancestor of $x_{s}$, which roots the $A_{2}$ clade, then the clade of $v$ also contains $A_{2}$. Thus, the clade of $v$ contains $A_{1}\cup A_{2}=A$. By assumption, $A$ is not a clade of $T$, so the clade of $v$ is a strict superset of $A$. Moreover, since $v$ is not a root, the clade of $v$ must be in $T$ because it was forced, implying the existence of a character in $\mathcal{C}$ that strictly contains $A$, contradicting its maximality, thereby establishing our fact. ∎
Observe that Fact 1 implies that in $T$, all the children of $x_{b}$ and $x_{s}$ are leaves. In fact, aside from the root, all the nodes on the path between $x_{b}$ and $x_{s}$ have a single non-leaf child, which is the one that leads to $x_{b}$ or $x_{s}$.
Now suppose that some $C\in\mathcal{C}$ is not forced by $\{A_{1},A_{2}\}$, but that there are $y,z\in C$ such that the parent $p_{y}$ of $y$ in $T$ is different from the parent $p_{z}$ of $z$ in $T$. Assume without loss of generality that the distance between $p_{y}$ and the root of $T$ is smaller than or equal to the distance between $p_{z}$ and the root. Since $p_{y}$ and $p_{z}$ are internal nodes of $T$, they are ancestors of $x_{b}$ or $x_{s}$ in $T$ by Fact 1. Figure 7 illustrates one possible scenario where $p_{y}$ is the root of $T$ and $p_{z}$ is on the path between the root and $x_{b}$. Note that $\{p_{y},p_{z}\}=\{x_{b},x_{s}\}$ is not possible, since otherwise $C$ would intersect with both $X_{b}\subseteq A_{1}$ and $X_{s}\subseteq A_{2}$, in which case $C$ should be in $\mathcal{X}$ and be a forced character. Thus at least one of $p_{y}$ or $p_{z}$ is a strict ancestor of $x_{b}$ or $x_{s}$ in $T$. Since $p_{y}$ is closer to the root, we may assume that this holds for $p_{y}$.
Next, consider $T^{*}$, and let $r=lca_{T^{*}}(x_{b},x_{s})$, that is, the lowest common ancestor of $x_{b}$ and $x_{s}$ in $T^{*}$. Note that because $T^{*}$ can have more clades than $T$, $r$ is not necessarily the root of $T^{*}$. Let $P$ be the set of nodes of $T^{*}$ on the path between $r$ and $x_{b}$, or on the path between $r$ and $x_{s}$, excluding $x_{b}$ and $x_{s}$ themselves.
Fact 2. Let $N$ be any galled-completion of $T^{*}$ that explains $\mathcal{C}$. Then there is an underlying cycle in $N$ that contains all the nodes in $P$, plus possibly nodes that descend from $x_{b}$ or $x_{s}$, but no other nodes.
[…]
Figure 8: An illustration of $T$ versus $T^{*}$. The white circles represent clades that are common to $T$ and $T^{*}$.
Our next step is to argue that because of $y$ and $z$, there is some other cycle that intersects with $P$. To this end, define $p_{y}^{\prime}$ as the first ancestor of $y$ in $T^{*}$ that has one of $x_{b}$ or $x_{s}$ as a descendant, and define $p_{z}^{\prime}$ as the first ancestor of $z$ in $T^{*}$ that has one of $x_{b}$ or $x_{s}$ as a descendant. Recall that $p_{y}$ is the parent of $y$ in $T$, but that $p_{y}$ is also in $T^{*}$ and represents the same clade, although $p_{y}$ might not be the parent of $y$ in $T^{*}$. But in $T^{*}$, $p_{y}$ has $y$ plus one of $x_{b}$ or $x_{s}$ as a descendant. Because of this, we have that $p_{y}^{\prime}$ is either equal to $p_{y}$, or it is a descendant of $p_{y}$ (in $T^{*}$). Likewise, $p_{z}^{\prime}$ is a descendant of $p_{z}$ in $T^{*}$. Figure 8 shows a case where $p_{y}$ is a strict ancestor of $p_{y}^{\prime}$ and $p_{z}=p^{\prime}_{z}$.
Fact 3. Let $N$ be any galled-completion of $T^{*}$ for $\mathcal{C}$. Then there is an underlying cycle in $N$ that contains an ancestor of $y$ not in $P$, plus all the nodes on the path between $p_{y}^{\prime}$ and $p_{z}^{\prime}$.
Proof Let us first argue that $p_{y}^{\prime}$ is not a FA node for $C$ in $T^{*}$: if this was the case, because $p_{y}^{\prime}$ has $x_{b}$ or $x_{s}$ as a descendant, having $p_{y}^{\prime}$ as a FA of $C$ would imply that $C$ contains $X_{b}$ or $X_{s}$ and it would be forced. Therefore, there is a FA node $f_{y}$ of $C$ that is a strict descendant of $p_{y}^{\prime}$ and an ancestor of $y$. Likewise, $p^{\prime}_{z}$ is not a FA of $C$ either, and there is a FA $f_{z}$ that is a strict descendant of $p^{\prime}_{z}$ and an ancestor of $z$.
Next, note that $y$ is not a descendant of $p_{z}$ in $T$, as otherwise $p_{y}$ would descend from $p_{z}$ and would be farther from the root. Since $p_{z}$ still represents the same clade in $T^{*}$, the same holds in $T^{*}$. Then in $T^{*}$, since $p_{z}^{\prime}$ is a descendant of $p_{z}$, we deduce that $p_{y}^{\prime}\neq p_{z}^{\prime}$ (otherwise, we would have $y\prec_{T^{*}}p_{y}^{\prime}=p_{z}^{\prime}\preceq_{T^{*}}p_{z}$). Thus, $f_{y}$ and $f_{z}$ are distinct nodes. By Lemma 3, there is a transfer edge between these two FA subtrees. Because $f_{y}$ and $f_{z}$ strictly descend from a node in $P$, the endpoints of this transfer edge are not in $P$, implying the existence of a cycle that contains the created transfer nodes, along with the path from that node to $p_{y}^{\prime}$, then the path from $p^{\prime}_{y}$ to $p_{z}^{\prime}$, followed by the path from $p^{\prime}_{z}$ to the transfer node in that subtree. ∎
We may conclude the proof with that last fact. Consider the cycle $H$ obtained from Fact 2 that contains the nodes of $P$ and possible descendants of $x_{s}$ and $x_{b}$. Then consider the cycle $H^{\prime}$ obtained from Fact 3. The latter has an ancestor of $y$ not in $P$, and because $y$ does not descend from $x_{b}$ or $x_{s}$, these two cycles are different. Moreover, $H^{\prime}$ contains $p^{\prime}_{y}$. If $p^{\prime}_{y}$ is a node of $P$, then $H$ and $H^{\prime}$ intersect, a contradiction. Otherwise, $p^{\prime}_{y}$ must be a strict ancestor of $r$. This can only occur if $p_{y}$ was the root of $T$, and that $p^{\prime}_{y}$ ended up as an ancestor of $lca_{T^{*}}(x_{b},x_{s})$ when refining $T$ to $T^{*}$. In this case however, $p_{z}$ could not be the root of $T$ as well, and $p_{z}$ must be a descendant of $r$. Therefore, $H^{\prime}$ contains $r$ and again, this implies that $H$ and $H^{\prime}$ intersect, which concludes the proof. ∎∎
By combining the elements gathered so far, we finally arrive at an algorithmically useful characterization of galled-compatible characters.
{lemmarep}
Let $A$ be a maximal character of $\mathcal{C}$ and let $\{A_{1},A_{2}\}$ be a partition of $A$. Then there is a tree that is completable for $\mathcal{C}$ and that contains the clades $A_{1}$ and $A_{2}$ if and only if all the following conditions hold:
  1. 1.
    Let $\mathcal{X}$ be the characters that intersect with $A_{1}$ and $A_{2}$. Then $\mathcal{X}$ forms an $(A_{1},A_{2})$-chain;
  2. 2.
    There exists a tree $T$ on leafset $\mathcal{S}$ whose set of clades is precisely the set of clades forced by $\{A_{1},A_{2}\}$ (plus the root clade and leaves).
  3. 3.
    Let $\mathcal{C}_{F}$ be the set of characters forced by $\{A_{1},A_{2}\}$. Then for any $C\in\mathcal{C}\setminus\mathcal{C}_{F}$, all the taxa of $C$ have the same parent in $T$.
  4. 4.
    For any subset $\mathcal{C}^{\prime}\subseteq\mathcal{C}\setminus\mathcal{C}_{F}$ such that all taxa that belong to some character of $\mathcal{C}^{\prime}$ have the same parent in $T$, $\mathcal{C}^{\prime}$ is galled-compatible.
{proofsketch}
In the ()($\Rightarrow$)( ⇒ ) direction, we know by all the previous lemmata that all the stated conditions must be satisfied by a galled-completable tree for $\mathcal{C}$. In the ()($\Leftarrow$)( ⇐ ) direction, suppose that all the conditions hold. Let $T$ be a tree that contains all the clades forced by $\{A_{1},A_{2}\}$. Then we can explain all the forced characters by adding a transfer edge just above the bottom of the $\mathcal{X}$ chain, going just above the stable side. As for the other characters, for a node $v$ of $T$, let $\mathcal{C}_{v}\subseteq\mathcal{C}$ be the non-forced characters having taxa whose parent is $v$. We can find a galled network $N_{v}$ that explains $\mathcal{C}_{v}$ and “attach” $N_{v}$ as a child of $v$. Doing this for every $v$ does not create intersecting cycles and explains all the remaining characters. ∎
Proof.
Suppose that there is a tree $T^{*}$ that contains the clades $A_{1},A_{2}$, such that $T^{*}$ is galled-completable for $\mathcal{C}$. Condition 1 holds because by Lemma 4, $\mathcal{X}$ forms an $(A_{1},A_{2})$-chain. For Condition 2, by Lemma 4 and 4, $T^{*}$ contains all the clades forced by $\{A_{1},A_{2}\}$, and therefore $T$ as described in the condition must exist. Condition 3 holds by Lemma 4. Finally, Condition 4 holds because by the existence of $T^{*}$, we have that $\mathcal{C}$ is galled-compatible. By heredity, any subset $\mathcal{C}^{\prime}\subseteq\mathcal{C}$ is galled-compatible, including those described in the condition.
In the converse direction, assume that all the conditions of our statement hold. We show how to obtain a tree $T^{*}$ with clades $A_{1}$ and $A_{2}$ that is galled-completable for $\mathcal{C}$. Let $T$ be the tree whose set of clades consists of the clades forced by $\{A_{1},A_{2}\}$. Let $\mathcal{X}$ be the characters that intersect $A_{1}$ and $A_{2}$. Let $X_{1}\in\mathcal{X}$ such that $X_{1}\cap A_{1}$ is the bottom of the chain and $X_{1}\setminus A_{1}=A_{2}$. We know that $T$ has the clades $X_{1}\cap A_{1}$ and $A_{2}$, since they are forced. For later reference, we let $x_{b}$ and $x_{s}$ be the roots of these clades in $T$, respectively.
Next, for each internal node $v\in V(T)$, let $\mathcal{C}_{v}$ be the characters whose taxa all have $v$ as their parent in $T$. We know that $\mathcal{C}_{v}$ is galled-compatible, so there is a tree $T_{v}$ that is galled-completable for $\mathcal{C}_{v}$. We modify $T$ as follows: remove every leaf that is a child of $v$, and add the root of $T_{v}$ as a new child of $v$. After doing this for every $v$, we obtain the tree $T^{*}$.
We claim that $T^{*}$ is galled-completable for $\mathcal{C}$. First, notice that $T^{*}$ contains all the clades of $T$, since we have only removed leaves (trivial clades) and replaced them by subtrees, which does not remove clades that were already present in $T$. Thus every clade forced by $\{A_{1},A_{2}\}$ is in $T^{*}$. In particular, $X_{1}\cap A_{1}$ and $A_{2}$ are still clades of $T^{*}$, still rooted at $x_{b}$ and $x_{s}$, respectively. Add a transfer edge from the parent branch of $x_{b}$ to the parent branch of $x_{s}$. This yields a network that we call $N$.
Let $\mathcal{C}_{F}$ be the characters forced by $\{A_{1},A_{2}\}$. We will add further edges to $N$, but for the moment we claim that $N$, with the single transfer edge $(p_{N}(x_{b}),p_{N}(x_{s}))$, explains the characters in $\mathcal{C}_{F}$. Recall that there are three types of forced clades, which we deal with as follows:
  • let $C\in\mathcal{C}_{F}$ such that $C$ contains the bottom of the chain, namely $X_{1}\cap A_{1}$. Then $C$ is a forced clade, which is in $T$ and therefore in $T^{*}$. Thus $T^{*}$ explains $C$ without requiring any transfer.
  • let $C\in\mathcal{C}_{F}$ such that $C$ contains the stable side of the chain, namely $A_{2}$. Again, $C$ is forced in $T$ and thus $T^{*}$ and is also explained.
  • let $C\in\mathcal{C}_{F}$ such that $C\in\mathcal{X}$. Then $C$ intersects both $A_{1}$ and $A_{2}$. By the definition of an $(A_{1},A_{2})$-chain, $C\cap A_{1}$ is either equal to $X_{1}\cap A_{1}$, or it is a strict superset of $X_{1}\cap A_{1}$. Also by the definition of chains, $C\cap A_{2}=A_{2}$. We know that $C\cap A_{1}$ and $C\cap A_{2}$ are forced by $\{A_{1},A_{2}\}$ and are thus present in $T$ and $T^{*}$. The network $N$ can explain $C$ by putting the origin at the root of the $C\cap A_{1}$ clade, and using the transfer edge $(p(x_{b}),p(x_{s}))$ to send the character to $C\cap A_{2}=A_{2}$. Note that this is possible since $C\cap A_{1}$ is a superset of $X_{1}\cap A_{2}$, and thus the root of the $C\cap A_{1}$ clade is an ancestor of $p(x_{b})$, or is equal to $x_{b}$ when $C=X_{1}$ (in which case the transfer can still be used).
It only remains to explain the characters not in $\mathcal{C}_{F}$. For each subtree $T_{v}$, replace $T_{v}$ by a galled completion $N_{v}$ that explains $\mathcal{C}_{v}$, which exists by assumption. This ensures that each character from each $\mathcal{C}_{v}$ is explained by the resulting network. Moreover, since the $T_{v}$’s are independent subtrees, all additional cycles created are entirely contained inside the $N_{v}$ subnetworks and, in particular, do not contain $v$. This implies that we do not create intersecting cycles that involve nodes from two distinct $N_{v}$ subnetworks, and that do not intersect with the cycle involving $(p(x_{b}),p(x_{s}))$. Therefore, the resulting network is a galled tree. This concludes the proof. ∎∎
Our strategy is then to find a maximal character $A$ and some $B$ it is incompatible with. We try to split $A$ into $A\setminus B$ and $A\cap B$ and if all the conditions of Lemma 4 pass, we have succeeded. Otherwise, it is $B$ that we split into $B\setminus A$ and $B\cap A$. We want to check the conditions on this split, but since Lemma 4 only apply to maximal characters, then $B$ must be maximal as well. We thus need the following.
{lemmarep}
Suppose that $\mathcal{C}$ has no maximal character that is compatible with all the others. Then there exists a pair of incompatible characters $\{A,B\}$ of $\mathcal{C}$ such that $A$ and $B$ are both maximal.
{proofsketch}
Let $A$ be maximal. Then $A$ has some incompatibility. Letting $B$ be the character of maximum cardinality such that $A,B$ are incompatible, one can argue that $B$ is also maximal. ∎
Proof.
Let $A$ be a maximal character of $\mathcal{C}$. By assumption, there is some $B\in\mathcal{C}$ such that $A$ and $B$ are incompatible. Choose such a $B$ of maximum cardinality. Note that $A\cap B$, $B\setminus A$, $A\setminus B$ are all non-empty. Suppose that $B$ is not maximal, i.e. there is $B^{\prime}\in\mathcal{C}$ such that $B\subset B^{\prime}$. We have $A\cap B^{\prime}\neq\emptyset$ because $B^{\prime}$ contains $B$. Also, $B^{\prime}\subset A$ is not possible since $B\setminus A$ is non-empty, and $A\subset B^{\prime}$ is not possible since $A$ is maximal. Thus $A$ and $B^{\prime}$ are incompatible, contradicting the choice of $B$. ∎∎
This summarizes the elements required for our algorithm. Due to space constraints, the full pseudo-code and analysis is relegated to the Appendix. However, its main steps can be summarized as follows:
  1. 1.
    If there is a maximal character compatible with every other, proceed recursively using Lemma 4.
  2. 2.
    Otherwise, find maximal incompatible characters $A,B$, guaranteed by Lemma 4.
    Initiate a tree whose clades are $A\cap B$ and $A\setminus B$.
  3. 3.
    Check whether Conditions 1-2-3 of Lemma 4 hold with respect to that partition. If so, consider the tree $T$ that contains all forced clades. For each $v\in V(T)$, list the set of characters $\mathcal{C}_{v}$ that only contain taxa whose parent is $v$. Solve for every such $\mathcal{C}_{v}$ recursively. If they all succeed, then we return that $\mathcal{C}$ is galled-compatible. If one of them fails, we may return that no solution is possible (because if one $\mathcal{C}_{v}$ is not galled-compatible, then $\mathcal{C}$ is not either).
  4. 4.
    If one of Condition 1-2-3 of Lemma 4 does not hold, then repeat the previous step but with a tree that contains clades $B\cap A,B\setminus A$. If this also fails, $\mathcal{C}$ is not galled-compatible.
Importantly, note that in this algorithm, we do not enter a recursion into $\mathcal{C}_{v}$ subsets for both partition attempts $\{A\cap B,A\setminus B\}$ and $\{B\cap A,B\setminus A\}$, as otherwise the complexity could be exponential. Since recursions are performed in at most one attempt of applying Step 3, we can argue that this takes polynomial time.
{toappendix}
Algorithm 2 is the full pseudocode of the galled-compatibility algorithm.
1 fonction getGalledTree($\mathcal{C}$)
2       if $\mathcal{C}=\emptyset$ then return $true$
3       if $\mathcal{C}$ has a maximal character $C$ compatible with every $C^{\prime}\in\mathcal{C}$ then
4             Let $\mathcal{C}_{1}=\{A\in\mathcal{C}:A\subset C\},\mathcal{C}_{2}=\{A\in\mathcal{C}:A\cap C=\emptyset\}$
5             return $getGalledTree(\mathcal{C}_{1})\wedge getGalledTree(\mathcal{C}_{2})$
6
7
8       Let $\{A,B\}$ be a pair of maximal incompatible characters from $\mathcal{C}$
9       Let $\{A_{1},A_{2}\}=\{A\setminus B,A\cap B\}$
10       $result=tryPartition(\mathcal{C},\{A_{1},A_{2}\})$
11       if $result=``yes"$ then return $true$
12       if $result=``no"$ then return $false$
13
14       //otherwise, $result=``invalid~{}partition"$, try the other partition
15       Let $\{B_{1},B_{2}\}=\{B\setminus A,B\cap A\}$
16       $result=tryPartition(\mathcal{C},\{B_{1},B_{2}\})$
17       if $result=``yes"$ then return $true$, otherwise return $false$
18
19 fonction tryPartition($C,\{A_{1},A_{2}\}$)
20       Let $\mathcal{X}$ be the characters that intersect $A_{1}$ and $A_{2}$
21       if $\mathcal{X}$ is not an $(A_{1},A_{2})$-chain nor an $(A_{2},A_{1})$-chain then return “invalid partition”
22       Let $\mathcal{C}_{F}$ and $F$ be the characters and clades, respectively, forced by $\{A_{1},A_{2}\}$
23       Let $T$ be the tree whose set of non-trivial clades is $F$
24       if $T$ does not exist then return “invalid partition”
25       if there is $C\in\mathcal{C}\setminus\mathcal{C}_{F}$ and $u,v\in C$ with distinct parents in $T$ then  return “invalid partition”
26       foreach $v\in V(T)$ do
27             Let $\mathcal{C}_{v}\subset\mathcal{C}\setminus\mathcal{C}_{F}$ be the characters only containing taxa whose parent is $v$
28             if $\mathcal{C}_{v}\neq\emptyset\wedge getGalledTree(\mathcal{C}_{v})=false$ then return “no”
29
30      return “yes”
31
32
Algorithm 2 Main galled-tree reconstruction algorithm.
The idea of the algorithm is as follows.
  1. 1.
    If $\mathcal{C}$ has a maximal character $C$ compatible with every character, then we know that we can simply recurse into $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ asin Lemma 4.
  2. 2.
    Otherwise, we must deal with maximal characters that are all incompatible. We choose incompatible $A,B$ that are both maximal (shown to exist in Lemma 4).
  3. 3.
    We know that a galled-completable tree, if one exists, must contain the clades $A\setminus B,A\cap B$, or $B\setminus A,B\cap B$. We do not know which one it is, so we try the former first. That is, first consider the partition $\{A_{1},A_{2}\}=\{A\setminus B,A\cap B\}$.
  4. 4.
    The function $tryPartition$ checks whether it is possible to satisfy Conditions 1,2,3 from Lemma 4 with $\{A_{1},A_{2}\}$ enforced. If one of those conditions fails, we know that $A_{1},A_{2}$ cannot lead to a solution, and $tryPartition$ returns “invalid partition”. When $getCalledTree$ receives this answer, it attempts to check whether using $\{B_{1},B_{2}\}=\{B\setminus A,B\cap A\}$ works instead, which is the only other possibility.
    Do note that Lemma 4 only applies to maximal characters. Since we potentially check them on $A$ and $B$, it is crucial to have both $A$ and $B$ as maximal characters.
  5. 5.
    If $\{A_{1},A_{2}\}$ passes the three tests in $tryPartition$, then we check recursively that all the $\mathcal{C}_{v}$’s are galled-compatible. If they all are, we return “yes” and we are done. If some $\mathcal{C}_{v}$ is not galled-compatible, then we know that $A_{1},A_{2}$ cannot lead to a solution.
    An important subtlety arises in this case. We do not return “invalid partition”, because this would lead to trying with $\{B_{1},B_{2}\}$ instead. This could lead to an exponential time algorithm, because we would recurse into too many cases on both $\{A_{1},A_{2}\}$ and $\{B_{1},B_{2}\}$. Instead, we return “no”, and the main algorithm returns $false$ without even trying $\{B_{1},B_{2}\}$. This is correct, because if we find a $\mathcal{C}_{v}$ that is not galled-compatible, we know that $\mathcal{C}$ itself is not galled-compatible and there is no point in trying $\{B_{1},B_{2}\}$.
The correctness and complexity details of the algorithm now follow.
{theoremrep}
The Galled Compatibility problem can be solved in time $O(n|\mathcal{C}|^{3})$.
Proof.
We first prove by induction on $|\mathcal{C}|$ that the algorithm always returns the correct answer. As a base case, when $|\mathcal{C}|=0$, $\mathcal{C}$ is trivially galled-compatible and the algorithm correctly returns true.
So suppose for the inductive step that $|\mathcal{C}|>0$. If $\mathcal{C}$ has a maximal compatible character $C$, then by Lemma 4, $\mathcal{C}$ is galled-compatible if and only if $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are galled-compatible. Since, by induction, the algorithm returns the correct answer on both $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, returning $getGalledTree(\mathcal{C}_{1})\wedge getGalledTree(\mathcal{C}_{2})$ is correct.
So suppose that $\mathcal{C}$ does not have a maximal compatible character. Note that by Lemma 4, the algorithm will find maximal incompatible $A$ and $B$. First assume that $\mathcal{C}$ is galled-compatible, in which case we need to argue that the algorithm returns $true$. Let $T^{*}$ be a tree that is galled-completable for $\mathcal{C}$. By Lemma 4, $T^{*}$ contains either the clades $\{A_{1},A_{2}\}=\{A\setminus B,A\cap B\}$ or $\{B_{1},B_{2}\}=\{B\setminus A,B\cap A\}$. Assume first that $T^{*}$ contains $A_{1},A_{2}$. Then all the conditions of Lemma 4 hold for $A_{1},A_{2}$. Thus, all the conditions verified by $tryPartition$ on input $\{A_{1},A_{2}\}$ will succeed (including the calls to $getGalledTree(\mathcal{C}_{v})$, which are assumed to return $true$ by induction). It follows that $tryPartition$ will return “yes” and that $getGalledTree$ will correctly return $true$.
Next, assume instead that $T^{*}$ contains the clades $\{B_{1},B_{2}\}$. If Algorithm 2 gets to call $tryPartition(\mathcal{C},\{B_{1},B_{2}\})$, then as in the previous case, we know that all the tests made by $tryPartition$ will pass and that it will return “yes”. However, we will not reach that point if the prior call $tryPartition(\mathcal{C},\{A_{1},A_{2}\})$ has returned “yes” or “no”. If that previous call returned “yes”, then $getGalledTree$ will return $true$, which is actually the correct answer. A problem occurs if this previous call returned “no” on input $\{A_{1},A_{2}\}$. By inspecting $tryPartition$, we see that this only occurs when there is a $\mathcal{C}_{v}$ on which $getGalledTree(\mathcal{C}_{v})$ returns $false$. By induction, this means that $\mathcal{C}_{v}$ is not galled-compatible, implying in turn that $\mathcal{C}$ is not galled-compatible, a contradiction. Thus, we may assume that $tryPartition$ on input $\{A_{1},A_{2}\}$ either returns “yes” (in which case we correctly return $true$), or “invalid partition”, in which case we correctly return $true$ by then trying $\{B_{1},B_{2}\}$.
In the converse direction, suppose that $\mathcal{C}$ is not galled-compatible. Then one of the four conditions of Lemma 4 must fail on both $\{A_{1},A_{2}\}$ and $\{B_{1},B_{2}\}$. This means that on either inputs, $tryPartition$ will either return “no” or “invalid partition”, which leads $getGalledTree$ to correctly returning $false$.
Now let us argue the complexity. Let us first analyze the recursive search tree $R$ created by the algorithm, where each node of $R$ corresponds to a call to $getGalledTree$. We show by induction on the height of $R$ that $R$ contains at most $3\cdot\max(1,|\mathcal{C}|)$ nodes. As a base case, when $R$ has height 00, it is a terminal case with $1\leq 3$ node, in which case our claim holds (even if $\mathcal{C}$ is empty). So assume that $R$ has higher height. Thus $\mathcal{C}$ is non-empty. If $\mathcal{C}$ has a maximal compatible character $C$, then we make recursive calls on disjoint strict subsets $\mathcal{C}_{1},\mathcal{C}_{2}$ of $\mathcal{C}$, none of which contains $C$. If $\mathcal{C}_{1},\mathcal{C}_{2}$ are non-empty, by induction the number of nodes of the recursion tree is at most $13|\mathcal{C}_{1}|+3|\mathcal{C}_{2}|+$ 1 (counting the root), which is at most $3(|\mathcal{C}|-1)+1\leq 3|\mathcal{C}|$ since $|\mathcal{C}_{1}|+|\mathcal{C}_{2}|<|\mathcal{C}|$. If, say, $\mathcal{C}_{1}$ is empty but not $\mathcal{C}_{2}$ (or vice-versa), the number of nodes in the recursion tree is at most $1+3|\mathcal{C}_{2}|+1<1+3(|\mathcal{C}|-1)+1\leq 3|\mathcal{C}|$. If both $\mathcal{C}_{1},\mathcal{C}_{2}$ are empty, the recursion tree has $3$ nodes, which is at most $3|\mathcal{C}|$.
If no maximal compatible $C$ exists, the algorithm makes recursive calls on $\mathcal{C}_{v}$ sets in $tryPartition$, either when it receives $\{A_{1},A_{2}\}$ or $\{B_{1},B_{2}\}$. Suppose that recursive calls are made when $\{A_{1},A_{2}\}$ is received. Again, we observe that we either return “yes” or “no”, and then $getGalledTree$ exits without attempting $\{B_{1},B_{2}\}$. It thus suffices to count the recursive calls in one call of $tryPartition$, on disjoint subsets $\mathcal{C}_{v}$ of $\mathcal{C}$ (which are non-empty since this is checked by the algorithm). Note that these subsets do not contain $A$. Hence, the number of nodes in $R$ is at most
$\sum_{v\in V(T)}3|\mathcal{C}_{v}|+1\leq 3(|\mathcal{C}|-1)+1\leq 3|\mathcal{C}|$
since all the $\mathcal{C}_{v}$’s are disjoint and none of them contains $A$ from $\mathcal{C}$. If instead some recursive calls are made when $\{B_{1},B_{2}\}$ is the input to $tryPartition$, we can repeat the same analysis and reach the same conclusion. This proves our claim.
It thus only remains to analyze the time needed to handle one node of the recursion tree. Recall that $n$ is the number of taxa. Testing compatibility of two characters requires computing set operations in time $O(n)$. Finding a maximal compatible character, or a pair of incompatible maximal characters, can be achieved by testing compatibility between each pair of characters, taking a total time of $O(n|\mathcal{C}|^{2})$. This also allows finding $A_{1},A_{2},B_{1},B_{2}$ in the same complexity.
During one recursion, we run $tryPartition$ at most once, say on $A_{1},A_{2}$. Computing $\mathcal{X}$ can be done in time $O(n|\mathcal{C}|)$ by computing two intersections for each character against $A_{1},A_{2}$. Testing the chain property can be done in time $O(n|\mathcal{X}|)=O(n|\mathcal{C}|)$ by sorting the $\mathcal{X}$’s by size (using e.g. bucket sort) and verifying the inclusions. It is straightforward to construct a tree $T$ in time $O(n|\mathcal{C}|^{2})$ from the forced clades $\mathcal{C}_{F}$ (or decide that it does not exist) by relating them by set inclusion and building the corresponding tree top-down from the maximal clades to the minimal ones. Checking the “same parent” condition and building the $\mathcal{C}_{v}$ sets is easily seen to not take more than $O(n|\mathcal{C}|^{2})$ time.
Overall, we spend time $O(n|\mathcal{C}|^{2})$ in one call to $getGalledTree$ and $tryPartition$. Since the recursion tree has $O(|\mathcal{C}|)$ nodes, the total time is $O(n|\mathcal{C}|^{3})$. ∎∎
5 Galled PTNs on functional characters: a case study
We now turn to a real-data case analysis of galled completion in order to gain better insight on its potential for the prediction horizontal gene transfers. Its applicability is not fully understood, mainly due to the lack of benchmarks for the interactions between transfers and characters. However, a few studies have now started exploring this territory, and here we take inspiration from [machine], which implemented machine learning approaches to infer HGTs from functional characters using the Kyoto Encyclopedia of Genes and Genomes (KEGG) Orthology database [KEGG] (see below). The authors used curated HGTs between bacteria and the list of functions of each taxa as features, allowing them to predict the presence or absence of HGT between any two bacterial species. They inferred 147,889 transfers between 6566 bacterial genomes, although some of these transfers could all be manifestations of one ancestral transfer in the phylogeny (but because these predictions are made in a pairwise manner, the time of these transfers cannot be inferred by the approach).
We also used functional characters from the KEGG database, which associates molecular functions of genes and proteins to orthologous groups. Table 1 lists the functions that we used, along with their hierarchical classification. The main type of entries in the database are KEGG Orthologs (KO), which are groups of genes sharing common functions (members of the same group are viewed as functional orthologs). Most groups are formed based on experimental evidence, with some generalizations to other organisms based on sequence similarity.
We took a subset of 45 species from the bacterial species used in [machine], which consists of species that were predicted to be involved in interphylum transfers. We obtained the corresponding species tree from NCBI Taxonomy Browser [schoch2020ncbi], noting that it is not completely resolved and is therefore non-binary. The whole annotated genomes of these species are contained in KEGG. As characters, we chose 23 KOs based on the features that shown as most important for the Graph Convolutional Network model presented in [machine]. These features include metabolism-related, information processing and antibiotic resistance KOs among other functions. The whole datasets and scripts used for this part are contained in the following repository: https://github.com/AliLopSan/ptns.
Character Brite Hierarchy Classification Function Description K18220 01504 Antimicrobial resistance genes ribosomal protection tetracycline resistance protein K02257 09100 Metabolism heme o synthase K01610 09101 Carbohydrate metabolism phosphoenolpyruvate carboxykinase (ATP) K04068 09191 Unclassified: metabolism anaerobic ribonucleoside -triphosphate reductase activating protein K00627 09101 Carbohydrate metabolism pyruvate dehydrogenase E2 component K00241 09101 Carbohydrate metabolism succinate dehydrogenase cytochrome b subunit K01679 09101 Carbohydrate metabolism fumarate hydratase, class II K13628 03016 Transfer RNA biogenesis iron-sulfur cluster assembly protein K01669 03400 DNA repair and recombination proteins deoxyribodipyrimidine photo-lyase K03980 09183 Protein families: signaling and cellular processes putative peptidoglycan lipidII flippase K06886 99996 General function prediction only hemoglobin K07305 99980 Enzymes with EC numbers peptide-methionine (R)-S-oxide reductase K01589 00230 Purine metabolism 5-(carboxyamino) imidazole ribonucleotide synthase K00561 03009 Ribosome biogenesis 23S rRNA (adenine-N6)-dimethyltransferase K07483 99976 Replication and Repair transposase K17836 01501 beta-Lactam resistance beta-Lactam resistance K02227 00860 Porphyrin metabolism adenosylcobinamide-phosphate synthase K18214 01504 Antimicrobial resistance genes MFS transporter, DHA3 family, tetracycline resistance protein K19310 09131 Membrane transport bacitracin transport system permease protein K19115 02048 Prokaryotic defense system CRISPR-associated protein Csh2 K02274 00190 Oxidative phosphorylation cytochrome c oxidase subunit I K03737 00720 Carbon fixation pathways in prokaryotes pyruvate-ferredoxin/flavodoxin oxidoreductase K00850 00010 Glycolysis / Gluconeogenesis 6-phosphofructokinase 1
Table 1: List of the complete set of characters used for our experiments which includes their function classification according to the KEGG BRITE database and a brief description of their function.
Figure 9: Distribution of the number of first-appearances per character. We distinguish those first-appearance nodes that are leaves and those that are inner nodes. Note that in general characters have more than two first-appearances, thus if there exists a PTN that can explain them, this is not necessarily galled. We note however that there exists a set of characters that have at most two first appearances:$\{K13628,K19115,K00561,K19310\}$ which are those that fall below the green dashed line.
Figure 10: An example that shows how allowing losses can lead to a completion that requires less transfers. We show the first appearances of character $K016699$. Note that in order to fulfill the connectivity requirements of Lemma 1 transfer edges should connect every colored block. Notice that with the no loss condition, there exist three clades at the right-hand side of (a) that need at least two transfers to remain connected since the species $T03770$ does not contain the character. In (b) we show that this could be explained by a loss event at species $T03770$, thus reducing by two the number of transfers that are needed to connect the updated first-appearance nodes.
Figure 11: (a) The species tree with 45 species with annotated leaves showing colored annotations by clade according to the presence of one of the characters: $\{K13628,K19115,K00561,K19310\}$. Note that between annotations of the same color there should exist a transfer edge that could join the parts of the tree where the characters are present. Leaf names correspond to the species id present in the KEGG orthology database. Notice that this network is not galled since the cycles resulting from joining the first appearances of characters $K13628$ and $K00561$ contain the root of the tree as a common node. However as shown in (b), by resolving one of the children of the root node, as the black dot, this network becomes galled. Tree annotations where visualized using iTOL tool  [itol].
Using the Fitch labeling algorithm described in [ptn], we first calculated the number of first-appearance nodes of each character as well as their position in the species tree. As established above, this serves as a first test for galled-completability, as characters with three of more such nodes are sufficient to discard this property. Figure 9 shows the distribution of the number of first-appearance nodes for each character. It is immediately clear that the “at most two first-appearances” condition is far from being satisfied, although it is insightful to understand why. For several characters, the majority of first-appearance nodes are leaves. An example of this phenomenon is shown in Figure 10. Notably, there is a clade in the top-right part of the tree in which every leaf has the character except one, resulting in three first-appearances in that clade. This prevents the common ancestor of the clade from being labeled with that character. It is plausible that one character was simply lost in one species, without invoking any transfer, as shown on the right. This strongly indicates that the “transfer-only” model is too rigid, and in future work we plan to work on approaches to identify situations in which losses are more likely than transfers.
Going back to Figure 9, we note that there does exist characters that posses at most two first-appearances. We checked whether this subset of four characters could at least be galled-completable, but it turns out that this is not even the case. This is shown in Figure 11 on the left. As explained in the caption, if we built the redundancy-free network for these characters and applied Lemma 3, we would add a transfer edge linking the branches above the roots of the blue character, and above the roots of the green character. These are non-redundant edges that create two cycles that intersect at the root. But one could hypothesize that the unresolved root is responsible for this: Figure 11 on the right partially resolves the tree by expanding the root, creating a branch in a way that the cycle of the blue characters now avoids the root. One can now check that the resulting redundancy-free network is galled, and that this set of characters is galled-completable.
To sum up, our ad hoc study leads to several insights and future directions. It shows that the possibility of gene losses cannot be ignored, and that these should either be predicted in a pre-processing step, or included as part of the model altogether. The analysis also shows that unresolved species trees may be an obstacle to galled completions. On the other hand, our approach may be useful in resolved such species trees.
6 Conclusion
In this work, we expanded the foundations on perfect phylogenies to galled trees. While compatible character in trees enjoy an elegant mathematical structure, it appears that the difficulty of characterizing compatibility ramps up very quickly as we move beyond trees. Nonetheless, this work can serve as a stepping-stone towards the understanding of character evolution on more complex structures. A small step in this direction would be to allow bidirectional transfers in galled trees. This change is not as trivial as it seems, since several of our key results do not hold when such edges are present. It will also be interesting to integrate broader classes of networks, for example level-$k$ networks, and to solve the open problem of finding a PTN with a minimum number of transfer edges.
Perhaps a more important direction, as exemplified by our case study, is to remove the “no-loss” assumption on characters. In the case of trees, one of the closest extension of perfect phylogenies is the Dollo parsimony model, where a character can be lost multiple times across the phylogeny, but can never be regained after it has been lost. Extending the perfect transfer networks to a form of Dollo parsimony may be interesting to reduce the “non-transfer noise” seen in our experimental study. It is worth noting though that any character can be explained with a single emergence at the root and losses under Dollo parsimony. That is, transfers are never needed to explain a character, and therefore some weighing scheme between transfers and losses will need to be developed before the model can become applicable on real data at a large scale. Examples of other directions include extensions to multi-state characters (as initiated by [nakhlehtesis]), finding maximal sets of characters that fit in a galled tree, or devising non-binary species tree resolution algorithms.

### v2

@@
 During the process, we provide several structural characterizations of characters that can be displayed on a given tree or a network.
-To improve the reading experience, we provide a sketch for most proofs, and refer to the appendix for detailed proofs.
-We also propose a case study of perfect transfer networks on a real dataset consisting of bacterial species and functional characters obtained the Kyoto Encyclopedia of Genes and Genomes (KEGG) database [KEGG].
+Althought most of our theoretical results are intuitive, the detailed proofs are sometimes involved.
+To improve the reading experience, we provide a sketch for most proofs, and refer to the appendix for the full details.
+We also propose a case study of perfect transfer networks on a real dataset consisting of bacterial species and functional characters obtained from the Kyoto Encyclopedia of Genes and Genomes (KEGG) database [KEGG].
 We show that the conditions required to explain a set of characters with a galled tree are difficult to achieve, as most characters prevent such an explanation, even when taken individually.
@@
 A tree $T$ is a network with no underlying cycle.
-We write $u\preceq_{T}v$ if $v$ is on the path from $\rho(T)$ to $u$, in which case $v$ is an ancestor of $u$ and $u$ a descendant of $v$.
+We write $u\preceq_{T}v$ if $v$ is on the path from $\rho(T)$ to $u$, in which case $v$ is an ancestor of $u$ and $u$ a descendant of $v$ (we may drop the $T$ subscript if unambiguous).
 Note that $v$ is an ancestor and descendant of itself.
-Two nodes $u$,$v$ are incomparable in $T$ if none descends from the other, i.e.
-if neither $u\preceq v$ nor $v\preceq u$.
+Two nodes $u$,$v$ are incomparable in $T$ if none descends from the other, i.e., if neither $u\preceq v$ nor $v\preceq u$.
 The parent of a node $v$ in $T$ is $p_{T}(v)$.
@@
 The set $L(v)$ is called a clade of $T$.
-An LGT network (where LGT comes from Lateral, or Horizontal, Gene Transfers[Cardona_2015] is a network $N=(V,E_{S}\cup E_{T})$, where $\{E_{S},E_{T}\}$ is a specified partition of the edge-set of $N$, such that the subgraph $\mathcal{T}_{N}:=(V,E_{S})$ is a tree with the same set of nodes as $N$.
+An LGT network (where LGT comes from Lateral Gene Transfers[Cardona_2015] is a network $N=(V,E_{S}\cup E_{T})$, where $\{E_{S},E_{T}\}$ is a specified partition of the edge-set of $N$, such that the subgraph $\mathcal{T}_{N}:=(V,E_{S})$ is a tree with the same set of nodes as $N$.
 The tree $\mathcal{T}_{N}$ is called the support tree of $N$.
@@
 We also assume that transfer nodes have exactly one out-neighbor in $\mathcal{T}_{N}$.
-The tree obtained from $\mathcal{T}_{N}$ by suppressing its subdivision nodes is called the base tree of $G$ [^1] Suppressing a subdivision node $u$ with parent $p$ and child $v$ consists of removing $u$ and adding an edge from $p$ to $v$.
+The tree obtained from $\mathcal{T}_{N}$ by suppressing its subdivision nodes is called the base tree of $N$ [^1] Suppressing a subdivision node $u$ with parent $p$ and child $v$ consists of removing $u$ and adding an edge from $p$ to $v$.
 For $v\in V(N)\setminus\{\rho(N)\}$, we use $p_{N}(v)$ to denote the unique in-neighbor of $v$ in the support tree of $N$.
@@
   3. 3.
-    for each $C\in\mathcal{C}$, there exists a unique node $v\in V$ that reaches every node $w$ satisfying $C\in l(w)$ (single origin).
+    for each $C\in\mathcal{C}$, there exists a unique node $v\in V$ that, in $N$, reaches every node $w$ satisfying $C\in l(w)$ (single origin).
 Furthermore, we call $N$ a perfect transfer network (PTN) for $\mathcal{C}$ if there exists a $\mathcal{C}$-labeling of $N$ that explains $\mathcal{C}$.
@@
 (b) A PTN with $T$ as base tree that is not a galled tree (two underlying cycles contain the left child of the root).
-(c) A galled-completion of $T$.
-Note that in (b) and (c) for every character $C$, there exists a node (in gray) that transmits to all its vertical descendants, and reaches every leaf in $C$.
+Notice for example the circle character $C_{2}$, which does not use any transfer and only needs to be transmitted vertically to its descendants after emergence.
+On the other hand, the square character $C_{1}$ needs to be transferred to two other ancestral species (left and right) after emergence.
+The triangle character $C_{3}$ needs one transfer, and uses the same arc going to the left as $C_{1}$.
+(c) A different PTN with $T$ as a base tree, which is a galled-completion of $T$.
+In short, each character $C\in\mathcal{C}$ must emerge at some node $v$, then be transmitted to every vertical descendant, and possibly horizontally to other species (and if so, such other species must in turn also transmit to their vertical descendants, and possibly transfer horizontally as well).
 Figure 1 shows two PTNs for the same set of characters.
@@
 {lemmarep}
-Let $N=(V,E_{S}\cup E_{T})$ be a galled LGT network and let $v\in V(N)$ be a node with two distinct descendants $x$ and $y$ that are transfer nodes (with $v\in\{x,y\}$ being possible).
+Let $N=(V,E_{S}\cup E_{T})$ be a galled LGT network and let $v\in V(N)$ be a node with two distinct descendants $x$ and $y$ in $\mathcal{T}_{N}$ that are transfer nodes (with $v\in\{x,y\}$ being possible).
 Then one of the transfer edges incident to $x$ or $y$ has both endpoints in the subtree $\mathcal{T}_{N}(v)$.
@@
 [warnow2024statistically] also independently developed galled tree reconstructions algorithms from character data (which are assumed to be SNPs in the paper).
-Therein, for each character, the species can be in one of two states, and the states are assumed to correspond to a bipartition of a galled tree $N$, where a bipartition is a partition of the leaves that can be obtained by taking a tree displayed by $N$, removing an edge, and taking the two sets of leaves of the resulting connected components (here, “displayed” means a tree that can be obtained by removing one parent edge from each node of in-degree two).
+Therein, for each character, the species can be in one of two states, and the states are assumed to correspond to a bipartition of a galled tree $N$.
+A bipartition is a partition of the leaves that can be obtained by taking a tree displayed by $N$, removing an edge, and taking the two sets of leaves of the resulting connected components (here, “displayed” means a tree that can be obtained by removing one parent edge from each node of in-degree two).
 They discuss algorithms to reconstruct a galled tree that contains a set of given characters, that is, that contains all the corresponding bipartitions.
@@
 If we interpret one state as “presence” and the other as “absence”, this is very similar to our formulation, with two notable exceptions.
-First, we assume an ordering in states from “absence” to “presence”, whereas no such ordering is imposed in the PPN model.
+First, in our work on PTNs we impose an ordering in states from “absence” to “presence”, whereas no such ordering is assumed in the PPN model.
 More importantly, the notion of displaying a tree does not distinguish between transfer and vertical descent edges.
@@
 In the model of [warnow2024statistically], this character is considered as explained, since the displayed tree on the right shows that the character indeed corresponds to a bipartition.
-On the other hand, this network does not explain the character under the PTN model (see caption).
+On the other hand, this network does not explain the character under the PTN model.
 Roughly speaking, displaying a tree allows removing a vertical edge, whereas we would require transmitting characters along that removed vertical edge.
@@
 The colored square represents a character present at the two rightmost leaves (or in the PPN formulation, there are two states: with or without a square).
-(b) A tree displayed by the network, obtained by deleting one edge (we left the resulting nodes with a single child, although they could be suppressed).
+(b) A tree displayed by the network, obtained by deleting one edge (we left the resulting subdivision node, although they could be suppressed).
 The illustrated character labeling shows that this network explains the character under the PPN model.
-As for PTNs, one can see from the definition of $F_{C}(N)$ that under the PTN formulation, no internal node of the initial network can contain the character, and thus that the network cannot explain it.
+As for PTNs, one can see from the definition of $F_{C}(N)$ that under the PTN formulation, no internal node of the initial network can contain the character, that there is no possible origin, and thus that the network cannot explain it.
 3 The Galled Completion Problem
@@
 The corresponding sets of FAs for every character are as follows: $\alpha_{T}(C_{1})=\{x,y\},\alpha_{T}(C_{2})=\{y,s_{2}\},\alpha_{T}(C_{3})=\{v\},\alpha_{T}(C_{4})=\{u\}$.
-For a character $C\in\mathcal{C}$, we will say that a node $v$ is an origin for $C$ if $v$ reaches every leaf in $C$ in $N-F_{C}(N)$.
-By Lemma 1, our goal is to ensure that every character has an origin.
+For a character $C\in\mathcal{C}$ and a network $N$, we will say that a node $v$ is an origin for $C$ if $v$ reaches every leaf in $C$ in $N-F_{C}(N)$.
+By Lemma 1, our goal is to add arcs to $T$ to ensure that every character has an origin.
 We first show that in any galled-completion of a tree $T$, an origin of a character $C$ must descend from a FA node (or it could be a transfer node added just above it), and it must “give” its character to all other FAs by transferring just above them.
@@
 Then there is $\alpha_{i}\in\alpha_{T}(C)$ such that both of the following hold:
-  • either $w\preceq\alpha_{i}$ or $w$ is a transfer node whose child in $\mathcal{T}_{N}$ is $\alpha_{i}$.
+  • either $w\preceq_{\mathcal{T}_{N}}\alpha_{i}$ or $w$ is a transfer node whose child in $\mathcal{T}_{N}$ is $\alpha_{i}$.
   • for every $\alpha_{j}\in\alpha_{T}(C)\setminus\{\alpha_{i}\}$ there is a transfer edge $(u,v)$ such that $u\preceq w$ and $v=p_{N}(\alpha_{j})$.
@@
 By definition, in $T$ the ancestors of FAs are in $F_{C}(N)$ since they have descendants not in $C$.
-Thus, an origin $w$ in $N$ must be in a FA subtree, or it could be a transfer node added above when adding transfer edges from $T$ to $N$.
-If there are multiple FAs, $w$ must be able to reach the other ones, and Lemma 2.2 lets us establish that transfer edges are needed to achieve this.
+Thus, an origin $w$ in $N$ must be in a FA subtree, or it could be a transfer node inserted above such an FA during the process of adding transfer edges from $T$ to $N$.
+If there are multiple FAs, $w$ must be able to reach the other ones, and Lemma 2.2 lets us establish that transfer edges as described are needed to achieve this.
 ∎
@@
 First observe that if $v\in V(T)$, then $v\in V(N)$ as well.
-Moreover, the parent $p_{v}$ of $v$ in $T$ is either also the parent of $v$ in $\mathcal{T}_{N}$, or the latter is a transfer node created to complete $T$ into $N$.
-In the latter case, note that at most one transfer node could could be present on the path between $p_{v}$ and $v$ in $\mathcal{T}_{N}$, as otherwise, there would be two consecutive transfer nodes which, as one can check, would violate Lemma 2.2.
-In other words, some edges of $T$ are subdivided to obtain $\mathcal{T}_{N}$, but each edge can be subdivided at most once.
+Moreover, the the parent of $v$ in $\mathcal{T}_{N}$ is either also the parent $p_{v}$ of $v$ in $T$, or it could be a transfer node created to complete $T$ into $N$.
+In the latter case, note that at most one transfer node could could be present on the path between $p_{v}$ and $v$ in $\mathcal{T}_{N}$, as otherwise, there would be two consecutive transfer nodes which, as one can check, would violate Lemma 2.2 ($p_{v}$ would have two descendants with outgoing transfers).
+This lets us establish that some edges of $T$ are subdivided to obtain $\mathcal{T}_{N}$, but each edge can be subdivided at most once.
 Let us next observe that if $w$ is an origin for $C$, by definition it must be a node in $N-F_{C}(N)$, which is a graph that only contains FA nodes of $T$, their support tree descendants in $N$, and possibly transfer nodes that are the parents of FA nodes.
@@
 Therefore, the first condition of our statement holds.
-For the second condition, as we know the origin $w$ descends from some $\alpha_{i}$ or is its parent.
+For the second condition, as we now know the origin $w$ descends from some $\alpha_{i}$ or is its parent.
 Let $\alpha_{j}$ be a FA other than $\alpha_{i}$.
@@
 Since $w$ is an origin, it reaches $l$ in $N-F_{C}(N)$ and thus also in $N$.
-As $w$ is not an ancestor of $l$, by Lemma 2.2 there is a transfer edge $(u,v)$ in $N$ such that $u\preceq w$ and $l\preceq v$ in $\mathcal{T}_{N}$.
+Since first-appearance nodes are incomparable in $T$ and $\mathcal{T}_{N}$, $w$ is not an ancestor of $l$ in $\mathcal{T}_{N}$.
+By Lemma 2.2 there is a transfer edge $(u,v)$ in $N$ such that $u\preceq w$ and $l\preceq v$ in $\mathcal{T}_{N}$.
 The endpoints of this transfer edge must also be in $N-F_{C}(N)$ for $w$ to be able to use it.
@@
 We next argue that the galled requirement places an important limitation on FAs, as there can be at most two per character.
-This means that a character $C$ must either be a clade of $T$, or it could be split in two clades.
+This means that a character $C$ must either be a clade of $T$, or it could be split in two clades, but not more.
+This is both useful in theory, but also in our experiments since it allows checking quickly whether an individual character can be explained by a galled-tree.
 {lemmarep}
-Let $T$ be a galled-completable tree for a character set $\mathcal{C}$. Then $\alpha_{T}(C)\leq 2$ for any character $C\in\mathcal{C}$.
+Let $T$ be a galled-completable tree for a character set $\mathcal{C}$. Then for any character $C\in\mathcal{C}$, ${\alpha_{T}(C)\leq 2}$.
 {proofsketch}
@@
 Likewise, there is a similar transfer connecting $b_{1}$ and $b_{2}$.
-Since $a_{2}$ and $b_{2}$ are incomparable, these two transfers must be distinct, they are both in $\mathcal{T}_{N}(a_{1})$ or $\mathcal{T}_{N}(p_{N}(a_{1}))$, and they both have external endpoints, contradicting Lemma 2.2.
+Since $a_{2}$ and $b_{2}$ are incomparable, these two transfers must be distinct, they are both in $\mathcal{T}_{N}(a_{1})$ or $\mathcal{T}_{N}(p_{N}(a_{1}))$, and they both have external endpoints since $b_{2}$ is not comparable with $a_{1}$, contradicting Lemma 2.2.
 Since neither case is possible, we must have $a_{2}=b_{2}$.
@@
 ∎∎
-Note that the first statement of the proof of Lemma 3 requires transfer edges to be unidirectional.
-If we allowed bidirectional transfer edges, then we can devise examples in which the lemma does not hold (not shown due to space constraints).
-Allowing such edges already leads to more complex structures.
+Note that in the detailed proof of the first statement of Lemma 3, we use the fact that transfer edges are unidirectional — the statement in the sketch that $b_{1}$ needs to use the “same” transfer as $a_{1}$ remains true, but they do not need to send the character to the same clade in a bidirectional setting, as the sending could go both ways.
+If we allowed bidirectional transfer edges, we can devise examples in which the lemma does not hold.
+This shows that even apparently minor changes to the model can lead to more complex structures.
 We can begin describing our algorithmic strategy.
@@
 {proofsketch}
-In the forward direction, if $T$ can be completed into some LGT network $N^{\prime}$, we can argue that for every transfer edge $(u,v)$ added to $N$, there is a corresponding transfer edge in $N^{\prime}$ whose endpoints are either $u$ and $v$, or their support tree descendants.
+In the forward direction, if $T$ can be completed into a galled tree $N^{\prime}$ that explains $\mathcal{C}$, we can argue that for every transfer edge $(u,v)$ the is present in the redundancy-free network $N$, there is a corresponding transfer edge in $N^{\prime}$ whose endpoints are either $u$ and $v$, or in their support tree descendants.
 Roughly speaking, this means that the cycles of $N$ have corresponding cycles in $N^{\prime}$, and since $N^{\prime}$ is a galled tree, so is $N$.
@@
 Since $N$ also subdivides every edge once, this allows us to assume that $N^{\prime}$ and $N$ have the same set of nodes, putting the nodes in correspondence without ambiguity.
-We first map each transfer edge $(u^{\prime},v^{\prime})$ of $N$ to a distinct transfer edge $(x,y)$ of $N^{\prime}$, such that $y\in\{u^{\prime},v^{\prime}\}$.
+We first build an injective map, that associates each transfer edge $(u^{\prime},v^{\prime})$ of $N$ with a distinct transfer edge $(x,y)$ of $N^{\prime}$, such that $y\in\{u^{\prime},v^{\prime}\}$.
 Let $(u^{\prime},v^{\prime})\in E_{T}$.
@@
 Either way, we map $(u^{\prime},v^{\prime})$ to $(x,y)$.
-Note that by Lemma 3, $u^{\prime}$ and $v^{\prime}$ are only incident to one transfer edge $(u^{\prime},v^{\prime})$.
-Therefore, only $(u^{\prime},v^{\prime})$ can be mapped to $(x,y)$, as $y\in\{u^{\prime},v^{\prime}\}$.
-Next, consider $(u^{\prime},v^{\prime})\in E_{T}$ and its corresponding edge $(x,y)$ in $N^{\prime}$.
+We claim that no other transfer edge of $N$ can be mapped to $(x,y)$ in this manner.
+Suppose that some other transfer edge $(u^{\prime\prime},v^{\prime\prime})$ of $N$ maps to $(x,y)$.
+According to our map, $y=u^{\prime\prime}$ or $y=v^{\prime\prime}$, but either way, $y$ is incident to two transfer arcs in $N$, contradicting Lemma 3.
+Therefore, only $(u^{\prime},v^{\prime})$ can be mapped to $(x,y)$.
+Next, consider $(u^{\prime},v^{\prime})\in E_{T}$ and its associated edge $(x,y)$ in $N^{\prime}$.
 Suppose that $(x,y)=(\tilde{u},v^{\prime})$ for some descendant $\tilde{u}$ of $u^{\prime}$ in $\mathcal{T}_{N^{\prime}}$.
@@
 The same idea applies if $y=u^{\prime}$ instead.
-Since all underlying cycles of $N^{\prime}$ are independent, we can apply the transformation from the previous paragraph to every mapped transfer edge of $N^{\prime}$.
-Notice that because each $(u^{\prime},v^{\prime})\in E_{T}$ maps to a distinct transfer edge of $N^{\prime}$, no edge of $N^{\prime}$ is “moved” twice by this process.
+Since all underlying cycles of $N^{\prime}$ are independent, we can apply the transformation from the previous paragraph to replace every mapped transfer edge of $N^{\prime}$ by a transfer edge of $N$, one after another, while maintaining the property that no cycles intersect.
+Notice that because each $(u^{\prime},v^{\prime})\in E_{T}$ maps to a distinct transfer edge of $N^{\prime}$, every edge of $N$ is incorporated, and no edge of $N^{\prime}$ is “moved” twice by this process.
 Therefore, we obtain a galled tree whose set of transfer edges contains $E_{T}$.
@@
 20            else
-21                   if  $c_{min}$ is marked as “visited” then
+21                   if  $c_{min}$ is marked as “possibly simple” then
 22                         Add $(p_{N}(v),p_{N}(c_{min}))$ to $N$.
@@
 24                  else
-25                         Mark $v$ as “visited”
+25                         Mark $v$ as “possibly simple”
 26
@@
 Thus the first for can be done in time $O(|\mathcal{C}||V|)$.
-On the other hand, the postorder traversal of $T$ that adds the transfer edges can be done on time $O(|V|+|\mathcal{C}|)$ since we know that every node can have at most $|\mathcal{C}|-1$ FA neighbors, as characters determine FA neighbors.
+On the other hand, the postorder traversal of $T$ that adds the transfer edges can be done in time $O(|V|+|\mathcal{C}|)$.
+This is because we traverse $O(|V|)$ nodes, and each $\{u,v\}$ relationship of FA neighborhood requires work at most twice throughout the algorithm, once for $u$ and once for $v$.
+Moreover, each character implies at most one FA relationship, the total work over the FA neighbor hood relationships is $O(|\mathcal{C}|)$.
+We note that checking for the incomparability or descendance of nodes can be achieved in constant time with standard pre-processing of the tree.
 Finally, the last verification on Line 1 can be done in time $O(|V|)$ (see e.g.
 [gusfield2014recombinatorics]).
-∎
+∎∎
 4 The Galled Compatibility Problem
@@
 Two characters $A,B\in\mathcal{C}$ are compatible if there exists a tree $T$ in which $A$ and $B$ are clades, and incompatible otherwise.
-It is well-known that $A,B$ are compatible if and only if one of $A\cap B=\emptyset$ or one of $A$ or $B$ is a subset of the other.
+It is well-known that $A,B$ are compatible if and only if either $A\cap B=\emptyset$ or one of $A$ or $B$ is a subset of the other.
 This means that if $A,B$ are incompatible, then the sets $A\cap B$, $A\setminus B,B\setminus A$ are all non-empty.
@@
 It follows that $A$ is in one of $\mathcal{C}_{1}$ or $\mathcal{C}_{2}$ and thus $\{\mathcal{C}_{1},\mathcal{C}_{2}\}$ is a partition of $\mathcal{C}\setminus\{C\}$.
-Moreover, by defining $S_{1}:=\bigcup_{A\in\mathcal{C}_{1}}A$ and $S_{2}:=\bigcup_{B\in\mathcal{C}_{2}}B$, we get that $S_{1}\cap C=\emptyset$ and $S_{2}\subseteq C$.
+Moreover, by defining $S_{1}:=\bigcup_{A\in\mathcal{C}_{1}}A$ and $S_{2}:=\bigcup_{B\in\mathcal{C}_{2}}B$, we get that $S_{1}\subseteq C$ and $S_{2}\cap C=\emptyset$.
 In particular, $S_{1}$ and $S_{2}$ are disjoint.
@@
 Then $L(N_{1})$ and $L(N_{2})$ are also disjoint.
-Consider the network $N$ obtained by (1) creating a root node $r$; (2) adding $r(N_{1})$ and $r(N_{2})$ as children of $r$; (3) for each $s\in C\setminus S_{2}$, adding $s$ as a leaf child of $r(N_{2})$.
+Consider the network $N$ obtained by (1) creating a root node $r$; (2) adding $r(N_{1})$ and $r(N_{2})$ as children of $r$; (3) for each $s\in C\setminus S_{1}$, adding $s$ as a leaf child of $r(N_{2})$.
 Observe that, since $N_{1}$ and $N_{2}$ are galled trees, by adding $r$ and its two child edges cannot create intersecting cycles, nor can adding leaves under $r(N_{2})$.
@@
 By incompatibility, $A$ and $B$ cannot both be clades of $T$.
-If only one of them is, say $A$ is a clade, then we are done since the only way to have two FAs for $B$ is to put $B\cap A$ inside of $A$, and $B\setminus A$ elsewhere.
+If only one of them is a clade, say $A$, then we are done since the only way to have two FAs for $B$ is to put $B\cap A$ inside of $A$, and $B\setminus A$ elsewhere.
 So assume that both $A$ and $B$ are split, having two FAs $a_{1},a_{2}$ and $b_{1},b_{2}$ each.
@@
 Notice that $b_{2}$ cannot be a descendant of $a_{1}$, since otherwise we would have $B_{2}\subseteq A_{1}$ and $B_{1}\cup B_{2}\subseteq A_{1}$, contradicting the incompatibility of $A$ and $B$.
-Since $b_{2}$ is incomparable with $b_{1}$, it is also incomparable with $a_{1}$.
+Since $b_{2}$ is incomparable with $b_{1}$, it also cannot be an ancestor of $a_{1}$, and thus $b_{1}$ is also incomparable with $a_{1}$.
 Then by Lemma 3 (part 1), $a_{2}=b_{2}$.
@@
 One can see that $\mathcal{X}$ forms an $(A_{1},A_{2})$-chain with bottom $\{d,e\}$ and stable side $\{f,g,h\}$.
+The concept of a chain is illustrated in Figure 7.
+Figure 7: An $(A_{1},A_{2})$-chain, where $X_{1},X_{2},X_{3}\in\mathcal{X}$ all intersect with both $A_{1}$ and $A_{2}$.
+On the $A_{1}$ side, the FAs are ordered by ancestry, and the only way to explain all these characters is to have them use the same transfer to reach the $A_{2}$ side.
 {lemmarep}
@@
 Proof.
-Let $a_{1},a_{2}$ be the FAs of $A$, which respectively correspond to the clades $A_{1}$ and $A_{2}$.
+Let $a_{1},a_{2}$ be the FAs of $A$ in $T$, which respectively correspond to the clades $A_{1}$ and $A_{2}$.
 Let $X\in\mathcal{X}\setminus\{A\}$ and note that although $X$ intersects with both $A_{1}$ and $A_{2}$, it cannot contain both, as otherwise $A$ would not be maximal.
@@
 As it turns out, the other characters can be handled in a recursive manner.
-To put this precisely, let $A$ be a maximal character of $\mathcal{C}$ and let $\{A_{1},A_{2}\}$ be a partition of $A$ into two non-empty sets.
+To put this precisely, let us gather all the information on the tree that has been shown to be forced so far.
+Let $A$ be a maximal character of $\mathcal{C}$ and let $\{A_{1},A_{2}\}$ be a partition of $A$ into two non-empty sets.
 Let $\mathcal{X}$ be the characters that intersect with $A_{1}$ and $A_{2}$ and suppose that $\mathcal{X}$ forms and $(A_{1},A_{2})$-chain.
@@
 Furthermore, a clade $Y\subseteq\mathcal{S}$ is forced by $\{A_{1},A_{2}\}$ if either: $Y=X\cap A_{1}$ or $Y=X\cap A_{2}$ for some $X\in\mathcal{X}$; or $Y=C$ for some character $C\in\mathcal{C}$ that contains the bottom or the stable side of the $\mathcal{X}$ chain.
+The first type of clades are enforced due to Lemma 4, and the second type due to Lemma 4.
 The next lemma is crucial: it shows that non-forced characters can be grouped and dealt with recursively according to the tree that contains the forced clades.
@@
 Let $x_{b}$ be root node of the $X_{b}$ clade in $T$ and let $x_{s}$ be the root node of the $X_{s}$ clade.
-By Lemma 4, $T^{*}$ contains all the clades forced by $\{A_{1},A_{2}\}$.
+By Lemma 4 and Lemma 4, $T^{*}$ contains all the clades forced by $\{A_{1},A_{2}\}$.
 Thus $T^{*}$ contains all the clades of $T$, plus possibly others.
@@
 Since any such clade contains $X_{b}$ or $X_{s}$, $v$ must be an ancestor of $x_{b}$ or $x_{s}$.
-Now assume that a non-root node $v$ has both $x_{b}$ and $x_{s}$ as descendants.
+Now assume that a non-root node $v$ has both $x_{b}$ and $x_{s}$ as descendants in $T$.
 Let $a_{1}$ be the root of the $A_{1}$ clade.
-Note that $a_{1}$ is an ancestor of $x_{b}$ but not $x_{s}$, so $a_{1}$ is on the path from $x_{b}$ to $x_{s}$.
+Note that $a_{1}$ is an ancestor of $x_{b}$ but not $x_{s}$, so $a_{1}$ is on the path from $x_{b}$ to $x_{s}$ (when viewing $T$ as an undirected graph).
 This implies that $v$ is an ancestor of $a_{1}$ and thus the clade corresponding to $v$ contains $A_{1}$.
@@
 Then there is an underlying cycle in $N$ that contains all the nodes in $P$, plus possibly nodes that descend from $x_{b}$ or $x_{s}$, but no other nodes.
-[…]
+Proof Since $X_{1}$ is split into $X_{b}$ and $X_{s}$ in $T^{*}$, by Lemma 3 there must be a transfer edge between $p_{N}(x_{b})$ or a descendant of $x_{b}$, and $p_{N}(x_{s})$ and a descendant of $x_{s}$, which implies that existence of the claimed cycle.
+∎
 Figure 9: An illustration of $T$ versus $T^{*}$. The white circles represent clades that are common to $T$ and $T^{*}$.
@@
 We will add further edges to $N$, but for the moment we claim that $N$, with the single transfer edge $(p_{N}(x_{b}),p_{N}(x_{s}))$, explains the characters in $\mathcal{C}_{F}$.
-Recall that there are three types of forced clades, which we deal with as follows:
+Recall that there are three types of forced characters, which we deal with as follows:
   • let $C\in\mathcal{C}_{F}$ such that $C$ contains the bottom of the chain, namely $X_{1}\cap A_{1}$.
@@
 ∎∎
-This summarizes the elements required for our algorithm. Due to space constraints, the full pseudo-code and analysis is relegated to the Appendix. However, its main steps can be summarized as follows:
-  1. 1.
-    If there is a maximal character compatible with every other, proceed recursively using Lemma 4.
-  2. 2.
-    Otherwise, find maximal incompatible characters $A,B$, guaranteed by Lemma 4.
-    Initiate a tree whose clades are $A\cap B$ and $A\setminus B$.
-  3. 3.
-    Check whether Conditions 1-2-3 of Lemma 4 hold with respect to that partition.
-If so, consider the tree $T$ that contains all forced clades.
-For each $v\in V(T)$, list the set of characters $\mathcal{C}_{v}$ that only contain taxa whose parent is $v$.
-Solve for every such $\mathcal{C}_{v}$ recursively.
-If they all succeed, then we return that $\mathcal{C}$ is galled-compatible.
-If one of them fails, we may return that no solution is possible (because if one $\mathcal{C}_{v}$ is not galled-compatible, then $\mathcal{C}$ is not either).
-  4. 4.
-    If one of Condition 1-2-3 of Lemma 4 does not hold, then repeat the previous step but with a tree that contains clades $B\cap A,B\setminus A$.
-If this also fails, $\mathcal{C}$ is not galled-compatible.
-Importantly, note that in this algorithm, we do not enter a recursion into $\mathcal{C}_{v}$ subsets for both partition attempts $\{A\cap B,A\setminus B\}$ and $\{B\cap A,B\setminus A\}$, as otherwise the complexity could be exponential.
-Since recursions are performed in at most one attempt of applying Step 3, we can argue that this takes polynomial time.
-{toappendix}
 Algorithm 2 is the full pseudocode of the galled-compatibility algorithm.
+For simplicity, it only solves the decision version of the problem (i.e., whether the characters are galled-compatible or not), but it can easily be adapted to reconstruct a network.
 1 fonction getGalledTree($\mathcal{C}$)
@@
 If both $\mathcal{C}_{1},\mathcal{C}_{2}$ are empty, the recursion tree has $3$ nodes, which is at most $3|\mathcal{C}|$.
-If no maximal compatible $C$ exists, the algorithm makes recursive calls on $\mathcal{C}_{v}$ sets in $tryPartition$, either when it receives $\{A_{1},A_{2}\}$ or $\{B_{1},B_{2}\}$.
+If no maximal compatible $C$ exists, the algorithm makes recursive calls on $\mathcal{C}_{v}$ sets in $tryPartition$, either when it receives $\{A_{1},A_{2}\}$ or $\{B_{1},B_{2}\}$ (but not both).
 Suppose that recursive calls are made when $\{A_{1},A_{2}\}$ is received.
@@
 5 Galled PTNs on functional characters: a case study
-We now turn to a real-data case analysis of galled completion in order to gain better insight on its potential for the prediction horizontal gene transfers.
-Its applicability is not fully understood, mainly due to the lack of benchmarks for the interactions between transfers and characters.
+We now turn to a real-data case analysis of galled-completion in order to gain better insight on its potential for the prediction horizontal gene transfers, as there is a general lack of benchmarks for the interactions between transfers and characters.
 However, a few studies have now started exploring this territory, and here we take inspiration from [machine], which implemented machine learning approaches to infer HGTs from functional characters using the Kyoto Encyclopedia of Genes and Genomes (KEGG) Orthology database [KEGG] (see below).
-The authors used curated HGTs between bacteria and the list of functions of each taxa as features, allowing them to predict the presence or absence of HGT between any two bacterial species.
+The authors of  [machine] used curated HGTs between bacteria, with the list of functions of each taxa as features, allowing them to predict the presence or absence of HGT between any two bacterial species.
+They therefore solve a binary classification task, making one prediction per species pair.
 They inferred 147,889 transfers between 6566 bacterial genomes, although some of these transfers could all be manifestations of one ancestral transfer in the phylogeny (but because these predictions are made in a pairwise manner, the time of these transfers cannot be inferred by the approach).
-We also used functional characters from the KEGG database, which associates molecular functions of genes and proteins to orthologous groups.
+In our case study, we also used functional characters from the KEGG database, which associates molecular functions of genes and proteins to orthologous groups.
 Table 1 lists the functions that we used, along with their hierarchical classification.
@@
 We obtained the corresponding species tree from NCBI Taxonomy Browser [schoch2020ncbi], noting that it is not completely resolved and is therefore non-binary.
-The whole annotated genomes of these species are contained in KEGG.
-As characters, we chose 23 KOs based on the features that shown as most important for the Graph Convolutional Network model presented in [machine].
+Observe that this is not a problem for our tree-completion procedure, as only the inserted transfer nodes are assumed to be binary in our model — not the nodes of the input tree.
+The whole annotated genomes of these species are contained in KEGG, and so as characters, we chose 23 KOs based on the features that were shown as most important for the Graph Convolutional Network model presented in [machine].
 These features include metabolism-related, information processing and antibiotic resistance KOs among other functions.
@@
 We distinguish those first-appearance nodes that are leaves and those that are inner nodes.
-Note that in general characters have more than two first-appearances, thus if there exists a PTN that can explain them, this is not necessarily galled.
+Most characters have more than two first-appearances, thus if there exists a PTN that can explain them, it will not be galled.
 We note however that there exists a set of characters that have at most two first appearances:$\{K13628,K19115,K00561,K19310\}$ which are those that fall below the green dashed line.
@@
 Leaf names correspond to the species id present in the KEGG orthology database.
-Notice that this network is not galled since the cycles resulting from joining the first appearances of characters $K13628$ and $K00561$ contain the root of the tree as a common node.
+Notice that the redundancy-free network is not galled since the cycles resulting from joining the first appearances of characters $K13628$ and $K00561$ contain the root of the tree as a common node.
 However as shown in (b), by resolving one of the children of the root node, as the black dot, this network becomes galled.
 Tree annotations where visualized using iTOL tool  [itol].
+Losses can obstruct galled-completability.
 Using the Fitch labeling algorithm described in [ptn], we first calculated the number of first-appearance nodes of each character as well as their position in the species tree.
-As established above, this serves as a first test for galled-completability, as characters with three of more such nodes are sufficient to discard this property.
+As established in Lemma 3, this serves as a first test for galled-completability, as finding three or more such nodes for a character is sufficient to discard the instance.
 Figure 10 shows the distribution of the number of first-appearance nodes for each character.
@@
 This prevents the common ancestor of the clade from being labeled with that character.
-It is plausible that one character was simply lost in one species, without invoking any transfer, as shown on the right.
+It is plausible that the character was transferred to the common ancestor of the clade, and then simply lost in one species, as shown on the right.
 This strongly indicates that the “transfer-only” model is too rigid, and in future work we plan to work on approaches to identify situations in which losses are more likely than transfers.
-Going back to Figure 9, we note that there does exist characters that posses at most two first-appearances.
+Unresolved species trees can obstruct galled-completability.
+Going back to Figure 10, we note that there do exist characters that possess at most two first-appearances.
 We checked whether this subset of four characters could at least be galled-completable, but it turns out that this is not even the case.

### v3

@@
 With the goal of recovering the simplest possible scenarios in this model, we introduce galled perfect transfer networks, which are PTNs that are galled trees.
-Such networks are useful for characters that are incompatible in terms of tree-like evolution, but that do fit in an almost-tree scenario.
+Such networks are useful for binary characters that are incompatible in terms of tree-like evolution, but that do fit in an almost-tree scenario.
 We provide polynomial-time algorithms for two problems: deciding whether one can add transfer edges to a tree to transform it into a galled PTN, and deciding whether a set of characters are galled-compatible, that is, they can be explained by some galled PTN.
@@
 introduced Perfect Phylogenetic Networks (PPNs) in [nakhleh, nakhlehtesis].
-These networks allow multi-state characters and require that, for each character $C$, the network contains a tree in which nodes in the same state are connected.
-This is a powerful model that is, unfortunately, difficult to work with, since even deciding whether a known network explains a set of characters is NP-hard.
+These networks allow multi-state characters and require that, for each character $C$, the network displays a tree in which nodes in the same state are connected.
+That is, it is possible to remove all but one of the incoming edges of each reticulation, then label the internal nodes of the resulting tree with a state, such that every state forms a connected component.
+This is a powerful model that is, unfortunately, difficult to work with, since even deciding whether a known network explains a set of characters is NP-hard [nakhlehtesis].
 We note that in parallel to our work, this question was studied in the context of binary characters on galled trees [warnow2024statistically], which we discuss in more detail below.
@@
 This implies knowledge of the vertical versus horizontal transmissions, and that these networks belong to the class of tree-based networks [Pons2018, Fischer2020] (in fact, they are LGT networks, see next section).
-In [ptn], we introduced perfect transfer networks (PTNs), a specialization of PPNs where characters satisfy the same assumptions as perfect phylogenies: they are binary, have a unique origin, and are assumed to always be inherited vertically once acquired.
+In [ptn], we introduced perfect transfer networks (PTNs), a specialization of PPNs where characters satisfy the same assumptions as perfect phylogenies: they are binary, have a unique origin, and characters can never be lost in the vertical descendants once they are acquired.
+The latter two conditions are often called the no-homoplasy and no-loss conditions, respectively.
 An important difference of PTNs is that one cannot choose which subtree of the network can be used to explain a character, as the tree of vertical inheritance is fixed.
-Because of this, as we showed in [ptn], even PPNs on binary characters are different from PTNs.
-In the latter, a character may or may not be transferred horizontally to other species.
 This can be useful for transferrable characters that are difficult to revert, such as material acquired horizontally from mitochondria or chloroplasts resulting from endosymbiotic events [Zachar2020, anselmetti2021gene], or in the case of metabolites when HGT plays a role in the generation of new metabolic pathways in bacteria [Goyal2022].
@@
 This raises the question of explaining a set of characters with the simplest possible network.
-One way would be to build a network or augment a tree with a minimum number of transfers, which is an unsolved complexity problem.
-Another direction is to impose structural conditions on the desired network.
+One way would be to build a network or augment a tree with a minimum number of transfers, or impose structural conditions on the desired network.
+Notably, there are several positive results on the reconstructibility of networks with bounded level that explain characters in the softwired sense, i.e., each character can be explained by switching off all but one incoming edges of reticulations [vanIersel2010, kelk2012].
+However, in PTNs, the no-loss condition implies that edges of vertical inheritance below a character’s origin cannot be switched off, and adapting the techniques developed in [vanIersel2010, kelk2012] to this additional restriction does not appear to be immediate (also see related works below).
 Our contribution.
-In this work, we focus on the second direction.
-We explore the evolutionary structure that many consider as the simplest beyond trees, namely galled trees.
-These consist of networks in which all underlying cycles are independent, and were first used in the context of hybridization and recombination [gusfield2004optimal].
+In this work, we explore the evolutionary structure that many consider as the simplest beyond trees, namely galled trees.
+These are also known as (binary) level-1 networks, and consist of networks in which all underlying cycles are independent, and were first used in the context of hybridization and recombination [gusfield2004optimal].
 When they fit the data, galled trees are desirable because of their parsimonious nature and ease of interpretation.
@@
 In the galled-completion problem, we ask whether it is possible to complete a given tree with transfers to obtain a galled PTN.
+We show that this verification can be done in polynomial time through the usage of an auxiliary structure.
 We also study the galled-compatibility problem in which, given a set of characters, we must decide whether a galled PTN can explain them.
+We show that it is possible to reconstruct a tree that can be augmented into a galled tree for a set of characters in polynomial time.
 During the process, we provide several structural characterizations of characters that can be displayed on a given tree or a network.
-Althought most of our theoretical results are intuitive, the detailed proofs are sometimes involved.
+Although most of our theoretical results are intuitive, the detailed proofs are sometimes involved.
 To improve the reading experience, we provide a sketch for most proofs, and refer to the appendix for the full details.
@@
 In a similar vein, in [vanIersel2010, kelk2012] the authors study the question of reconstructing a network that displays a set of characters in the softwired sense, meaning that for each character, some tree contained in the network contains it as a clade (characters are called clusters therein).
-In particular, one can reconstruct in polynomial time a level-1 network that explains a set of characters, if one exists.
-As level-1 networks are closely related to galled trees, it is likely that the latter can be used in our setting.
-But, as also argued in [ptn, Figure 2], softwired characters can explain different sets of characters than PTNs.
-The main reason is that distinguishing between horizontal and vertical edges does not allow choosing which tree of the network should be used to explain a character.
+It is known that for any fixed $k$, one can reconstruct in polynomial time a level-$k$ network that explains a set of characters, if one exists.
+In particular, level-1 networks are closely related to galled trees, so it is possible that these approaches can be used in our setting.
+But, as also argued in [ptn, Figure 2], softwired characters can explain more sets of characters than PTNs, mainly because of the no-loss condition.
 2 Preliminaries
 A phylogenetic network, or simply a network for short, is a directed acyclic graph $N=(V,E)$ with one node $\rho(N)$ of in-degree zero, called the root.
+We use $V(N)$ and $E(N)$ to denote the sets of nodes and edges of $N$, respectively.
+We say that a node $u\in V(N)$ reaches a node $v\in V(N)$ if there exists a directed path from $u$ to $v$ in $N$.
 A node of in-degree one and outdegree zero is a leaf, and $L(N)$ denotes the set of leaves of $N$.
@@
 A network $N$ is a galled tree if no two distinct underlying cycles of $N$ contain a common node (see Figure 1.(c)).
+Observe that two cycles may contain the same nodes but not the same edges, in which case they are considered distinct.
 A tree $T$ is a network with no underlying cycle.
@@
 Two nodes $u$,$v$ are incomparable in $T$ if none descends from the other, i.e., if neither $u\preceq v$ nor $v\preceq u$.
-The parent of a node $v$ in $T$ is $p_{T}(v)$.
+The parent of a non-root node $v$ in $T$ is $p_{T}(v)$.
+We shall only use the notions of $\preceq$, ancestors, descendants, and incomparable on trees, as they are not defined on networks.
 For $v\in V(T)$, we use $T(v)$ for the subtree of $T$ rooted at $v$, that is, $T(v)$ contains $v$ and all of its descendants.
@@
 We assume that for each transfer edge $(u,v)\in E_{T}$, the nodes $u$ and $v$ are incomparable in $\mathcal{T}_{N}$.
-We also assume that transfer nodes have exactly one out-neighbor in $\mathcal{T}_{N}$.
+We also assume that transfer nodes have exactly one child in $\mathcal{T}_{N}$.
 The tree obtained from $\mathcal{T}_{N}$ by suppressing its subdivision nodes is called the base tree of $N$ [^1] Suppressing a subdivision node $u$ with parent $p$ and child $v$ consists of removing $u$ and adding an edge from $p$ to $v$.
-For $v\in V(N)\setminus\{\rho(N)\}$, we use $p_{N}(v)$ to denote the unique in-neighbor of $v$ in the support tree of $N$.
+For $v\in V(N)\setminus\{\rho(N)\}$, we use the shorthand $p_{N}(v):=p_{\mathcal{T}_{N}}(v)$ to denote the parent of $v$ in the support tree of $N$.
 For simplicity, an LGT network that is also a galled tree will be called a galled LGT network.
+Let us emphasize that in an LGT network, the partition $\{E_{S},E_{T}\}$ is specified.
+That is, there may be multiple ways of defining a tree within $N$ along with transfer edges, but $\{E_{S},E_{T}\}$ gives the unique desired way to do so [^2] Notation-wise, it may be more accurate to define an LGT network as a triple $(V,E,f)$ where $f:E\rightarrow\{support,transfer\}$ specifies the type of each edge, but we prefer to use the more convenient partition notation as in [Cardona_2015]..
+Therefore, the support tree $\mathcal{T}_{N}$ is defined unambiguously.
+This contrasts with so-called tree-based networks, where a partition into support and transfer edges is known to exist, but is not specified and may not be unique.
+Also note that galled trees always admit such a partition and are thus tree-based, but calling them galled LGT networks clarifies that the partition into support and transfer edges is given.
 2.1 Perfect transfer networks
@@
 An example of this problem is shown on Figure 1.
-For a set of characters $\mathcal{C}$ on taxa $\mathcal{S}$, we say that $\mathcal{C}$ is galled-compatible if there exists a galled tree on leafset $\mathcal{S}$ that explains $\mathcal{C}$.
+For a set of characters $\mathcal{C}$ on taxa $\mathcal{S}$, we say that $\mathcal{C}$ is galled-compatible if there exists a galled LGT network on leafset $\mathcal{S}$ that explains $\mathcal{C}$.
 Our problems of interest are the following:
@@
 We begin by stating properties of galled PTNs that will be useful throughout.
-Let $N$ be a network that explains a set of characters $\mathcal{C}$.
-Observe that if a leaf $l$ does not possess a character $C\in\mathcal{C}$, then by the “never lost once acquired” condition, the parent of $l$ in $\mathcal{T}_{N}$ cannot possess $C$ either (otherwise, the parent would be required to transfer $C$ to $v$).
-In fact, no ancestor of $l$ in $\mathcal{T}_{N}$ can possess $C$.
+Let $N$ be an LGT network that explains a set of characters $\mathcal{C}$.
+Observe that if a leaf $s$ does not possess a character $C\in\mathcal{C}$, then by the “never lost once acquired” condition, the parent of $s$ in $\mathcal{T}_{N}$ cannot possess $C$ either (otherwise, the parent would be required to transmit $C$ vertically to $v$).
+In fact, no ancestor of $s$ in $\mathcal{T}_{N}$ can possess $C$.
 This lets us deduce a subset $F_{C}$ of nodes that forbid $C$, as follows:
-$F_{C}(N)=\{v\in V(N):\exists l\in L_{\mathcal{T}_{N}}(v)\mbox{ such that $l\notin C$}\}.$
+$F_{C}(N)=\{v\in V(N):\exists s\in L_{\mathcal{T}_{N}}(v)\mbox{ such that $s\notin C$}\}.$
+As an example, consider the network $N$ in Figure 1.(b).
+Here, $F_{C_{1}}(N)$ consists of every node from the root to the rightmost leaf (including the root and the leaf), because the character is not in that rightmost leaf.
+The set $F_{C_{2}}(N)$ contains all nodes on a path from the root to one of the two rightmost leaves, because those two do not have $C_{2}$ (so here, $F_{C_{2}}(N)$ includes one transfer node).
+Note that the same logic applies to subfigure (c).
 It follows that the origin of $C$ must be a node of $N-F_{C}$.
@@
 We also describe two useful generic properties of galled LGT networks.
-The first one says that a node $v$ cannot have two descending transfer nodes that each go outside of the $v$ subtree (otherwise, $v$ would be part of two distinct cycles).
+The first one says that a node $v$ cannot have two descending transfer nodes that each go outside of the descendants of $v$ in the support tree (otherwise, $v$ would be part of two distinct cycles).
 {lemmarep}
 Let $N=(V,E_{S}\cup E_{T})$ be a galled LGT network and let $v\in V(N)$ be a node with two distinct descendants $x$ and $y$ in $\mathcal{T}_{N}$ that are transfer nodes (with $v\in\{x,y\}$ being possible).
-Then one of the transfer edges incident to $x$ or $y$ has both endpoints in the subtree $\mathcal{T}_{N}(v)$.
+Then one of the transfer edges of $N$ incident to $x$ or $y$ has both endpoints in the subtree $\mathcal{T}_{N}(v)$.
 {proofsketch}
-If $v$ has two descendants $x,y$ with transfer neighbors that are external to the $v$ subtree of $\mathcal{T}_{N}$, then the two outgoing transfers inevitably create two distinct cycles that both contain $v$, contradicting the galled property.
+Suppose that $v$ has two distinct descendants $x,y$ in $\mathcal{T}_{N}$ that are both transfer nodes in $N$.
+Let $x^{\prime}$ (respectively $y^{\prime}$) be the other endpoint of the transfer edge of $N$ containing $x$ (resp.
+$y$), that is, $(x,x^{\prime})\in E_{T}$ or $(x^{\prime},x)\in E_{T}$ (same for $y$ and $y^{\prime}$).
+If both $x^{\prime}$ and $y^{\prime}$ are not descendants of $v$ in $\mathcal{T}_{N}$, then the two transfer edges containing $x$ and $y$ go outside of $\mathcal{T}_{N}(v)$ and create two cycles that both contain $v$, and these cycles are distinct because they use a different transfer edge, contradicting the galled property.
 ∎
@@
 By our supposition, in $\mathcal{T}_{N}$ we have that $v$ is on the path from $lca(x,x^{\prime})$ to $x$, and from $lca(y,y^{\prime})$ to $y$.
-This implies the existence of two underlying cycles in $N$: there is a cycle formed by $lca(x,x^{\prime})-x^{\prime}-x-lca(x,x^{\prime})$ and the cycle formed by $lca(y,y^{\prime})-y^{\prime}-y-lca(y,y^{\prime})$ (where the edges between and $x^{\prime},x$, and between $y^{\prime},y$ are used, and the rest use edges of $\mathcal{T}_{N}$).
+This implies the existence of two underlying cycles in $N$: there is a cycle formed by $lca(x,x^{\prime})-x-x^{\prime}-lca(x,x^{\prime})$ and the cycle formed by $lca(y,y^{\prime})-y-y^{\prime}-lca(y,y^{\prime})$ (where here, the dashes represent paths; the edges between and $x,x^{\prime}$, and between $y,y^{\prime}$ are used, and the rest use edges of $\mathcal{T}_{N}$).
 These cycles are distinct since they use a different (unique) transfer edge.
@@
 ∎∎
-The next property states that if some $x$ is able to reach a node $y$ that does not descend from it, then $x$ must of course achieve this through a transfer edge.
-{lemmarep}
+The next property states that if some $x$ is able to reach a node $y$ in $N$ but $y$ is not a vertical descendant of $x$, then $x$ must reach $y$ through a path that uses exactly one transfer edge.
+Lemma 2.
 Let $N=(V,E_{S}\cup E_{T})$ be a galled LGT network.
-Let $x,y\in V$ be two nodes that are not comparable in $\mathcal{T}_{N}$ and such that $x$ reaches $y$ in $N$.
-Then there exists an edge $(x^{\prime},y^{\prime})\in E_{T}$ such that $x^{\prime}\preceq_{\mathcal{T}_{N}}x$ and $y\preceq_{\mathcal{T}_{N}}y^{\prime}$.
-{proofsketch}
-Consider the first transfer edge $(x^{\prime},y^{\prime})$ on a path from $x$ to $y$ in $N$.
+Let $x,y\in V$ be two nodes that are incomparable in $\mathcal{T}_{N}$ and such that $x$ reaches $y$ in $N$.
+Then there exists a transfer edge $(x^{\prime},y^{\prime})\in E_{T}$ such that $x^{\prime}\preceq_{\mathcal{T}_{N}}x$ and $y\preceq_{\mathcal{T}_{N}}y^{\prime}$.
+Proof.
+Consider the first transfer edge $(x^{\prime},y^{\prime})$ on a path from $x$ to $y$ in $N$, which must exist.
 Note that $x^{\prime}$ must descend from $x$ in $\mathcal{T}_{N}$.
-If $y$ is not a descendant of $y^{\prime}$, then from $y^{\prime}$ the path needs to borrow another transfer edge that goes out of the $y^{\prime}$ subtree of $\mathcal{T}_{N}$.
-Thus, the $y^{\prime}$ subtree has two transfers with external endpoints, contradicting Lemma 2.2.
-∎
-Proof.
-Suppose for a contradiction that such $(x^{\prime},y^{\prime})$ does not exist in $N$.
-Since $x$ reaches $y$ in $N$, we know that there must exist at least another node $w\in V(N)$ incomparable to both $x$ and $y$ that serves as an intermediary node from $x$ to $y$.
-More specifically, there exists a transfer edge with an endpoint in $\mathcal{T}_{N}(x)$ and whose receiving end is $w$.
-Nevertheless, $y$ is not a descendant of $w$ (otherwise, $w=y^{\prime}$ would end the proof), which implies the existence of a transfer edge from some descendant of $w$ to a node incomparable with $w$.
-However this creates a situation where $\mathcal{T}_{N}(w)$ is a subtree with two transfer edges that have their endpoints outside of the subtree, which is forbidden by Lemma 2.2 ∎∎
+If $y$ is not a descendant of $y^{\prime}$ in $\mathcal{T}_{N}$, then from $y^{\prime}$ the path needs to borrow another transfer edge that goes out of the $y^{\prime}$ subtree of $\mathcal{T}_{N}$.
+Thus $y^{\prime}$ has two distinct descendants in $\mathcal{T}_{N}(y^{\prime})$ that are transfer nodes with external endpoints, namely itself and the next transfer node on the path, contradicting Lemma 2.2.
+∎∎
 2.3 Differences with reconstructions from bipartitions
 As we pointed out in the introduction, Warnow et al.
-[warnow2024statistically] also independently developed galled tree reconstructions algorithms from character data (which are assumed to be SNPs in the paper).
+[warnow2024statistically] also independently developed galled tree reconstruction algorithms from character data (which are assumed to be SNPs in the paper).
+With the above definitions in mind, we can now clearly state how PTNs differ from this work.
 Therein, for each character, the species can be in one of two states, and the states are assumed to correspond to a bipartition of a galled tree $N$.
@@
 The set of FA nodes for $C$ in $T$ is denoted as $\alpha_{T}(C)$.
-In other words, $v$ is a FA for $C$ if it roots a maximal subtree of taxa that contain $C$. An example of this definition is shown in Figure 3.
+In other words, $v$ is an FA for $C$ if it roots a maximal subtree of taxa that contain $C$. An example of this definition is shown in Figure 3.
 Figure 3: A tree $T$ on species $\mathcal{S}=\{s_{1},s_{2},s_{3},s_{4},s_{5},s_{6},s_{7}\}$ with character set $\mathcal{C}=\{C_{1},C_{2},C_{3},C_{4}\}$.
@@
 The corresponding sets of FAs for every character are as follows: $\alpha_{T}(C_{1})=\{x,y\},\alpha_{T}(C_{2})=\{y,s_{2}\},\alpha_{T}(C_{3})=\{v\},\alpha_{T}(C_{4})=\{u\}$.
-For a character $C\in\mathcal{C}$ and a network $N$, we will say that a node $v$ is an origin for $C$ if $v$ reaches every leaf in $C$ in $N-F_{C}(N)$.
-By Lemma 1, our goal is to add arcs to $T$ to ensure that every character has an origin.
+For a character $C\in\mathcal{C}$ and an LGT network $N$, we say that a node $v$ is an origin for $C$ if $v$ reaches every leaf in $C$ in $N-F_{C}(N)$.
+By Lemma 1, our goal is to add transfer edges to $T$ to ensure that every character has an origin.
 We first show that in any galled-completion of a tree $T$, an origin of a character $C$ must descend from a FA node (or it could be a transfer node added just above it), and it must “give” its character to all other FAs by transferring just above them.
@@
 {proofsketch}
-By definition, in $T$ the ancestors of FAs are in $F_{C}(N)$ since they have descendants not in $C$.
-Thus, an origin $w$ in $N$ must be in a FA subtree, or it could be a transfer node inserted above such an FA during the process of adding transfer edges from $T$ to $N$.
-If there are multiple FAs, $w$ must be able to reach the other ones, and Lemma 2.2 lets us establish that transfer edges as described are needed to achieve this.
+First note that $V(T)\subseteq V(N)$, by the definition of a completion.
+For each FA $\alpha\in V(T)$ of character $C$, and each strict ancestor $z$ of $\alpha$ in $T$, i.e., with $\alpha\prec_{T}z$, we have that $L_{T}(z)$ has leaves not containing character $C$.
+The same is true for $L_{\mathcal{T}_{N}}(z)$, and so $z$ is in $F_{C}(N)$.
+This means that in $N$, an origin $w$ for $C$ cannot be $p_{T}(v)$ nor an ancestor of $p_{T}(v)$ in $\mathcal{T}_{N}$.
+So, $w$ must either descend from an FA node $\alpha_{i}$ in $\mathcal{T}_{N}$, or $w$ could be a transfer node inserted above $\alpha_{i}$ while adding transfer edges from $T$ to $N$.
+For example in Figure 3, in a galled-completion $N$ of $T$, an origin of $C_{2}$ could never be $x$ nor any of its ancestors in $\mathcal{T}_{N}$, because those nodes have $s_{1}$ as a vertical descendant (likewise, $p_{T}(y)$ and its ancestors cannot be origins because of $s_{7}$).
+This justifies the first part of the statement.
+For the second part, if there are multiple FAs, $w$ must reach the ones other than $\alpha_{i}$.
+Using the definition of FAs, we can deduce that $w$ is incomparable to all nodes of $\alpha_{T}(C)\setminus\{\alpha_{i}\}$ in $\mathcal{T}_{N}$, and Lemma 2 lets us establish that transfer edges as described are needed to achieve this.
 ∎
@@
 We focus on the first condition.
-First observe that if $v\in V(T)$, then $v\in V(N)$ as well.
-Moreover, the the parent of $v$ in $\mathcal{T}_{N}$ is either also the parent $p_{v}$ of $v$ in $T$, or it could be a transfer node created to complete $T$ into $N$.
-In the latter case, note that at most one transfer node could could be present on the path between $p_{v}$ and $v$ in $\mathcal{T}_{N}$, as otherwise, there would be two consecutive transfer nodes which, as one can check, would violate Lemma 2.2 ($p_{v}$ would have two descendants with outgoing transfers).
+Before proceeding, notice that if $N=(V,E_{S}\cup E_{T})$ is a galled-completion of $T$, then $\mathcal{T}_{N}$ is obtained from $T$ by subdividing some edges of $T$ to add transfer nodes, possibly multiple times on the same edge.
+We claim that $\mathcal{T}_{N}$ can be obtained from $T$ by subdividing each edge at most one time.
+In other words, no two consecutive transfer nodes can be added along an edge of $T$.
+To see this, let $v\in V(T)$ be a non-root node of $T$ and let $p_{v}=p_{T}(v)$ be the parent of $v$ in $T$.
+Notice that $v,p_{v}\in V(N)$.
+Suppose for contradiction that from $T$ to $\mathcal{T}_{N}$, the edge $(p_{v},v)$ is subdivided twice or more.
+That is, suppose that in $N$, there are support edges $(p_{v},x),(x,y)\in E_{S}$, such that $x\neq v$ and $y\neq v$ (and $v\prec_{\mathcal{T}_{N}}y$).
+Then $x$ and $y$ are both transfer nodes, and are both incident to a distinct transfer edge whose other endpoint is outside of the subtree $\mathcal{T}_{N}(x)$.
+This contradicts Lemma 2.2 (notice that $x$ and $y$ have a single child in $\mathcal{T}_{N}$ and that the lemma still applies to this case).
 This lets us establish that some edges of $T$ are subdivided to obtain $\mathcal{T}_{N}$, but each edge can be subdivided at most once.
 Let us next observe that if $w$ is an origin for $C$, by definition it must be a node in $N-F_{C}(N)$, which is a graph that only contains FA nodes of $T$, their support tree descendants in $N$, and possibly transfer nodes that are the parents of FA nodes.
-Thus either $w$ descends from some $\alpha_{i}\in\alpha_{T}(C)$, or results from a subdivision of $(p_{T}(\alpha_{i}),\alpha_{i})$.
-In the latter case, $w$ is the parent of $\alpha_{i}$ since an edge can only be subdivided once, as explained above.
+Thus either $w$ descends from some $\alpha_{i}\in\alpha_{T}(C)$ in $\mathcal{T}_{N}$, or results from a subdivision of $(p_{T}(\alpha_{i}),\alpha_{i})$.
+In the latter case, $w$ is the parent of $\alpha_{i}$ in $\mathcal{T}_{N}$ since an edge can only be subdivided once, as explained above.
 Therefore, the first condition of our statement holds.
-For the second condition, as we now know the origin $w$ descends from some $\alpha_{i}$ or is its parent.
-Let $\alpha_{j}$ be a FA other than $\alpha_{i}$.
-Let $l$ be a leaf descending from $\alpha_{j}$ in $\mathcal{T}_{N}$.
-Since $w$ is an origin, it reaches $l$ in $N-F_{C}(N)$ and thus also in $N$.
-Since first-appearance nodes are incomparable in $T$ and $\mathcal{T}_{N}$, $w$ is not an ancestor of $l$ in $\mathcal{T}_{N}$.
-By Lemma 2.2 there is a transfer edge $(u,v)$ in $N$ such that $u\preceq w$ and $l\preceq v$ in $\mathcal{T}_{N}$.
+For the second condition, as we now know the origin $w$ descends from some $\alpha_{i}$ or is its parent in $\mathcal{T}_{N}$.
+Let $\alpha_{j}$ be an FA other than $\alpha_{i}$.
+Let $x$ be a leaf descending from $\alpha_{j}$ in $\mathcal{T}_{N}$.
+Since $w$ is an origin, it reaches $x$ in $N-F_{C}(N)$ and thus also in $N$.
+Since first-appearance nodes are incomparable in $T$ and $\mathcal{T}_{N}$, $w$ is not an ancestor of $x$ in $\mathcal{T}_{N}$.
+By Lemma 2 there is a transfer edge $(u,v)$ in $N$ such that $u\preceq w$ and $x\preceq v$ in $\mathcal{T}_{N}$.
 The endpoints of this transfer edge must also be in $N-F_{C}(N)$ for $w$ to be able to use it.
-Next, note that $p_{T}(\alpha_{j})$ has descendants not in $C$.
-Thus in $N$, $p_{T}(\alpha_{j})$ and its support tree ancestors are in $F_{C}(N)$, and so it follows that $v$ descends from $\alpha_{j}$ or is a transfer node on its parent edge, in which case $v=p_{N}(\alpha_{j})$ (again since edges can only be subdivided once).
+Next, note that $p_{T}(\alpha_{j})$ has descendants not in $C$, both in $T$ and in $\mathcal{T}_{N}$.
+Thus in $N$, $p_{T}(\alpha_{j})$ and its support tree ancestors are in $F_{C}(N)$, and $v\in V(N-F_{C}(N))$ implies that either $v\prec_{\mathcal{T}_{N}}\alpha_{j}$ or $v$ is a transfer node above $\alpha_{j}$, i.e., $v=p_{N}(\alpha_{j})$ (again since edges can only be subdivided once).
 If $\alpha_{j}$ is a leaf, the latter must hold and we are done.
-So assume that $\alpha_{j}$ is not a leaf and let $l_{1},l_{2}$ be two leaves descending from $\alpha_{j}$ in $\mathcal{T}_{N}$.
-Then there are two transfer edges $(u_{1},v_{1})$ and $(u_{2},v_{2})$ that exist in $N-F_{C}(N)$, where $u_{1},u_{2}$ descend from $w$ and $v_{1},v_{2}$ are ancestors of $l_{1}$ and $l_{2}$, respectively.
+So assume that $\alpha_{j}$ is not a leaf and let $x_{1},x_{2}$ be two leaves descending from $\alpha_{j}$ in $\mathcal{T}_{N}$.
+Then there are two transfer edges $(u_{1},v_{1})$ and $(u_{2},v_{2})$ that exist in $N-F_{C}(N)$, where $u_{1},u_{2}$ descend from $w$ in $\mathcal{T}_{N}$ and $v_{1},v_{2}$ are ancestors of $x_{1}$ and $x_{2}$ in $\mathcal{T}_{N}$, respectively.
 Since these two transfers go out of the $\mathcal{T}_{N}(w)$ subtree, they must in fact be equal by Lemma 2.2.
@@
 {lemmarep}
-Let $T$ be a galled-completable tree for a character set $\mathcal{C}$. Then for any character $C\in\mathcal{C}$, ${\alpha_{T}(C)\leq 2}$.
+Let $T$ be a galled-completable tree for a character set $\mathcal{C}$. Then for any character $C\in\mathcal{C}$, ${|\alpha_{T}(C)|\leq 2}$.
 {proofsketch}
@@
 Let $N$ be a galled-completion of $T$ that explains $\mathcal{C}$.
-Suppose for contradiction that $\alpha_{T}(C)\geq 3$ for some $C\in\mathcal{C}$.
+Suppose for contradiction that $|\alpha_{T}(C)|\geq 3$ for some $C\in\mathcal{C}$.
 Let $w$ be an origin for $C$ in $N-F_{C}(N)$.
-By Lemma 3, $w$ descends from or is the parent of one of the FAs, say $x\in\alpha_{T}(C)$.
+By Lemma 3, there is $x\in\alpha_{T}(C)$ such that $w\preceq_{\mathcal{T}_{N}}x$ or $w=p_{N}(x)$.
 Let $y,z$ be two other FAs of $C$.
-Note that, by definition of FAs, $x,y,z$ are pairwise incomparable in $T$.
-By the second statement of Lemma 3, there are transfer edges $(u_{1},p_{N}(y))$ and $(u_{2},p_{N}(z))$ where $u_{1}$ and $u_{2}$ descend from $w$ in $\mathcal{T}_{N}$.
+Note that, by definition of FAs, $x,y,z$ are pairwise incomparable in $T$, and thus in $\mathcal{T}_{N}$.
+By the second statement of Lemma 3, there are transfer edges $(u_{1},p_{N}(y))$ and $(u_{2},p_{N}(z))$ in $N$ where $u_{1}$ and $u_{2}$ descend from $w$ in $\mathcal{T}_{N}$.
 This implies that the subtree $\mathcal{T}_{N}(w)$ contains two outgoing transfers, which contradicts Lemma 2.2.
@@
   1. 1.
-    If $b_{1}\prec a_{1}$ and $b_{2}$ is not comparable to $a_{1}$, then $a_{2}=b_{2}$.
+    If $b_{1}\prec_{T}a_{1}$ and $b_{2}$ is not comparable to $a_{1}$ in $T$, then $a_{2}=b_{2}$.
   2. 2.
-    If $a_{1}=b_{1}$, then either $b_{2}\prec a_{2}$ or $a_{2}\prec b_{2}$.
+    If $a_{1}=b_{1}$, then either $b_{2}\prec_{T}a_{2}$ or $a_{2}\prec_{T}b_{2}$.
 {proofsketch}
-For the first statement, if $b_{1}\prec a_{1}$ and $b_{2}$ is outside of the $a_{1}$ subtree, then $b_{1}$ needs to use the same transfer edge as $a_{1}$ to exchange material (because of the galled property).
-This is only possible if they send the character to the same clade, that is, if $a_{2}=b_{2}$.
+For the first statement, by Lemma 3 in any galled-completion $N$ of $T$ there is a transfer edge with one end that is either $p_{N}(a_{1})$ or descends from $a_{1}$ in $\mathcal{T}_{N}$, and the other end is outside of $\mathcal{T}_{N}(a_{1})$.
+Likewise, some transfer edge has one end that is $p_{N}(b_{1})$ or descends from $b_{1}$, and the other end is $p_{N}(b_{2})$ or it descends from $b_{2}$.
+That other end is also outside of $\mathcal{T}_{N}(a_{1})$, since $a_{1}$ and $b_{2}$ are incomparable (in both $T$ and $\mathcal{T}_{N}$) by assumption.
+Hence by Lemma 2.2 the two transfer nodes must be equal, i.e., $A$ and $B$ use the same transfer edge in $N$ to transfer the character.
+With some case analysis, we can show that this is only possible if the origins of both $A$ and $B$ are on the $a_{1}$ and $b_{1}$ side, and they send their character to $a_{2}$ and $b_{2}$ using that transfer edge.
+The second part of Lemma 3 lets us deduce that the receiving end must be $p_{N}(a_{2})=p_{N}(b_{2})$ and that $a_{2}=b_{2}$.
 For the second statement, if $a_{1}=b_{1}$, then again the two characters must use the same transfer edge to exchange material.
-In fact, one can argue that $a_{1}=b_{1}$ must be the receiving end of the transfer, and that $a_{2},b_{2}$ must be comparable to be able to use the same node to send the character.
+In fact, one can argue that $p_{N}(a_{1})=p_{N}(b_{1})$ must be the receiving end of the transfer, and that $a_{2},b_{2}$ must be comparable to be able to use the same transfer edge to send the character.
 ∎
@@
   • Case 1: $a_{2}$ and $b_{2}$ are comparable.
-Suppose $b_{2}\prec a_{2}$.
+Suppose $b_{2}\prec_{T}a_{2}$.
 Let $w$ be an origin for $A$ in $N-F_{A}(N)$.
-By Lemma 3, $w$ descends from $a_{1}$ or $a_{2}$, or is a transfer node with one of those as a child.
-Suppose first that $w\preceq a_{1}$ or $w=p_{N}(a_{1})$.
-We know by Lemma 3 that there exists a transfer edge from some descendant of $w$ that has $p_{N}(a_{2})$ as endpoint.
-Also by Lemma 3, $p_{N}(b_{2})$ or a descendant of $b_{2}$ is a transfer node whose other endpoint is $p_{N}(b_{1})$ or a descendant of $b_{1}$.
-This implies that there are two transfer nodes in the subtree $\mathcal{T}_{N}(p_{N}(a_{2}))$ whose other endpoint is outside, which is forbidden by Lemma 2.2.
-The case where $w\preceq a_{2}$ or is the parent of $a_{2}$ is symmetric.
-    Next suppose that $a_{2}\prec b_{2}$.
+By Lemma 3, $w$ descends from $a_{1}$ or $a_{2}$ in $\mathcal{T}_{N}$, or is a transfer node with one of those as a child.
+Suppose first that $w\preceq_{\mathcal{T}_{N}}a_{1}$ or $w=p_{N}(a_{1})$.
+We know by Lemma 3 that there exists a transfer edge from some descendant of $w$ in $\mathcal{T}_{N}$ that has $p_{N}(a_{2})$ as endpoint.
+Also by Lemma 3, $p_{N}(b_{2})$ or a descendant of $b_{2}$ in $\mathcal{T}_{N}$ is a transfer node whose other endpoint is $p_{N}(b_{1})$ or a descendant of $b_{1}$.
+But $b_{2}\prec_{\mathcal{T}_{N}}a_{2}$ implies that these two transfer edges are distinct, and so there are two transfer nodes in the subtree $\mathcal{T}_{N}(p_{N}(a_{2}))$ whose other endpoint is outside, which is forbidden by Lemma 2.2.
+The case where $w\preceq_{\mathcal{T}_{N}}a_{2}$ or is the parent of $a_{2}$ is symmetric.
+    Next suppose that $a_{2}\prec_{\mathcal{T}_{N}}b_{2}$.
 By Lemma 3, one of $p_{N}(a_{1})$ or $p_{N}(a_{2})$ is the receiving end of a transfer, whose sending end is on the other side.
-If $p_{N}(a_{1})$ is a transfer node, we also know that $p_{N}(b_{1})$ or one of its descendants is a transfer node, with the other endpoint on the $b_{2}$ side.
-Either way, $\mathcal{T}_{N}(p_{N}(a_{1}))$ has two descending transfer nodes with external endpoints, contradicting Lemma 3.
+If $p_{N}(a_{1})$ is a transfer node, we also know that $p_{N}(b_{1})$ or one of its descendants in $\mathcal{T}_{N}$ is a transfer node, with the other endpoint on the $b_{2}$ side.
+Therefore, $\mathcal{T}_{N}(p_{N}(a_{1}))$ has two descending transfer nodes with external endpoints, contradicting Lemma 3.
 Thus $p_{N}(a_{2})$ is the receiving end of a transfer.
 By a symmetric argument, $p_{N}(b_{1})$ is the receiving end of a transfer.
-This implies that there are transfers between $\mathcal{T}_{N}(a_{1})$ and $\mathcal{T}_{N}(b_{2})$ (or their parents) going in opposite directions.
+This implies that there are transfers edges between nodes of $\mathcal{T}_{N}(a_{1})$ and nodes of $\mathcal{T}_{N}(b_{2})$ (or their parents) going in opposite directions.
 Because transfer edges are unidirectional, these transfers are distinct, from which it can easily be seen that there are two underlying cycles intersecting.
@@
   • Case 2: $a_{2}$ and $b_{2}$ are incomparable.
-By Lemma 3, there is one transfer with one endpoint being $p_{N}(a_{1})$ or a descendant of $a_{1}$, and the other being $p_{N}(a_{2})$ or a descendant of $a_{2}$.
-Likewise, there is a similar transfer connecting $b_{1}$ and $b_{2}$.
-Since $a_{2}$ and $b_{2}$ are incomparable, these two transfers must be distinct, they are both in $\mathcal{T}_{N}(a_{1})$ or $\mathcal{T}_{N}(p_{N}(a_{1}))$, and they both have external endpoints since $b_{2}$ is not comparable with $a_{1}$, contradicting Lemma 2.2.
+By Lemma 3, there is one transfer edge with one endpoint being $p_{N}(a_{1})$ or a descendant of $a_{1}$ in $\mathcal{T}_{N}$, and the other being $p_{N}(a_{2})$ or a descendant of $a_{2}$ in $\mathcal{T}_{N}$.
+Likewise, there is a similar transfer edge connecting $b_{1}$ and $b_{2}$.
+Since $a_{2}$ and $b_{2}$ are incomparable, these two transfers edges must be distinct.
+Because $b_{1}\prec_{\mathcal{T}_{N}}a_{1}$, they are both in $\mathcal{T}_{N}(a_{1})$ or $\mathcal{T}_{N}(p_{N}(a_{1}))$, and they both have an outside endpoint since $b_{2}$ is not comparable with $a_{1}$, contradicting Lemma 2.2.
 Since neither case is possible, we must have $a_{2}=b_{2}$.
@@
 Let $N$ be a galled-completion of $T$ that explains $\mathcal{C}$.
-Suppose for a contradiction that $a_{1}=b_{1}$, but neither $a_{2}\prec b_{2}$ nor $b_{2}\prec a_{2}$ holds i.e.
-$a_{2}$ and $b_{2}$ are incomparable in $\mathcal{T}_{N}$.
-By Lemma 3, there is a transfer edge between descendants of $a_{1}$ and $a_{2}$ (or their parents), and a transfer edge between descendants of $b_{1}$ and $b_{2}$ (or their parents).
+Suppose for a contradiction that $a_{1}=b_{1}$, but neither $a_{2}\prec_{T}b_{2}$ nor $b_{2}\prec_{T}a_{2}$ holds, i.e., $a_{2}$ and $b_{2}$ are incomparable in $\mathcal{T}_{N}$.
+By Lemma 3, there is a transfer edge between descendants of $a_{1}$ and $a_{2}$ (or their parents) in $\mathcal{T}_{N}$, and a transfer edge between descendants of $b_{1}$ and $b_{2}$ (or their parents) in $\mathcal{T}_{N}$.
 Since $a_{2}$ and $b_{2}$ are incomparable, these transfers must be distinct.
-However, they imply the existence of two transfer nodes descending from $a_{1}=b_{1}$ (or its parent) that have an external endpoint, contradicting Lemma 2.2.
+However, they imply the existence of two transfer nodes descending from $a_{1}=b_{1}$ (or its parent) in $\mathcal{T}_{N}$ that have an external endpoint, contradicting Lemma 2.2.
 ∎∎
-Note that in the detailed proof of the first statement of Lemma 3, we use the fact that transfer edges are unidirectional — the statement in the sketch that $b_{1}$ needs to use the “same” transfer as $a_{1}$ remains true, but they do not need to send the character to the same clade in a bidirectional setting, as the sending could go both ways.
+Note that in the detailed proof of the first statement of Lemma 3, we use the fact that transfer edges are unidirectional — the statement in the sketch that $A$ and $B$ need to use the “same” transfer edge remains true, but they do not need to send the character in the same direction in a bidirectional setting, as the sending could go both ways.
 If we allowed bidirectional transfer edges, we can devise examples in which the lemma does not hold.
 This shows that even apparently minor changes to the model can lead to more complex structures.
+An algorithm using redundancy-free networks
 We can begin describing our algorithmic strategy.
 The first step is to locate and count the FAs for each character to verify the condition established by Lemma 3.
-A subtree which is rooted at a FA for a specific character $C$ is in fact a maximal clade for $C$.
+A subtree which is rooted at an FA for a specific character $C$ is in fact a maximal clade for $C$.
 An intuitive way of joining these clades is to add transfer edges between the different subtrees to fulfill the connectivity requirement in Lemma 1.
@@
 A pair of FA nodes $\{v,w\}$ is called simple if $w$ is the unique FA neighbor of $v$ and $v$ is the unique FA neighbor of $w$.
+See Figure 4.b for an example, where FA neighborhood relationships are shown in dotted lines.
+It turns out that these relationships tell us where transfer edges should be added.
 Definition 3.
@@
 Let $T^{\prime}$ be the tree obtained by subdividing every edge of $T$ once.
-Then the redundancy-free network for $T$ is the LGT-network $N=(V,E_{S}\cup E_{T})$ with $T^{\prime}$ as support tree obtained as follows:
-  • for each $v\in V(T)$ with at least two FA neighbors, and for every minimal FA neighbor $w$ of $v$, add the transfer edge $(p_{T^{\prime}}(w),p_{T^{\prime}}(v))$;
-  • for each pair $\{v,w\}$ of simple FA nodes, add one of the transfer edges $(p_{T^{\prime}}(v),p_{T^{\prime}}(w))$ or $(p_{T^{\prime}}(w),p_{T^{\prime}}(v))$ arbitrarily (but not both).
+Then a redundancy-free network for $T$ is an LGT-network $N=(V,E_{S}\cup E_{T})$ with $T^{\prime}$ as support tree obtained as follows:
+  • for each $v\in V(T)$ with at least two FA neighbors, and for every minimal FA neighbor $w$ of $v$, add the transfer edge $(p_{T^{\prime}}(w),p_{T^{\prime}}(v))$ to $E_{T}$;
+  • for each pair $\{v,w\}$ of simple FA nodes, add to $E_{T}$ one of the transfer edges $(p_{T^{\prime}}(v),p_{T^{\prime}}(w))$ or $(p_{T^{\prime}}(w),p_{T^{\prime}}(v))$ arbitrarily (but not both).
+Note that there may be multiple redundancy-free networks $N$ for a tree $T$, since the direction of transfer edges added in the second step is arbitrary.
+Let us also point out that if $w$ is an FA neighbor of $v$, then $v,w$ are FAs of some character $C$ and are therefore incomparable.
+Hence, edges of $N$ are only between incomparable nodes.
 Figure 4: (a) A given tree $T$ on four characters.
 (b) The FA neighbors between every FAs for each character are indicated by the dashed lines.
-(c) The redundancy free network of $T$.
+In this example, $y$ is an FA for $C_{1}$ and $C_{2}$ and has two FA neighbors $x$ and $z$ ($x$ is an FA for $C_{1}$ and $z$ an FA for $C_{2}$).
+Here $z$ is a minimal FA neighbor of $y$ because no strict descendant of $z$ is an FA neighbor of $y$, but $x$ is not a minimal FA neighbor of $y$ (because $z\prec_{T}x$).
+(c) A redundancy free network of $T$.
 Note that $C_{3}$ has simple FAs, so the direction for the transfer edge is arbitrary.
 On the other hand, for $C_{1}$ and $C_{2}$ note that $z$ is the minimal FA neighbor of $y$ so the direction of the transfer edge is not arbitrary.
-The idea of this network is that if a node $y$ has multiple FA neighbors, as in Figure 4, then there are many characters with $y$ as a FA.
-By adding a transfer edge whose sender is the parent of the minimal FA neighbor $z$ of $y$, again as in the figure, all those characters can have their origin on the $z$ side and transfer to $y$.
-In the figure, we add a transfer from the parent of $z$ to that of $y$ for character $C_{2}$.
-At the same time, $C_{1}$ can use that same transfer edge because it also wants $y$ and its descendants to inherit the character.
-For characters whose FAs correspond to a simple pair, no such transfer re-use is needed and the direction is arbitrary.
-Before proving our main characterization in terms of redundancy-free networks, we need an intermediary result.
+To illustrate the idea of a redundancy-free network and the importance of minimal FA neighbors, consider the example provided in Figure 4.
+Therein, node $y$ is an FA for two characters $C_{1},C_{2}$ and has two FA neighbors $x$ and $z$, which are respectively FAs of $C_{1}$ and $C_{2}$.
+Here, $z$ is a minimal FA neighbor of $y$ but not $x$, as seen on subfigure (b).
+By Lemma 3, in any galled-completion $N$ of $T$, because of $C_{1}$ either $p_{N}(x)$ or a descendant of $x$ in $\mathcal{T}_{N}$ is a transfer node, and because of $C_{2}$ we must have that $p_{N}(z)$ is a transfer node (because in the example $z$ is a leaf and cannot have descending transfer nodes).
+By the same lemma, both these transfer nodes are incident to an edge whose other end is $p_{N}(y)$ or a descendant of $y$, and to satisfy Lemma 2.2 these two transfer edges must be equal (otherwise we create intersecting cycles).
+In other words, $C_{1}$ and $C_{2}$ must use the same transfer edge to connect their FAs, and one end of that transfer edge must be the parent of $z$, the minimal FA neighbor of $y$.
+One can then work out that the edge must be from $p_{N}(z)$ to $p_{N}(y)$, since otherwise $x$ has no way to receive the character $C_{1}$ from a transfer to transmit it vertically.
+This is precisely the edge that we add to the redundancy-free network.
+This idea generalizes as follows.
+Suppose that some node $y$ of $T$ has multiple FA neighbors $x_{1}\prec_{T}x_{2}\prec_{T}\ldots\prec_{T}x_{k}$, where $x_{1}$ is a minimal FA neighbor of $y$, because of some character $C$.
+As argued above, for each $i\in[k]$ some transfer edge is needed between $p_{N}(x_{i})$ or a descendant of $x_{i}$, and $p_{N}(y)$ or a descendant of $y$.
+Still by Lemma 2.2, all these transfer edges must be equal, and one end of that transfer must be $p_{N}(x_{1})$ or a descendant of $x_{1}$.
+Hence, the minimal FA neighbor $x_{1}$ of $y$ gives the “highest” possible location at which a transfer node can be added.
+Then, the direction from the $x_{1}$ side to the $y$ side is forced since all the $x_{i}$’s must use that transfer edge to send their character to all descendants of $y$.
+We add such an edge in the construction of a redundancy-free network.
+Before proving our main characterization in terms of redundancy-free networks, we need an intermediary result, which essentially states that if $T$ satisfies all the established properties so far, then each node of $N$ needs to send its characters to at most one other node.
+In particular, we never add bidirectional edges, that is, for two nodes $u,v$ at most one of $(u,v)$ or $(v,u)$ is added.
 {lemmarep}
 Let $T$ be a tree on leafset $\mathcal{S}$ and character set $\mathcal{C}$.
-If $T$ is galled-completable, then in the redundancy-free network $N=(V_{T},E_{T})$ of $T$, every node is incident to at most one transfer edge.
+If $T$ is galled-completable, then in a redundancy-free network $N=(V_{T},E_{S}\cup E_{T})$ of $T$, every node is incident to at most one transfer edge.
 {proofsketch}
-Suppose that some node $u^{\prime}$ of $N$ is incident to two distinct transfer edges, with $v^{\prime}$ and $w^{\prime}$ as the other endpoints (where $u^{\prime},v^{\prime},w^{\prime}$ are subdivision nodes added when transforming $T$ to $N$).
-Since $N$ only adds transfers between minimal FA neighbors, this means that $v^{\prime}$ and $w^{\prime}$ are incomparable nodes of $\mathcal{T}_{N}$ (a special case arises when $v^{\prime}=w^{\prime}$, see the full proof for details).
-Roughly speaking, this means that two characters need to exchange material between the clades of $u^{\prime}$ and $v^{\prime}$, and also between the clades of $u^{\prime}$ and $w^{\prime}$.
-This is only possible in a galled-completion if the $u^{\prime}$ subtree contains at least two transfers, one exchanging with $v^{\prime}$ and the other with $w^{\prime}$.
-This creates two transfers below $u^{\prime}$ with external endpoints, contradicting the galled property by Lemma 2.2.
+Suppose that some node $u^{\prime}$ of $N$ is incident to two distinct transfer edges, with $v^{\prime}$ and $w^{\prime}$ as the other endpoints.
+Here $u^{\prime},v^{\prime},w^{\prime}$ are subdivision nodes added when transforming $T$ to $N$, whose children in $\mathcal{T}_{N}$ are $u,v,w\in V(T)$, respectively.
+The presence of the two transfer edges implies that $\{u,v\}$ and $\{u,w\}$ are both sets of FAs of distinct characters.
+Moreover, since we only add edges between minimal FA neighbors, we get that $v$ and $w$ must be incomparable, as otherwise one of them would not be a minimal FA neighbor of $u$ (this requires case checking depending on the direction of the edges, along with a special case that arises when $v^{\prime}=w^{\prime}$, see the full proof for details).
+The pairs $\{u,v\}$ and $\{u,w\}$ of FAs then contradict Lemma 3, second part.
 ∎
@@
 We can further assume that $w$ is a minimal FA neighbor of $u$ (otherwise, we choose $w$ to be such a neighbor, which does not affect the presence of the edge between $u^{\prime}$ and $v^{\prime}$), which guarantees that $(w^{\prime},u^{\prime})\in E_{T}$.
-By the construction of $N$, the edge $(v^{\prime},u^{\prime})$ cannot be in $E_{T}$, as $u^{\prime}$ only receives transfers from its minimal neighbors.
+By the construction of $N$, the edge $(v^{\prime},u^{\prime})$ cannot be in $E_{T}$, as $u^{\prime}$ only receives transfers from the parents of the minimal FA neighbors of $u$.
 Therefore, the transfer between $u^{\prime}$ and $v^{\prime}$ is in the $(u^{\prime},v^{\prime})$ direction.
@@
 This means that there is another character $C_{vz}$ with FAs $\{v,z\}$, where $z\neq u$.
-Using Lemma 3 in the same way as before, we get $u\prec_{T}z$.
+Using Lemma 3 in the same way as before, we get $u\prec_{T}z$ (and not $z\prec_{T}u$ because $u$ is a minimal FA neighbor of $v$).
 We now have two pairs of FAs $\{z,v\}$ and $\{u,w\}$, where $u\prec_{T}z$, $w$ is not comparable to $z$ (because $v$ is not), and $w\neq v$.
 By putting $\{z,v\}=\{a_{1},a_{2}\}$ and $\{u,w\}=\{b_{1},b_{2}\}$, we obtain a contradiction of the first part of Lemma 3.
-So suppose that $v=w$, and hence $C_{uv}=C_{uw}$.
+So suppose that $v=w$.
+Recall that $u,v$ are the FAs of character $C_{uv}$, and by definition, this means that $C_{uv}=L(T(u))\cup L(T(v))$.
+Likewise, $C_{uw}=L(T(u))\cup L(T(w))$, and because $v=w$ we deduce that $C_{uv}=C_{uw}$ (also recall that we deal with sets of character, no so character is repeated).
 This means that both edges $(u^{\prime},v^{\prime})$ and $(v^{\prime},u^{\prime})$ are present.
@@
 {lemmarep}
-A tree $T$ on leafset $\mathcal{S}$ with character set $\mathcal{C}$ is galled-completable if and only if its redundancy-free network $N=(V,E_{S}\cup E_{T})$ is a galled tree.
+A tree $T$ on leafset $\mathcal{S}$ with character set $\mathcal{C}$ is galled-completable if and only if any of its redundancy-free networks $N=(V,E_{S}\cup E_{T})$ is a galled tree.
 {proofsketch}
-In the forward direction, if $T$ can be completed into a galled tree $N^{\prime}$ that explains $\mathcal{C}$, we can argue that for every transfer edge $(u,v)$ the is present in the redundancy-free network $N$, there is a corresponding transfer edge in $N^{\prime}$ whose endpoints are either $u$ and $v$, or in their support tree descendants.
+In the forward direction, if $T$ can be completed into a galled LGT network $N^{\prime}$ that explains $\mathcal{C}$, we can argue that for every transfer edge $(u,v)$ that is present in a redundancy-free network $N$, there is a corresponding transfer edge in $N^{\prime}$ whose endpoints are either $u$ and $v$, or $u$ is a support tree descendant of the sending end of that transfer edge in $N$.
 Roughly speaking, this means that the cycles of $N$ have corresponding cycles in $N^{\prime}$, and since $N^{\prime}$ is a galled tree, so is $N$.
@@
 ($\Rightarrow$) Suppose that $T$ is galled-completable and let $N^{\prime}=(V^{\prime},E_{S}^{\prime}\cup E_{T}^{\prime})$ be a galled-completion for $T$ that explains $\mathcal{C}$.
+Let $N=(V,E_{S}\cup E_{T})$ be any redundancy-free network of $T$.
 We need to show that $N$ is a galled tree.
@@
 Since $N$ also subdivides every edge once, this allows us to assume that $N^{\prime}$ and $N$ have the same set of nodes, putting the nodes in correspondence without ambiguity.
-We first build an injective map, that associates each transfer edge $(u^{\prime},v^{\prime})$ of $N$ with a distinct transfer edge $(x,y)$ of $N^{\prime}$, such that $y\in\{u^{\prime},v^{\prime}\}$.
+We first build an injective map, that associates each transfer edge $(u^{\prime},v^{\prime})$ of $N$ with a distinct transfer edge $(x,y)$ of $N^{\prime}$, such that $y\in\{u^{\prime},v^{\prime}\}$, and such that $x$ descends from the node of $\{u^{\prime},v^{\prime}\}\setminus\{y\}$ in $\mathcal{T}_{N^{\prime}}$.
 Let $(u^{\prime},v^{\prime})\in E_{T}$.
-Note that by the construction of $N$, $u^{\prime},v^{\prime}$ are subdivision nodes of $\mathcal{T}_{N}$ and are parents of FAs $u,v$.
+Note that by the construction of $N$, $u^{\prime},v^{\prime}$ are subdivision nodes of $\mathcal{T}_{N}$ and, in that support tree, they are parents of nodes $u,v\in V(T)$.
 Moreover, $u,v$ are FA neighbors, and thus there is some character $C$ with $\alpha_{T}(C)=\{u,v\}$.
-By Lemma 3, in $N^{\prime}$ there is a transfer edge $(x,y)$, where either $x\preceq_{\mathcal{T}_{N}}u^{\prime}$ and $y=v^{\prime}$, or $x\preceq_{\mathcal{T}_{N}}v^{\prime}$ and $y=u^{\prime}$.
-Either way, we map $(u^{\prime},v^{\prime})$ to $(x,y)$.
+By Lemma 3, in $N^{\prime}$ there is a transfer edge $(x,y)\in E^{\prime}_{T}$, where either $x\preceq_{\mathcal{T}_{N}}u^{\prime}$ and $y=v^{\prime}$, or $x\preceq_{\mathcal{T}_{N}}v^{\prime}$ and $y=u^{\prime}$.
+Either way, we map $(u^{\prime},v^{\prime})$ to $(x,y)$ (note that our desired properties on the map hold).
 We claim that no other transfer edge of $N$ can be mapped to $(x,y)$ in this manner.
 Suppose that some other transfer edge $(u^{\prime\prime},v^{\prime\prime})$ of $N$ maps to $(x,y)$.
-According to our map, $y=u^{\prime\prime}$ or $y=v^{\prime\prime}$, but either way, $y$ is incident to two transfer arcs in $N$, contradicting Lemma 3.
-Therefore, only $(u^{\prime},v^{\prime})$ can be mapped to $(x,y)$.
-Next, consider $(u^{\prime},v^{\prime})\in E_{T}$ and its associated edge $(x,y)$ in $N^{\prime}$.
-Suppose that $(x,y)=(\tilde{u},v^{\prime})$ for some descendant $\tilde{u}$ of $u^{\prime}$ in $\mathcal{T}_{N^{\prime}}$.
+According to our map, $y=u^{\prime\prime}$ or $y=v^{\prime\prime}$, but either way, $y$ is incident to two transfer edge $(u^{\prime},v^{\prime})$ and $(u^{\prime\prime},v^{\prime\prime})$ in $N$, contradicting Lemma 3.
+Therefore, only $(u^{\prime},v^{\prime})$ can be mapped to $(x,y)$ and the mapping is indeed injective.
+Next, consider $(u^{\prime},v^{\prime})\in E_{T}$ and its associated edge $(x,y)\in E^{\prime}_{T}$ in $N^{\prime}$.
+According the the properties of our map, suppose first that $(x,y)=(\tilde{u},v^{\prime})$ for some descendant $\tilde{u}$ of $u^{\prime}$ in $\mathcal{T}_{N^{\prime}}$.
 Consider the network $\tilde{N}$ obtained from $N^{\prime}$ by removing $(\tilde{u},v^{\prime})$ and inserting $(u^{\prime},v^{\prime})$ (which does nothing if $\tilde{u}=u^{\prime}$).
@@
 Since $H$ did not intersect with other underlying cycles in $N^{\prime}$, this new underlying cycle does not either, and it follows that $\tilde{N}$ is a galled tree as well.
-The same idea applies if $y=u^{\prime}$ instead.
-Since all underlying cycles of $N^{\prime}$ are independent, we can apply the transformation from the previous paragraph to replace every mapped transfer edge of $N^{\prime}$ by a transfer edge of $N$, one after another, while maintaining the property that no cycles intersect.
+The same idea applies if $y=u^{\prime}$ and $x\preceq_{\mathcal{T}_{N^{\prime}}}v^{\prime}$ instead, that is, we can remove $(x,y)$ from $N^{\prime}$ and add $(u^{\prime},v^{\prime})$ and the result $\tilde{N}$ is still galled.
+Since all underlying cycles of $N^{\prime}$ are independent, we can apply the transformation from the previous paragraph to insert into $N^{\prime}$ every transfer edge of $N$, one after another, while maintaining the property that no cycles intersect.
 Notice that because each $(u^{\prime},v^{\prime})\in E_{T}$ maps to a distinct transfer edge of $N^{\prime}$, every edge of $N$ is incorporated, and no edge of $N^{\prime}$ is “moved” twice by this process.
@@
 It then follows that $N$ is a galled tree as well.
-()($\Leftarrow$)( ⇐ ) Suppose that the redundancy-free network $N$ is a galled tree.
+()($\Leftarrow$)( ⇐ ) Take some redundancy-free network $N$ of $T$ and assume that $N$ is a galled tree.
 We show that for every character in $C\in\mathcal{C}$ there exists a node in $N-F_{C}(N)$ that reaches every leaf in $C$ as required by Lemma 1.
@@
 Denote $p_{u}=p_{\mathcal{T}_{N}}(u)$ and $p_{v}=p_{\mathcal{T}_{N}}(v)$, where $p_{u}$ and $p_{v}$ are transfer nodes since every edge of $T$ was subdivided to produce $N$.
-Note that by definition, in $\mathcal{T}_{N}$ all the descendants of $p_{u}$ and of $p_{v}$ are in $C$, and thus both of them appear in $N-F_{C}(N)$.
-Suppose that one of the edges $(p_{u},p_{v})$ or $(p_{v},p_{u})$ is present.
+Note that by definition, in $\mathcal{T}_{N}$ all the descendants of $p_{u}$ and of $p_{v}$ are in $C$, and thus both $p_{u}$ and $p_{v}$ appear in $N-F_{C}(N)$.
+Suppose that one of the transfer edges $(p_{u},p_{v})$ or $(p_{v},p_{u})$ is present in $N$.
 In the first case, $p_{u}$ is an origin for $C$ and in the second case, $p_{v}$ is an origin.
 So suppose that neither transfer edge is present.
-Then by construction, either $u$ or $v$ has at least two FA neighbors (otherwise, we would have added one of the two edges between their parent in an arbitrary direction and be in the previous case).
+Because $u$ and $v$ are FA neighbors, this is only possible if either $u$ or $v$ has at least two FA neighbors (otherwise, we would have added one of the two edges between their parent in an arbitrary direction and be in the previous case).
 Suppose without loss of generality that $u$ has at least two FA neighbors.
@@
 2       //$T$ is a tree on taxa set $\mathcal{S}$, $\mathcal{C}$ is the character set.
-3       Let $T^{\prime}$ be the tree that results from subdividing every edge in $T$, let $N=T^{\prime}$
-4       // $FA(v)$ is a set that will keep track of FA neighbors present in $v$. for  $C\in\mathcal{C}$ do
+3       Let $T^{\prime}$ be obtained from $T$ by subdividing every edge, let $N=T^{\prime}$
+4       Initialize $FA(v)=\emptyset$ for all $v\in V(N)$, used to store the FA neighbors of $v$.
+5       for  $C\in\mathcal{C}$ do
 6             Compute $\alpha_{T}(C)$, the set of FAs of $C$ in $T$
@@
 8             if $\alpha_{T}(C)=\{u,v\}$ then add $u$ to $FA(v)$ and add $v$ to $FA(u)$
-8
+9
 10      for $v$ in postorder($T$) do
@@
 14                         return “not galled-completable”
-14
-15                   if $u\prec c_{min}$ then set $c_{min}=u$
-16
+15
+16                   if $u\prec_{T}c_{min}$ then set $c_{min}=u$
+17
 18            if $|FA(v)|\geq 2$ then
 19                   Add $(p_{N}(c_{min}),p_{N}(v))$ to $N$
-19
+20
 21            else
@@
 23                         Add $(p_{N}(v),p_{N}(c_{min}))$ to $N$.
-23
+24
 25                  else
 26                         Mark $v$ as “possibly simple”
-26
 27
+28
 29      if  $N$ is a galled tree then return $N$
 30       else return “not galled-completable”
-30
+31
 Algorithm 1 Check if a given tree $T$ is galled-completable.
 The previous lemma implies a polynomial time verification algorithm, detailed in Algorithm 1.
-We build the redundancy-free network and verify that it is a galled tree.
+We build a redundancy-free network and verify that it is a galled tree.
 We can calculate the set of FAs for each character and use this information to assign the FA neighbors to each node in $T$.
@@
 The rest of the algorithm only adds a time of $O(|V(T)|+|\mathcal{C}|)$.
+Note that verifying whether a given network is a galled tree, as required in Line 1 can be done in time $O(|V|)$.
+Indeed, in a galled tree $N=(V,E)$ the number of edges is $O(|V|)$.
+Moreover, galled trees are the (binary) networks whose biconnected components contain at most one reticulation node (see e.g.
+[gusfield2014recombinatorics, Chapter 8], a biconnected component is a subgraph that cannot be disconnected by removing a single vertex, and a reticulation is a node of indegree $2$).
+We can thus use Tarjan’s algorithm to find biconnected components in linear time [tarjan1972depth], then verify that each of them contains at most one reticulation node.
+Since those components are vertex-disjoint, going through all of them takes time $O(|V|)$.
 We leave the problem of computing FAs in linear time, if at all possible, for a future discussion.
@@
 However, if $v$ has two incomparable FA neighbors, then $v$ must also have at least two minimal incomparable FA neighbors.
-This then implies that the redundancy-free network $N$ will have two transfer edges incident to $p_{N}(v)$, a contradiction of Lemma 3.
+This then implies that a redundancy-free network $N$ will have two transfer edges incident to $p_{N}(v)$, a contradiction of Lemma 3.
 Thus, Line 1 allows us to rule out these cases.
@@
 Finally, the last verification on Line 1 can be done in time $O(|V|)$ (see e.g.
-[gusfield2014recombinatorics]).
+[gusfield2014recombinatorics] and main text).
 ∎∎
 4 The Galled Compatibility Problem
-Let us recall the galled compatibility problem: we are given a set of characters $\mathcal{C}$ on a set of taxa $\mathcal{S}$ and must decide whether a galled tree $N$ explains $\mathcal{C}$.
+Let us recall the galled compatibility problem: we are given a set of characters $\mathcal{C}$ on a set of taxa $\mathcal{S}$ and must decide whether a galled LGT network $N$ explains $\mathcal{C}$.
 Note that if such an $N$ exists, then its base tree is galled-completable.
@@
 Figure 5: An example instance with characters $\mathcal{C}=\{A,B,C,X,Y,P,Q,R\}$ on taxa $\mathcal{S}=\{a,b,c,d,e,f,g,h,i,j,k,l,m\}$ used to illustrate the main steps 1-4 of the algorithm.
+Note that the maximal characters are: $A,B,C,Q$ and $R$.
 Step (1) splits $A$ into $A\setminus B,A\cap B$.
-Step (2) integrates the clades forced by $X,Y$, which intersect the two clades of $A$.
+Step (2) integrates the characters $X,Y$ that intersect the two clades of $A$.
 Step (3) integrates $B$ and $C$ as clades.
-Step (4) solves for $P$ recursively, and then for $Q,R$ recursively, then the transfer required by $A,X$, and $Y$ is added.
+Step (4) solves for $P$ recursively, and then for $Q,R$ recursively.
+Note that transfer edges are not part of the output of the algorithm, as it only returns a galled-completable tree.
+The transfer edges were added in the subfigure to show that the resulting tree was indeed galled-completable — for instance the transfer edge from $\{d,e\}$ to $\{f,g,h\}$ is required to explain $A,X,Y$.
+We now present our polynomial-time algorithm to reconstruct a galled-completable tree $T$ for $\mathcal{C}$, if one exists.
+For each character $C\in\mathcal{C}$, we must decide whether $C$ should be a clade in $T$, or whether we want to split $C$ into two different clades that will eventually be joined by a transfer edge in a completion of $T$.
 The detailed algorithm is somewhat involved, but the main idea can be described in a few steps that are illustrated in Figure 5:
@@
 (1) otherwise, every maximal character $A$ has some incompatibility with some $B$, as in Figure 5.
-We can show that $T$ must either split $A$ into clades $\{A\setminus B,A\cap B\}$, or split $B$ into clades $\{B\setminus A,B\cap A\}$.
+We choose an arbitrary pair of incompatible characters $\{A,B\}$ such that both $A$ and $B$ are maximal (such a pair is shown to exist).
+We can show that $T$ must either split $A$ into clades $\{A\setminus B,A\cap B\}$ and keep $B$ as a clade, or split $B$ into clades $\{B\setminus A,B\cap A\}$ and keep $A$ as a clade.
 We try the first option, and if it leads to a dead-end we try the other option.
@@
 This means that in any completion, there will be a transfer edge between the two clades to provide an origin for $A$;
-(2) each character that intersects both clades must use that same transfer edge in a completion.
-Such characters enforce new clades that refine the previous ones.
-For example in Figure 5.2, $X$ enforces the clade $\{d,e\}$, which is the intersection $X\cap\{a,b,c,d,e\}$, and similarly $Y$ enforces $\{c,d,e\}$.
-These enforced new clades must be ordered by inclusion;
-(3) other characters can enforce further clades, namely those that contain the minimal clades enforced so far.
-For example, the character $C$ from Figure 5 contains $\{d,e\}$, and so the clade $C$ is enforced.
+(2) We next scan for clades that are forced by the above choice, i.e., clades that must be present in any galled-completable tree that contains clades $A\setminus B,A\cap B$.
+We show that if a character intersects both these clades, then both intersections are forced clades.
+For example in Figure 5.2, $X$ enforces the clades $X\cap(A\setminus B)=\{d,e\}$ and $X\cap(A\cap B)=\{f,g,h\}$.
+Similarly $Y$ enforces $Y\cap(A\setminus B)=\{c,d,e\}$ and $Y\cap(A\cap B)=\{f,g,h\}$;
+(3) We can find further forced clades.
+Namely, the characters that contain clades enforced so far are shown to be forced clades.
+For example, the character $C$ from Figure 5 is a superset of the clade $\{d,e\}$ enforced in the previous step, and so the clade $C$ is enforced.
 Likewise, $B$ is forced since it contains $\{f,g,h\}$.
 These are added in Figure 5.3;
-(4) it turns out that if $\mathcal{C}$ is galled-compatible, then any unforced character so far represents a set of leaves that have the same parent $v$ in $T$.
+(4) it turns out that if $\mathcal{C}$ is galled-compatible, then any character that has not implied a forced clade so far represents a set of leaves that have the same parent $v$ in $T$.
 See $P,Q,R$ in Figure 5.3.
-We can recurse into the leaf children of each $v$ and replace the leaves by the resulting network.
-In Figure 5.4, the leaf set $\{i,j,k\}$ is replaced by a network that explains $Q,R$, and $\{f,g,h\}$ by a network that explains $P$.
+We can recurse into the leaf children of each $v$ and replace the leaves by a galled-completable subtree with respect to that set of leaves.
+In Figure 5.4, the leaf set $\{i,j,k\}$ is replaced by a tree that is galled-completable for $Q,R$, and $\{f,g,h\}$ by a tree that is galled-completable for $P$.
 An important subtlety arises in Step (1).
@@
 {proofsketch}
-If $C$ as stated exists, we can start constructing a tree whose root has two children, one with $C$ as a clade, and the other with everything not in $C$.
+The forward direction is immediate.
+For the converse, if $C$ as stated exists, we can start constructing a tree whose root has two children, one with $C$ as a clade, and the other with everything not in $C$.
 This automatically explains $C$.
 We can then recursively solve for $\mathcal{C}_{1}$ on the $C$ side, and for $\mathcal{C}_{2}$ on the non-$C$ side, and replace the child clades with the corresponding networks.
-Because $C$ is compatible, every character will be on one side, this solves for every character.
+Because $C$ is compatible with every character, they will all be in either $\mathcal{C}_{1}$ or $\mathcal{C}_{2}$.
+If the latter two are galled-compatible, we can easily merge the networks that explains each subset as in Figure 6, since characters in $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ have no taxa in common.
 ∎
@@
 So assume that both $A$ and $B$ are split, having two FAs $a_{1},a_{2}$ and $b_{1},b_{2}$ each.
-Because $A$ and $B$ intersect, with some effort it can be shown that these FAs must be related by ancestry (i.e., $a_{1}\prec b_{1}$ or $b_{1}\prec a_{1}$, and also $a_{2}\prec b_{2}$ or $b_{2}\prec a_{2}$).
+Because $A$ and $B$ intersect, with some effort it can be shown that these FAs must be related by ancestry (i.e., $a_{1}\prec_{T}b_{1}$ or $b_{1}\prec_{T}a_{1}$, and also $a_{2}\prec_{T}b_{2}$ or $b_{2}\prec_{T}a_{2}$).
 The proof shows that each possible case leads to requiring two distinct transfers to explain $A$ and $B$, which create intersecting cycles.
@@
 Let $a_{1},a_{2}$ and $b_{1},b_{2}$ be the FA nodes of $A$ and $B$, respectively, and let $A_{1},A_{2},B_{1},B_{2}$ be the respective clades of $a_{1},a_{2},b_{1},b_{2}$.
-We claim that some $a_{i}$ is a strict ancestor of some $b_{j}$ or vice-versa.
+We claim that in $T$, some $a_{i}$ is a strict ancestor of some $b_{j}$ or vice-versa.
 Note that $A\cap B\neq\emptyset$, so we may assume without loss of generality that $A_{1}\cap B_{1}\neq\emptyset$.
@@
 If $a_{1}\neq b_{1}$, our claim holds.
-If $a_{1}=b_{1}$, then $a_{2}\prec b_{2}$ or $b_{2}\prec a_{2}$ by Lemma 3 (part 2), and again our claim holds.
+If $a_{1}=b_{1}$, then $a_{2}\prec_{T}b_{2}$ or $b_{2}\prec_{T}a_{2}$ by Lemma 3 (part 2), and again our claim holds.
 We may therefore assume, without loss of generality, that $a_{1}$ is a strict ancestor of $b_{1}$, which implies $B_{1}\subseteq A_{1}$.
 Notice that $b_{2}$ cannot be a descendant of $a_{1}$, since otherwise we would have $B_{2}\subseteq A_{1}$ and $B_{1}\cup B_{2}\subseteq A_{1}$, contradicting the incompatibility of $A$ and $B$.
-Since $b_{2}$ is incomparable with $b_{1}$, it also cannot be an ancestor of $a_{1}$, and thus $b_{1}$ is also incomparable with $a_{1}$.
+Since $b_{2}$ is incomparable with $b_{1}$, it also cannot be an ancestor of $a_{1}$, and thus $b_{2}$ is also incomparable with $a_{1}$.
 Then by Lemma 3 (part 1), $a_{2}=b_{2}$.
@@
 Let $b$ be the (unique) FA of $B$.
-If $b\preceq a_{1}$ or $b\preceq a_{2}$, then $B\subseteq A$ and $A,B$ would be compatible.
+If $b\preceq_{T}a_{1}$ or $b\preceq_{T}a_{2}$, then $B\subseteq A$ and $A,B$ would be compatible.
 Moreover, $B$ intersects with at least one of $A_{1}$ and $A_{2}$ since it intersects with $A$, and thus $b$ is an ancestor of $a_{1}$ or $a_{2}$.
@@
 This can be formalized as follows.
+Definition 4 ($(A_{1},A_{2})$-chains.).
 Let $A$ be a character and let $\{A_{1},A_{2}\}$ be a partition of $A$.
@@
 We call $X_{1}\cap A_{1}$ the bottom of the chain, and we call $A_{2}$ the stable side of the chain.
-In Figure 5.3, $A_{1}=\{a,b,c,d,e\},A_{2}=\{f,g,h\}$ and $\mathcal{X}=\{X,Y,A\}$.
-One can see that $\mathcal{X}$ forms an $(A_{1},A_{2})$-chain with bottom $\{d,e\}$ and stable side $\{f,g,h\}$.
-The concept of a chain is illustrated in Figure 7.
+Consider Figure 5.3, with character $A$ partitioned into $A_{1}=\{a,b,c,d,e\},A_{2}=\{f,g,h\}$ and $\mathcal{X}=\{X,Y,A\}$.
+One can see that $\mathcal{X}$ forms an $(A_{1},A_{2})$-chain.
+Indeed, we have $X\cap A_{1}\subset Y\cap A_{1}\subset A\cap A_{1}$ because these intersections give $\{d,e\}\subset\{c,d,e\}\subset\{a,b,c,d,e\}$.
+Moreover, $X\setminus A_{1}=Y\setminus A_{1}=A\setminus A_{1}=\{f,g,h\}$.
+The bottom of this chain is $\{d,e\}$ and the stable side $\{f,g,h\}$.
+The general concept of a chain is illustrated in Figure 7 (the transfer edge can be ignored).
 Figure 7: An $(A_{1},A_{2})$-chain, where $X_{1},X_{2},X_{3}\in\mathcal{X}$ all intersect with both $A_{1}$ and $A_{2}$.
 On the $A_{1}$ side, the FAs are ordered by ancestry, and the only way to explain all these characters is to have them use the same transfer to reach the $A_{2}$ side.
+Note, the transfer edge is not part of the chain nor the reconstruction, we add it to emphasize how the characters forming the chain can be explained.
 {lemmarep}
@@
 Then, after possibly exchanging the subscripts of $A_{1}$ and $A_{2}$, $\mathcal{X}$ is an $(A_{1},A_{2})$-chain.
-Moreover, for every $X\in\mathcal{X}$, the clades $X\cap A_{1}$ and $X\cap A_{2}$ are in $T$.
+Moreover, for every $X\in\mathcal{X}$, the clades $X\cap A_{1}$ and $X\cap A_{2}=A_{2}$ are in $T$.
 {proofsketch}
@@
 Suppose that $T$ is a galled-completable tree for $\mathcal{C}$ in which $A$ is split into the clades $A_{1}$ and $A_{2}$.
-Let $\mathcal{X}\subseteq\mathcal{C}$ be the subset of characters that intersect with both $A_{1}$ and $A_{2}$ and suppose that $\mathcal{X}$ is an $(A_{1},A_{2})$-chain.
+Let $\mathcal{X}\subseteq\mathcal{C}$ be the subset of characters that intersect with both $A_{1}$ and $A_{2}$ and suppose that $\mathcal{X}$ is a $(A_{1},A_{2})$-chain.
 Let $X_{1}\cap A_{1}$ be the bottom of the chain and let $A_{2}$ be the stable side of the chain.
@@
 To put this precisely, let us gather all the information on the tree that has been shown to be forced so far.
+Definition 5 (Forced characters and clades.).
 Let $A$ be a maximal character of $\mathcal{C}$ and let $\{A_{1},A_{2}\}$ be a partition of $A$ into two non-empty sets.
@@
 Furthermore, a clade $Y\subseteq\mathcal{S}$ is forced by $\{A_{1},A_{2}\}$ if either: $Y=X\cap A_{1}$ or $Y=X\cap A_{2}$ for some $X\in\mathcal{X}$; or $Y=C$ for some character $C\in\mathcal{C}$ that contains the bottom or the stable side of the $\mathcal{X}$ chain.
-The first type of clades are enforced due to Lemma 4, and the second type due to Lemma 4.
+For example in Figure 5.2, we made the decision of partitioning $A$ into $A_{1}$ and $A_{2}$, and $\mathcal{X}=\{X,Y,A\}$ forms an $(A_{1},A_{2})$-chain.
+The definition states that all characters of $\mathcal{X}$ are forced.
+Moreover, character $C$ is forced because it is a superset of the bottom of the chain $\{d,e\}$, and character $B$ is forced because it contains the stable side $\{f,g,h\}$.
+We call these characters forced because by Lemma 4 and Lemma 4, they imply the existence of clades in $T$.
+In turn, the clades that are implied are called forced.
+For example, character $X$ in the figure implies the existence of the forced clade $X\cap A_{1}=\{d,e\}$ in $T$, and character $C$ implies the existence of forced clade $C$ in $T$.
 The next lemma is crucial: it shows that non-forced characters can be grouped and dealt with recursively according to the tree that contains the forced clades.
@@
 We assume that some characters $B,D,E$ enforced other clades.
-If we assume that some character $C$ has elements with distinct parents, the situation on the right is unavoidable.
-A transfer will be needed to link the bottom and stable side, and another transfer to explain $C$.
+If we assume that some character $C$ has two taxa with distinct parents (gold circles in the figure), the situation on the right is unavoidable.
+That is, in any galled-completion of $T$, a transfer will be needed to link the bottom and stable side, and another transfer to explain $C$.
 These two transfers create intersecting underlying cycles.
@@
 As for the other characters, for a node $v$ of $T$, let $\mathcal{C}_{v}\subseteq\mathcal{C}$ be the non-forced characters having taxa whose parent is $v$.
-We can find a galled network $N_{v}$ that explains $\mathcal{C}_{v}$ and “attach” $N_{v}$ as a child of $v$.
+We can find a galled LGT network $N_{v}$ that explains $\mathcal{C}_{v}$ and “attach” $N_{v}$ as a child of $v$.
 Doing this for every $v$ does not create intersecting cycles and explains all the remaining characters.
@@
 {lemmarep}
-Suppose that $\mathcal{C}$ has no maximal character that is compatible with all the others.
+Suppose that $\mathcal{C}$ has no maximal character that is compatible with all the other characters.
 Then there exists a pair of incompatible characters $\{A,B\}$ of $\mathcal{C}$ such that $A$ and $B$ are both maximal.
@@
 For simplicity, it only solves the decision version of the problem (i.e., whether the characters are galled-compatible or not), but it can easily be adapted to reconstruct a network.
-1 fonction getGalledTree($\mathcal{C}$)
+1 function getGalledTree($\mathcal{C}$)
 2       if $\mathcal{C}=\emptyset$ then return $true$
@@
 18
-19 fonction tryPartition($C,\{A_{1},A_{2}\}$)
+19 function tryPartition($C,\{A_{1},A_{2}\}$)
 20       Let $\mathcal{X}$ be the characters that intersect $A_{1}$ and $A_{2}$
@@
   3. 3.
-    We know that a galled-completable tree, if one exists, must contain the clades $A\setminus B,A\cap B$, or $B\setminus A,B\cap B$.
+    We know that a galled-completable tree, if one exists, must contain the clades $A\setminus B,A\cap B$, or $B\setminus A,A\cap B$.
 We do not know which one it is, so we try the former first.
@@
 When $getCalledTree$ receives this answer, it attempts to check whether using $\{B_{1},B_{2}\}=\{B\setminus A,B\cap A\}$ works instead, which is the only other possibility.
-    Do note that Lemma 4 only applies to maximal characters. Since we potentially check them on $A$ and $B$, it is crucial to have both $A$ and $B$ as maximal characters.
+    Do note that Lemma 4 only applies to maximal characters.
+Since we potentially check the four conditions of the lemma on $A$ both and $B$, it is crucial to have both $A$ and $B$ as maximal characters.
   5. 5.
@@
 Leaf names correspond to the species id present in the KEGG orthology database.
-Notice that the redundancy-free network is not galled since the cycles resulting from joining the first appearances of characters $K13628$ and $K00561$ contain the root of the tree as a common node.
+Notice that the resulting redundancy-free network is not galled since the cycles resulting from joining the first appearances of characters $K13628$ and $K00561$ contain the root of the tree as a common node.
 However as shown in (b), by resolving one of the children of the root node, as the black dot, this network becomes galled.
@@
 This is shown in Figure 12 on the left.
-As explained in the caption, if we built the redundancy-free network for these characters and applied Lemma 3, we would add a transfer edge linking the branches above the roots of the blue character, and above the roots of the green character.
+As explained in the caption, if we built a redundancy-free network for these characters and applied Lemma 3, we would add a transfer edge linking the branches above the roots of the blue character, and above the roots of the green character.
 These are non-redundant edges that create two cycles that intersect at the root.
@@
 6 Conclusion
-In this work, we expanded the foundations on perfect phylogenies to galled trees.
-While compatible character in trees enjoy an elegant mathematical structure, it appears that the difficulty of characterizing compatibility ramps up very quickly as we move beyond trees.
+In this work, we expanded the foundations on perfect phylogenies to galled perfect transfer networks.
+While compatible characters in trees enjoy an elegant mathematical structure, it appears that the difficulty of characterizing compatibility ramps up very quickly as we move beyond trees.
 Nonetheless, this work can serve as a stepping-stone towards the understanding of character evolution on more complex structures.
@@
 It will also be interesting to integrate broader classes of networks, for example level-$k$ networks, and to solve the open problem of finding a PTN with a minimum number of transfer edges.
+In that regards, it is possible that the algorithms from [vanIersel2010, kelk2012] for explaining softwired clusters could be adapted to PTN, as only differ by restricting which reticulation edges can be switched off or not.
 Perhaps a more important direction, as exemplified by our case study, is to remove the “no-loss” assumption on characters.